### v1

# Contextual Stochastic Optimization for Order Fulfillment Optimization

[…]
Contextual Stochastic Optimization for Order Fulfillment Optimization
[…]
Contextual Stochastic Optimization for Omnichannel Multi-Courier Order Fulfillment Under Delivery Time Uncertainty
[…]
Tinghan Ye, Sikai Cheng, Amira Hijazi, Pascal Van Hentenryck
[…]
NSF AI Institute for Advances in Optimization (AI4OPT), Georgia Institute of Technology, Atlanta, GA
[…]
Problem definition: The paper studies a large-scale order fulfillment problem for a leading e-commerce company in the United States. The challenge involves selecting fulfillment centers and shipping carriers with observational data only to efficiently process orders from a vast network of physical stores and warehouses. The company’s current practice relies on heuristic rules that choose the cheapest fulfillment and shipping options for each unit, without considering opportunities for batching items or the reliability of carriers in meeting expected delivery dates. Methodology / results: The paper develops a data-driven Contextual Stochastic Optimization (CSO) framework that integrates distributional forecasts of delivery time deviations with stochastic and robust order fulfillment optimization models. The framework optimizes the selection of fulfillment centers and carriers, accounting for item consolidation and delivery time uncertainty. Validated on a real-world data set containing tens of thousands of products, each with hundreds of fulfillment options, the proposed CSO framework significantly enhances the accuracy of meeting customer-expected delivery dates compared to current practices. It provides a flexible balance between reducing fulfillment costs and managing delivery time deviation risks, emphasizing the importance of contextual information and distributional forecasts in order fulfillment. Managerial implications: This is the first paper that studies the omnichannel multi-courier order fulfillment problem with delivery time uncertainty through the lens of contextual optimization, fusing machine learning and optimization. The results provide actionable insights for e-commerce companies and online retailers to enhance service quality and customer satisfaction through efficient order fulfillment strategies that accounts for delivery time uncertainty. Adopting this framework can balance cost effectiveness with customer service, potentially leading to higher customer retention and profitability.
[…]
logistics, order fulfillment, machine learning, contextual stochastic optimization, contextual robust optimization
1 Introduction
Over the past decade, online shopping has surged dramatically. The e-commerce share of total retail sales in the United States has risen from 6% in 2015 to 16% in 2024, exceeding 1.1 trillion dollars (US Department of Commerce 2024). Amid intense competition in pricing, service, and marketing, driven by the lucrative e-commerce market, logistical performance has become a key element to success. Last-mile fulfillment, in particular, is critical as it accounts for 55% of the total transportation costs in the fulfillment process (Viswanathan 2023). This evolving market landscape poses unprecedented challenges for e-commerce companies striving to deliver satisfactory customer services. Fortunately, the growing volume of data also offers opportunities for data-driven operational improvements, enabling companies to boost revenue and cut costs (Fisher et al. 2019).
In addition, after the COVID-19 pandemic, the demand for timely home delivery has also increased significantly. As a result, omnichannel fulfillment and the use of multiple couriers, including crowd-shipping, have become essential components of the fulfillment landscape, drawing significant attention both in industry and academia (Dethlefs et al. 2022, Mohri et al. 2023, Das et al. 2023). The omnichannel approach combines distribution centers (DCs) and brick-and-mortar stores to fulfill online orders. Unlike the traditional method that uses DCs for online orders and stores for in-store purchases separately, the omnichannel approach leverages the proximity of physical stores to delivery addresses. This can potentially offer more cost-effective and faster delivery options for online customers (Acimovic and Farias 2019). Meanwhile, in the crowd-shipping model, retailers recruit non-dedicated drivers, such as gig couriers like Roadie, to handle pick-up and delivery tasks for online customers. This approach can alleviate the workload of employed drivers during peak hours and offer faster delivery to customers, resulting in higher service satisfaction (Zehtabian et al. 2022, Behrendt et al. 2023). The problem considered in this paper involves fulfilling orders by choosing from various types of fulfillment centers, such as warehouses and stores, and different couriers, including traditional shipping companies and gig couriers.
Given that transportation and shipping are significant cost drivers in e-commerce (Dethlefs et al. 2022, Kuhn and Sternbeck 2013, Hübner et al. 2013), delivery consolidation has emerged as an effective strategy for fulfilling orders with multiple items. This approach has been shown to enhance consumer satisfaction (Wagner et al. 2023), while also reducing shipping costs for online retailers (Chen et al. 2024). In addition to cutting parcel expenses, item consolidation enhances the customer experience by eliminating the need to receive multiple packages for a single order. This approach not only improves convenience but also promotes a more sustainable and environmentally friendly fulfillment process by decreasing the number of packages shipped and the number of delivery trips needed. This reduction in shipments and trips leads to lower carbon emissions and less packaging waste, contributing to greener logistics practices (Ülkü 2012). This paper also investigates multi-item order consolidation in its omnichannel fulfillment problem.
Most existing studies on order fulfillment optimization have not explicitly considered the uncertainty in delivery timeliness. However, Freedman (2019) highlight the strong desire of customers for accurate delivery time promises and faster service. The survey indicates that around 44% of online shoppers have abandoned their shopping carts because the items they wanted would not arrive on time, while 20% abandon orders with unclear delivery dates. Furthermore, 28% of shoppers are willing to pay for expensive, expedited shipping to receive products at their desired time. The statistics clearly show that customers prioritize service quality, sometimes even more than the product value, when making purchase decisions. Moreover, Salari et al. (2022) point out the importance of considering delivery time uncertainty in order fulfillment to better meet customer expectations and improve satisfaction. Their proposed promised delivery time policy demonstrates a 6.1% increase in sales volume in numerical experiments. Notably, overpromising delivery times can lead to higher return rates, lower repurchase rates, and dissatisfaction with delayed service. Additionally, customers can have mixed sentiments over underpromised delivery times, as early deliveries may not align with their availability. Therefore, missing the exact promised delivery date can negatively impact short-term and long-term sales, both implicitly and explicitly (Chan et al. 2018, Cui et al. 2020, Salari et al. 2022). Given the crucial role of delivery timeliness, the paper incorporates delivery time uncertainty into the omnichannel multi-courier order fulfillment optimization problem through a contextual stochastic optimization (CSO) framework.
One of the key challenges of Omnichannel Multi-Courier Order Fulfillment Optimization comes from the nature of historical data: only observational data is available, i.e., only the delivery time deviations of the selected options are available in the data set. In addition, in real time, these deviations are only observable once the decisions have been taken. The availability of observation data only raises fundamental challenges that are typically not addressed in prior work.
To address these challenges, this paper proposes a Contextual Stochastic Optimization (CSO) framework, encompassing both risk-neutral and robust approaches. Unlike traditional stochastic programming methods that depend on static, predefined probability distributions for uncertainties, CSO combines machine learning techniques with mathematical optimization to dynamically predict uncertainties based on contextual information. Thanks to the availability of contextual information, CSO has the flexibility of generating more tailored and effective decisions, leveraging the most current and relevant information. Such flexibility is particularly beneficial in the dynamic environment of online retail, where conditions and customer behaviors can change rapidly, requiring a more responsive approach to decision making. The effectiveness of CSO has been validated in the existing operations research and management science literature (Mišić and Perakis 2020, Sadana et al. 2024). This approach has proven successful in various applications, including inventory control (Bertsimas et al. 2016, Meller et al. 2018), price and revenue management (Ito and Fujimaki 2016, Perakis et al. 2023), and, more recently, school redistricting (Guan et al. 2024).
1.1 Contributions
The key novelty in this paper is a CSO framework for an omnichannel multi-courier order fulfillment optimization with observational data only. Existing studies on CSO typically assume the availability of a dataset $\mathcal{D}=\{(\mathbf{s}_{1},\mathbf{c}_{1}),(\mathbf{s}_{2},\mathbf{c}_{2}),\ldots\}$, pairing observed covariates with complete uncertain parameter vectors. However, real-world scenarios often involve high-dimensional uncertain parameters where individual data points capture only partial information, particularly when uncertainty measurement depends on the chosen solution, which is the case in the application considered in this paper.
To address the challenge of observational data, this paper introduces a generalized CSO framework addressing these real-world complexities. The novel CSO framework is applied to a sophisticated stochastic omnichannel order fulfillment optimization problem with multiple carrier options and delivery time uncertainties. In collaborating with a major US online retailer, the framework is rigorously evaluated on an industrial order fulfillment dataset featuring a vast, intricate fulfillment network. This application demonstrates the framework capabilities and its potential impact on large-scale, real-world optimization challenges. The contributions of the paper can be summarized as follows.
  1. 1.
    To the best of the authors’ knowledge, this paper is the first study of the order fulfillment problem in the context of omnichannel and multi-courier logistics through the lens of data-driven contextual stochastic optimization.
  2. 2.
    A mixed-integer linear program (MILP) is developed for a novel order fulfillment optimization problem that jointly considers the locations of fulfillment centers and the selection of shipping carriers for online order fulfillment. This MILP accounts for the consolidation of items within each order.
  3. 3.
    Two classes of machine learning models are proposed to learn the delivery time deviations: probabilistic multi-class classification models and quantile regression models that provide distributional forecasts of delivery time deviations.
  4. 4.
    The proposed MILP and machine learning models are integrated into a generalized CSO framework. The framework integrates the conditional distributions of delivery time deviations with stochastic and robust optimization models via sample average approximation (SAA) and tractable reformulations.
  5. 5.
    The paper applies the CSO framework to a large-scale real-world instance, demonstrating its impact on balancing reduced fulfillment costs and increased on-time order deliveries. The results validate the practical effectiveness of contextual optimization in an e-commerce setting, bringing significant benefits compared to the baselines.
1.2 Structure of the Paper
The remaining sections of the paper are organized as follows. Section 2 reviews the related literature. Section 3 formalizes the generic CSO problem addressed in this paper and proposes SAA and robust optimization methods. The methods are applied to an unique order fulfillment problem in Section 4, which describes the MILP model and the machine learning models. Section 5 describes the case study, provides experimental results, and discusses actionable insights obtained from the experiments based on a real dataset. Section 6 concludes the paper.
2 Related Literature
There are several streams of literature related to this paper in terms of methodology and problem context.
2.1 Data-Driven Decision Making
This paper adds to the growing body of literature on data-driven decision-making, particularly at the intersection of machine learning and optimization, which has led to the development of CSO. Sadana et al. (2024) put forth a comprehensive survey on contextual optimization. Tutorials on various methodologies and applications of CSO models in logistics and operations management are available in the works of Mišić and Perakis (2020), Qi and Shen (2022), and Tian et al. (2023). Among the body of research on CSO, the most relevant to this paper are those that focus on learning conditional distributions of uncertain parameters and incorporating these forecasts into optimization models. Data-driven approaches typically build these forecasts from empirical residual errors (Deng and Sen 2022, Kannan et al. 2023, Perakis et al. 2023) or weighted empirical distributions based on proximity to the training data (Bertsimas and Kallus 2020, Notz and Pibernik 2022). However, these studies assume full observation of uncertain parameters for each data point, which differs from the context of this paper. Recent research has also explored integrating downstream optimization effects into ML model training. Elmachtoub and Grigas (2022) introduce the Smart Predict then Optimize (SPO) loss and its convex surrogate for point prediction models, with subsequent extensions to conditional distributional estimation (Qi et al. 2021, Kallus and Mao 2023). However, the SPO loss framework is not directly applicable to this paper’s context, as the dataset analyzed here is observational. In such datasets, outcomes are only partially observed based on decisions made, presenting unique challenges not addressed by these existing methods.
This paper thus contributes to the field of data-driven decision-making that handles observational data, a domain explored by prior research, e.g., in Bertsimas et al. (2019) and Jo et al. (2021) using optimal prescriptive trees for prescriptive analytics in personalized treatment. While these methods effectively learn policies mapping features to actions, they cannot be directly applied to this fulfillment problem considered in this paper due to the presence of instance-specific hard constraints on decisions. The proposed framework addresses these limitations, offering a novel approach to CSO problems with observational data and constrained decision spaces.
Unlike earlier CSO applications that focus on risk-neutral objectives, a growing stream of research is concerned with risk-averse optimization models that incorporate contextual information (Bayram et al. 2022, Peršak and Anjos 2023, Patel et al. 2024, Sun et al. 2024). Building on these studies, this paper also adapts traditional robust optimization models to tackle a unique CSO problem that differs from those explored in other research.
2.2 Order Fulfillment Problem
The novelty in this paper is the application of CSO to a unique instance of the order fulfillment problem. The order fulfillment problem has been extensively studied, and Acimovic and Farias (2019) provide a tutorial on related algorithms developed for addressing the fulfillment optimization problem. In terms of problem context, the most relevant works are those related to omnichannel and multi-item order fulfillment (Xu et al. 2009, Acimovic and Graves 2015, Jasin and Sinha 2015, Zhao et al. 2020, Wei et al. 2021, Ma 2023, Chen et al. 2024). However, these works do not address fulfillment involving a variety of carrier shipping options.
Among the research on order fulfillment, demand uncertainty is the most commonly studied source of randomness (Jasin and Sinha 2015, Zhao et al. 2020, Das et al. 2023, DeValve et al. 2023, Ma 2023). In this paper, demand uncertainty is not considered because the focus is on a single-period order fulfillment problem, rather than a forward-looking approach that spans multiple periods.
A few studies do address delivery time uncertainty at various stages of the order fulfillment process or throughout the entire process. For instance, Raj et al. (2024) propose an integrated queuing model that accounts for delivery time uncertainty in different parts of the order life cycle. They incorporate these uncertainties into an optimization model that minimizes costs while adhering to a delivery probability constraint to ensure on-time delivery.
In contrast to using queuing models, this study is particularly relevant to data-driven, ML-based methods that address for delivery time uncertainty in order fulfillment. Specifically, Liu et al. (2021) integrate point predictions of travel time with a last-mile delivery order assignment optimization problem. They propose both a SAA model and a distributionally robust optimization (DRO) model, with the latter constructing a moment-based ambiguity set. Similarly, Kandula et al. (2021) explore a last-mile delivery problem, focusing on delivery success as the uncertainty. They map success probabilities to delivery times, which are then integrated into a Vehicle Routing Problem with Time Windows. Bayram et al. (2022) study a robust order batching optimization problem in warehouse picking and packing. They use random forest to predict order processing time to form an uncertainty set within a robust optimization model. This paper builds on similar concepts, developing tractable approximations and reformulations of integer programming-based optimization models. More precisely, the uncertainty in the order fulfillment problem considered in this paper is the potential order delivery deviations, which can occur at various stages of the fulfillment process. A similar uncertainty addressed using data-driven approach can be found in Salari et al. (2022), who develop a quantile regression forest-based method to generate distributional forecasts of order delivery times. Their approach then applies a decision rule, designed to balance the asymmetric costs of early and late deliveries, to produce expected arrival times for customers. This paper differs in two key aspects. First, the distributional forecasts are directly incorporated into contextual optimization models, eliminating the need for a separate decision rule to generate point predictions. Second, the CSO framework proposed in this paper is more comprehensive, accommodating multiple carrier options, whereas Salari et al. (2022) focus on a single-carrier scenario.
3 Problem and Methodology
This section describes a generic contextual stochastic optimization problem (denoted by ${\cal P}$) for fulfillment applications where only an observational dataset is available. It then outlines the overall approach taken to address the problem.
3.1 Problem Description
Problem ${\cal P}$ consists in fulfilling a set of incoming requests sequentially, each of which contains orders for a set of products. Each request $o$ has an associated quantity $q_{io}$ for some product $i\in\mathcal{I}$. These products can be fulfilled from a set $\mathcal{V}$ of selection options. When a request $o$ arrives and $o$ contains a subset of products in $\mathcal{I}$, the goal is to determine the quantity of product $i$ assigned to selection option $v$. The decision variables can thus be captured by a vector $\mathbf{z}_{o}=(z_{ivo})_{i\in\mathcal{I},v\in\mathcal{V}}$, where each $z_{ivo}$ indicates the quantity of product $i$ assigned to selection option $v$ in request $o$. Each such assignment must satisfy operational constraints and the set of feasible assignments is denoted by $\mathcal{Z}_{o}$. For each product $i$, each selection option $v$ is associated with an uncertain cost $\tilde{c}_{ivo}\in\mathbb{R}$. The full vector of uncertain costs, governed by a probability distribution $\mathbb{P}$, is denoted by $\mathbf{\tilde{c}}_{o}=(\tilde{c}_{ivo})_{i\in\mathcal{I},v\in\mathcal{V}}$. The objective function $g(\mathbf{z}_{o},\mathbf{\tilde{c}_{o}})$ of the problem depends on these uncertain costs and the decision variables. Problem ${\cal P}$ is thus a risk-neutral stochastic optimization which can be formulated as follows for request $o$:
$\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\mathbf{\tilde{c}}_{o}\sim\mathbb{P}}[g(\mathbf{z}_{o},\mathbf{\tilde{c}}_{o})],$ (1)
where $\mathbb{E}_{\mathbf{\tilde{c}}_{o}\sim\mathbb{P}}[\cdot]$ represents the expectation of the objective function over the probability distribution $\mathbb{P}$.
3.2 Contextual Stochastic Optimization (CSO)
In Problem (1), the distribution $\mathbb{P}$ is not readily available. However, side information is typically available for product $i$ and selection option $v$. Let the vector of covariates for request be $\mathbf{s}_{ivo}$. In the context of an omnichannel order fulfillment problem, these covariates may represent the state of the fulfillment network, the weather, and the current demand for products. All these covariate vectors are aggregated into state vector $\mathbf{s}_{o}=(\mathbf{s}_{ivo})_{i\in\mathcal{I},v\in\mathcal{V}}$, which encompasses all contextual information available to the request at the time of the assignment decision.
The distribution $\mathbb{P}$ is then approximated by the conditional distribution $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ of the cost vector given the observed state vector. This leads to the following Contextual Stochastic Optimization (CSO) problem:
$\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\tilde{\mathbf{c}}_{o}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}[g(\mathbf{z}_{o},\mathbf{\tilde{c}}_{o})],$ (2)
where $\mathbb{E}_{\tilde{\mathbf{c}}_{o}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}[\cdot]$ is the conditional expectation of the objective function given the state vector $\mathbf{s}_{o}$. Let the uncertain cost vector and the aggregated covariates for selection option $v$ be $\tilde{\mathbf{c}}_{vo}=(\tilde{c}_{ivo})_{i\in\mathcal{I}}$ and $\mathbf{s}_{vo}=(\mathbf{s}_{ivo})_{i\in\mathcal{I}}$, respectively. This paper assumes that option costs are independent, i.e.,
$\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})=\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo}),$
where $\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo})$ is the conditional distribution of the uncertain cost $\tilde{\mathbf{c}}_{vo}$ given the covariates $\mathbf{s}_{vo}$. This assumption is reasonable as the uncertain cost for a specific option depends solely on the contextual features of that option, not on the contextual features of other options. Note that, although these conditional distributions are considered independent, selection options that share common covariates (e.g., weather) and thus their distributions will tend to be similar. Problem (2) can then be rewritten as
$\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\tilde{\mathbf{c}}_{o}\sim\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo})}[g(\mathbf{z}_{o},\mathbf{\tilde{c}}_{o})].$ (3)
3.3 Learning the Contextual Distribution
The conditional distribution $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ is also generally unknown but, in practice, it can often be learned from historical data available. In many real-world settings however, each data point is observational, i.e., only those costs involved in the actual decision can be observed. As a result, the dataset available to estimate $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ is of the form $\mathbb{D}=\{(\mathbf{s}_{o},\mathbf{C}_{o})\}_{o=1}^{|\mathbb{D}|}$, where $\mathbf{C}_{o}$ are the observable costs for order $o$. The learning task consists in finding a machine learning $\mathcal{M}$ that approximates $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ in a supervised manner using the dataset $\mathbb{D}$: $\mathcal{M}$ receives a context $\mathbf{s}$ as input and returns a probability distribution $\mathcal{M}(\mathbf{s})$ of costs. Since learning $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ amounts to learning $\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo})$ for each option $v$, the dataset $\mathbb{D}$ is partitioned into subsets, with $\mathbb{D}_{v}=\{(\mathbf{s}_{v1},\mathbf{c}_{v1}),(\mathbf{s}_{v2},\mathbf{c}_{v2}),\ldots\}$ representing the dataset for option $v$. A machine learning model $\mathcal{M}_{v}$ is associated with each option $v$ and $\mathcal{M}(\mathbf{\mathbf{s}})=(\mathcal{M}_{1}(\mathbf{s_{1}}),\ldots,\mathcal{M}_{|\mathcal{V}|}(\mathbf{s}_{|\mathcal{V}|}))$.
3.4 Contextual Sample Average Approximation (C-SAA)
Problem (3) is approximated by generalizing the Sample Average Approximation (SAA) method Kleywegt et al. (2002) to contextual stochastic optimization. The core of the Contextual Sample Average Approximation (C-SAA) for a request $o$ with covariates $\mathbf{s}_{o}$ amounts to
  1. 1.
    sampling the distribution $\mathcal{M}(\mathbf{s}_{o})$ $N$ times to obtain a set $\bm{\xi}_{o}^{n}=(\bm{\xi}_{vo}^{n})_{v\in\mathcal{V}}$ of cost vectors $(n\in[N])$;
  2. 2.
    solving the C-SAA problem
    (C-SAA) $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\frac{1}{N}\sum_{n=1}^{N}g(\mathbf{z}_{o},\bm{\xi}_{o}^{n}).$ (4)
Following Kleywegt et al. (2002), the C-SAA procedure can be outlined as follows:
  1. 1.
    Solve $Q$ independent C-SAA problems, each of sample size $N=N_{1}$, to produce candidate solutions $\mathbf{z}_{1},\ldots,\mathbf{z}_{Q}$ with corresponding objective values $\underaccent{\bar}{f}_{1},\ldots\underaccent{\bar}{f}_{Q}$. Calculate the average objective value as $\underaccent{\bar}{f}=\frac{1}{Q}\sum_{i=1}^{Q}\underaccent{\bar}{f}_{i}$. Let $f^{*}$ be the true optimal objective value of Problem (2). Since $\underaccent{\bar}{f}$ is an unbiased estimator of $\mathbb{E}_{{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}}[\underaccent{\bar}{f}]$, and because $\mathbb{E}_{{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}}[\underaccent{\bar}{f}]\leq f^{*}$, $\underaccent{\bar}{f}$ serves as a statistical lower bound on $f^{*}$ (Kleywegt et al. 2002).
  2. 2.
    Then, re-evaluate the solutions $\mathbf{z}_{1},\ldots,\mathbf{z}_{Q}$ using a larger sample size $N=N_{2}\gg N_{1}$, yielding new objective values $\bar{f}(\mathbf{z}_{1}),\ldots,\bar{f}(\mathbf{z}_{Q})$. Since each candidate solution is feasible, its re-estimated objective values $\bar{f}(\cdot)$ provides an upper bound on $f^{*}$. Therefore, the approximated upper bound on the optimality gap for each candidate solution is $\bar{f}(\cdot)-\underaccent{\bar}{f}$. Finally, select the best solution $\mathbf{z}^{*}$ from the candidates, which is the one with the smallest optimality gap.
It is interesting to compare C-SAA with the approach proposed in Bertsimas and Kallus (2020). A key idea in Bertsimas and Kallus (2020) is to use the historial data (the dataset $\mathbb{D}$ in this paper, as the samples for the stochastic optimization and to weight them properly. Unfortunately, this is not possible in this problem $\mathcal{P}$ considered here, because only observational data is available, i.e., only the costs of the selected options are revealed. One way to proceed would be to define the sample set for the cost vector as all possible combinations of data points across the selection options. However, as the number of selection options and historical data points increase, this approach suffers from the curse of dimensionality and becomes intractable for large dataset with big selection space, which is the case in the application considered in this paper.
3.5 Contextual Robust Optimization (C-RO)
The performance of C-SAA largely hinges on the accuracy of the estimated conditional distribution. To mitigate the risk of model misspecification in forecasting, robust optimization can be employed to account for worst-case scenarios, making it a viable approach for contextual stochastic optimization. Let $\mathcal{U}(\mathbf{s}_{o})$ represent the contextual uncertainty set for the cost vector $\tilde{\mathbf{c}}_{o}$. The Contextual Robust Optimization (C-RO) problem for (3) is then formulated as follows:
(C-RO) $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\max_{\bm{\xi}_{o}\in\mathcal{U}(\mathbf{s}_{o})}g(\mathbf{z}_{o},\bm{\xi}_{o}).$ (5)
To effectively leverage the conditional distributional forecasts and create tractable reformulations of Problem (5), two types of uncertainty sets are considered: a discrete uncertainty set $\mathcal{U}_{d}(\mathbf{s}_{o})$ and a budgeted interval uncertainty set $\mathcal{U}_{b}(\mathbf{s}_{o})$.
3.5.1 Contextual Discrete Uncertainty Set
With a discrete uncertainty set $\mathcal{U}_{d}(\mathbf{s}_{o})$, Problem (5) can be reformulated into a single-level problem by enumerating all possible values of $\tilde{\mathbf{c}}_{o}$ within $\mathcal{U}_{d}(\mathbf{s}_{o})$. This is a common technique in robust optimization models (Goerigk and Khosravi 2024):
[…]
s.t. $t\geq g(\mathbf{z}_{o},\bm{\xi}_{o}),\quad\forall\bm{\xi}_{o}\in\mathcal{U}_{d}(\mathbf{s}_{o}).$
A discrete distribution can be generated as mentioned before, by combining all the datapoints for the selection options. However, this will make the problem intractable as mentioned earlier. For comparison purposes, the results will provide a sampling counterpart to that approach in the same spirit as C-SAA. Since it is not quite satisfactory, another uncertainty set is considered next.
3.5.2 Contextual Budgeted Interval Uncertainty Set
A commonly used uncertainty set is the budgeted interval set $\mathcal{U}_{b}(\mathbf{s}_{o})$, which is defined by an uncertainty budget $B$. This budget limits the extent to which all components of the uncertain parameter $\tilde{\mathbf{c}}_{o}$ can simultaneously deviate to their upper bounds. This section assumes that the uncertain cost $\tilde{\mathbf{c}}_{o}$ is element-wise positive and $g(\mathbf{z}_{o},\tilde{\mathbf{c}}_{o})$ is a piecewise linear function of the following form:
$g(\mathbf{z}_{o},\tilde{\mathbf{c}}_{o})=(\tilde{\mathbf{c}}_{o}+\mathbf{a}_{o}^{j})^{T}\mathbf{z}_{o}\quad\text{if}~{}\mathbf{z}_{o}\in[\mathbf{b}^{j-1}_{o},\mathbf{b}^{j}_{o}],~{}j=1,\ldots n,$ (7)
where $\{\mathbf{b}^{0}_{o},\mathbf{b}^{1}_{o},\ldots,\mathbf{b}^{n}_{o}\}$ are the breakpoints of the piecewise linear function, and $\mathbf{a}_{o}^{j}=(a_{ivo}^{j})_{i\in\mathcal{I},v\in\mathcal{V}}$ is a deterministic cost vector corresponding to the segment $[\mathbf{b}^{j-1}_{o},\mathbf{b}^{j}_{o}]=[(b_{ivo}^{j-1})_{i\in\mathcal{I},v\in\mathcal{V}},(b_{ivo}^{j})_{i\in\mathcal{I},v\in\mathcal{V}}]$. Given these assumptions, the typical approach to ensure tractability when using a budgeted interval set is to dualize the inner maximization problem, resulting in the robust counterpart of the original problem (Bertsimas and Sim 2003, Bayram et al. 2022, Goerigk and Khosravi 2024).
Let the conditional prediction interval of $\tilde{\mathbf{c}}_{vo}$ be $[\bm{\underline{\xi}}_{vo},\bm{\underline{\xi}}_{vo}+\hat{\bm{\xi}}_{vo}]$, where $\bm{\underline{\xi}}_{vo}=(\underline{\xi}_{ivo})_{i\in\mathcal{I}}$ and $\bm{\underline{\xi}}_{vo}+\hat{\bm{\xi}}_{vo}=(\underline{\xi}_{ivo}+\hat{\xi}_{ivo})_{i\in\mathcal{I}}$ represent the lower and upper estimated conditional quantiles of the cost vector, respectively. The budgeted interval set $\mathcal{U}_{b}(\mathbf{s}_{o})$ can then be expressed as follows:
$\mathcal{U}_{b}(\mathbf{s}_{o},B)=\{\bm{\xi}_{o}\in\mathbb{R}^{|\mathcal{I}||\mathcal{V}|}:\xi_{ivo}=\underline{\xi}_{ivo}+\hat{\xi}_{ivo}\delta_{iv},\sum_{i\in\mathcal{I},v\in\mathcal{V}}\delta_{iv}\leq B,\delta_{iv}\in[0,1]\}.$ (8)
For a budgeted uncertainty set, it should be noted that learning the entire conditional distribution is unnecessary. Instead, learning the lower and upper conditional quantiles is sufficient to construct the conditional prediction interval. For instance, to create a 95% conditional prediction interval, only the 2.5% and 97.5% conditional quantiles are required. Let $q^{low}$ and $q^{upp}$ be the lower and upper quantiles for the conditional prediction interval of selection option $v$, respectively. Then, the lower bound $\bm{\underline{\xi}}_{vo}$ is defined as $\inf\{\mathbf{u}:\mathbb{P}_{v}(\mathbf{c}_{v}\leq\mathbf{u}\mid{\mathbf{s}_{vo}})\geq q^{low}\}$, and the upper bound $(\bm{\underline{\xi}}_{vo}+\hat{\bm{\xi}}_{vo})$ is defined as $\inf\{\mathbf{u}:\mathbb{P}_{v}(\mathbf{c}_{v}\leq\mathbf{u}\mid{\mathbf{s}_{vo}})\geq q^{upp}\}$.
With a fixed $\mathbf{z}$, the inner maximization problem of Problem (5) is expressed as follows:
$\max\quad$ $\sum_{i\in\mathcal{I}}\sum_{v\in\mathcal{V}}\hat{\xi}_{ivo}z_{ivo}\delta_{iv}$ (9)
s.t. $\sum_{i\in\mathcal{I},v\in\mathcal{V}}\delta_{iv}\leq B$ (10)
$\delta_{iv}\in[0,1],\quad\forall i\in\mathcal{I},v\in\mathcal{V}.$ (11)
By taking the dual of the inner maximization problem and introducing auxiliary variables, a tractable reformulation of the robust counterpart of Problem (5) is obtained as follows:
[…]
s.t. $\pi+\lambda_{iv}\geq\hat{\xi}_{ivo}z_{ivo},\quad\forall i\in\mathcal{I},v\in\mathcal{V}$
$\pi\geq 0$
$\lambda_{iv}\geq 0,\quad\forall i\in\mathcal{I},v\in\mathcal{V}$
$z_{ivo}\leq b^{j}_{ivo}+M(1-y_{j}),\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
$z_{ivo}\geq b^{j-1}_{ivo}-M(1-y_{j}),\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
$w_{ivo}^{j}\geq(\underline{\xi}_{ivo}+a_{ivo}^{j})z_{ivo}-M(1-y_{j}),\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
$y_{j}\in\{0,1\},\quad\forall j\in[n]$
$w_{ivo}^{j}\geq 0,\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
where $\pi$ and $(\lambda_{iv})_{i\in\mathcal{I},v\in\mathcal{V}}$ are the dual variables corresponding to Constraints (10) and (11), respectively, and $M$ is a reasonably large constant.
4 Application: The Ominichannel Multi-Courier Order Fulfillment Problem
This sections introduces a real-world order fulfillment problem encountered by the industrial partner. The proposed CSO framework is then applied to this problem, demonstrating significant advantages over the current practice and common baseline methods.
4.1 High-Level Problem Description
This section provides a high-level overview of the online order fulfillment challenges faced by the industrial partner. Similar to other major e-commerce companies, tens of thousands of orders are placed continuously throughout the day on the website and await processing. Each incoming order is organized by stock-keeping units (SKUs), resulting in one or more order lines. For each order line, there are hundreds of fulfillment centers across the company network eligible to source the SKUs. These centers include primarily physical stores, as well as regional distribution centers and metro e-commerce centers. Each location is equipped with multiple eligible carriers capable of providing different levels of service (LOS), such as Ground, Same-Day, and Overnight shipping. The set of available carriers spans from major shipping companies to crowdsourced gig carriers. Figure 1 illustrates the order fulfillment life cycle within the operations of the industrial partner.
Figure 1: The Life Cycle of the Order Fulfillment Process.
Before making an order fulfillment decision, the company has access to both the transactional details of the order and the real-time status of the fulfillment network. Specifically, the order information includes the date and time the order was placed, the type, dimensions, weight, and quantities of the SKUs, the shipping destination, and the desired delivery date. The network status data provides details on the on-hand inventory levels of the requested SKUs and the remaining open capacities at various fulfillment centers. In addition, each carrier provides an estimated table for shipping costs and planned time-in-transit.
Given this information, the problem is to decide which fulfillment center to source each order line from and which carrier to use for shipping from that location. The goal is to identify the most cost-effective fulfillment option (a location-carrier pair) while ensuring the customer’s desired delivery date is met. The uncertainty of this problem lies in the delivery timeliness performance of each fulfillment option.
4.1.1 Overview of the Current Practice
Currently, the industrial partner uses a sourcing engine that employs a straightforward greedy algorithm. For each unit in an order, the sourcing engine sorts the eligible location-carrier pairs by the estimated shipping costs and transit times.
There are several limitations with this approach. First, since the engine makes decisions for each unit independently, it misses the opportunity to consolidate multiple units of the same or different products within an order. This can lead to higher shipping costs, as the potential to reduce the number of packages—and thereby lower shipping expenses—is not fully realized.
Second, the method relies solely on static transit time estimates provided by the carriers based on their service levels to select the best carrier for meeting the promised delivery date. This approach assumes that these estimates are always accurate. However, static transit times may not accurately reflect the actual delivery performance of carriers. Historical data from the industrial partner reveals that over 10% of orders experience varying degrees of early or late deliveries, which are called deviations in this paper. Given the complexity of the entire fulfillment process, these deviations can be attributed to numerous factors, such as warehouse pick-and-pack performance, late carrier arrivals, no-shows, and external factors like extreme weather or other unexpected circumstances. Such deviations from customer expectations can negatively impact their experience, potentially harming the company in the long run.
4.2 The Nominal Optimization Problem
This section introduces the nominal MILP model developed for optimizing the omnichannel multi-courier order fulfillment problem, which is a special case of the generic problem described in Section 3.1. A comprehensive list of the notations presented in this section is provided in Appendix for reference.
4.2.1 Main Decision Variables
Let $\mathcal{I}$ be the set of SKUs and $\mathcal{V}$ be the set of fulfillment options. Here, $\mathcal{V}$ is defined as the Cartesian product $\mathcal{K}\times\mathcal{L}$, where $\mathcal{K}=\{1,\ldots,K\}$ denotes the set of carriers, and =$\mathcal{L}=\{1,\ldots,L\}$ denotes the set of locations.
Each order $o$ needs to solve an independent optimization problem to decide the selections of carrier-location pairs for each SKU. The primary variables are denoted by the vector $\mathbf{z_{o}}=(z_{ik\ell o})_{i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}}$, where each $z_{ik\ell o}\in\mathbb{Z}$ indicates the quantity of SKU $i$ sourced from location $\ell$ and shipped by carrier $k$ in order $o$.
4.2.2 Objective Function
The objective of the optimization problem for order $o$ is to minimize the overall fulfillment costs while ensuring timely delivery. The fulfillment costs for an order include the fixed sourcing (non-parcel) costs and the shipping (parcel) costs, with consolidation discounts applied when multiple units are assigned to the same carrier-location pair. Delivery timeliness is measured by deviations from the desired delivery date, whether early or late. To encourage timeliness, the objective imposes penalties for both early and late deliveries.
The objective function is formally defined as follows. Let $c_{ik\ell o}^{ship}>0$ be the per-unit shipping cost when using location $\ell$ and carrier $k$ to fulfill SKU $i$ in order $o$. For each location $\ell$, let $c_{\ell}^{fixed}>0$ be the per-unit fixed sourcing cost. For each carrier $k$, let $\beta_{k}\in(0,1)$ be the discount percentage applied to shipping costs when multiple units are shipped using carrier $k$ from the same fulfillment center. Let $\gamma^{+}\geq 0$ and $\gamma^{-}\geq 0$ be constants that convert late and early delivery penalties into per-unit costs, respectively.
In an ideal setting where all carrier-location deviations are known at the time of fulfillment, let the vector $\mathbf{d}_{o}=(d_{k\ell o})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ represent the realized uncertain delivery deviations for order $o$, where each $d_{k\ell o}\in\mathbb{R}$ denotes the number of days that the carrier-location pair $(k,\ell)$ deviates from desired delivery date. A positive deviation indicates late delivery, a zero deviation indicates on-time delivery, and a negative deviation indicates early delivery.
The weighted objective function for order $o$ can be constructed as follows:
$g(\mathbf{z}_{o},\mathbf{d}_{o})=\begin{cases}\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}[c_{\ell}^{fixed}+(1-\beta_{k})c_{ik\ell o}^{ship}+\gamma^{+}d_{k\ell o}^{+}+\gamma^{-}d_{k\ell o}^{-}]z_{ik\ell o}&\text{if }\sum_{i\in\mathcal{I}}z_{ik\ell o}\geq 2\\ \sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell o}^{ship}+\gamma^{+}d_{k\ell o}^{+}+\gamma^{-}d_{k\ell o}^{-})z_{ik\ell o}&\text{otherwise},\end{cases}$ (13)
where $d_{k\ell o}^{+}=\max\{0,d_{k\ell o}\}$ and $d_{k\ell o}^{-}=\max\{0,-d_{k\ell o}\}$. This objective function encourages item consolidation within an order by offering a carrier-specific discount when applicable. In addition, it leverages an asymmetric penalty cost function for delivery deviations, with $\gamma^{+}\gg\gamma^{-}$ to primarily discourage late deliveries while also disincentivizing early deliveries to a lesser extent. The objection function is piecewise linear and can be linearized using standard rewritings.
4.2.3 Operational Constraints
Several operational constraints ensure that the selection of fulfillment center locations and carriers is feasible in practice. For location $\ell\in\mathcal{L}$, let $\operatorname{inv}_{i\ell o}\geq 0$ be the inventory level of SKU $i$ at location $\ell$ and $\operatorname{cap}_{\ell o}\geq 0$ be the available capacity at location $\ell$ when order $o$ is placed. For SKU $i$ in order $o$, let $q_{io}\geq 0$ be the quantity of SKU $i$ in order $o$, and let $e_{ik\ell o}\in\{0,1\}$ be a binary indicator of whether the location-carrier pair $(\ell,k)$ is an eligible fulfillment option.
  • Each unit of an SKU must be sourced from exactly one location and shipped by exactly one carrier.
    $\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}z_{ik\ell o}=q_{io},\quad\forall i\in\mathcal{I}$ (14)
  • Each SKU can only be fulfilled by eligible candidate location-carrier pairs, which may vary depending on the carrier’s coverage and package requirements.
    $z_{ik\ell o}\leq q_{io}e_{ik\ell o},\quad\forall i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}$ (15)
  • Only locations with available inventory of the requested SKUs can be chosen.
    $\sum_{k\in\mathcal{K}}z_{ik\ell o}\leq\operatorname{inv}_{i\ell o},\quad\forall i\in\mathcal{I},\ell\in\mathcal{L}$ (16)
  • Each location has a daily capacity limit on the total number of units it can process.
    $\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}z_{ik\ell o}\leq\operatorname{cap}_{\ell o},\quad\forall\ell\in\mathcal{L}$ (17)
4.3 Applying the CSO Framework to the Problem
In practice, the delivery deviations of each carrier-location pair cannot be perfectly estimated in advance. To account for this uncertainty for each order $o$, this section models the vector of uncertain delivery deviations $\tilde{\mathbf{d}}_{o}=(\tilde{d}_{k\ell o})_{k\in\mathcal{K},\ell\in\mathcal{L}}\sim\mathbb{P}$ as a $(KL)$-dimensional random variable governed by the probability distribution $\mathbb{P}$. In addition, let $\mathbf{s}_{o}=(\mathbf{s}_{k\ell o})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ denote the state vector observable for order $o$ before the fulfillment decision is made, where each $\mathbf{s}_{k\ell o}$ is a vector of covariates encompassing information about the fulfillment network status and order details related to carrier-location pair $(k,\ell)$. The conditional distribution of the deviation vector, upon observing $\mathbf{s}_{o}$, is denoted by $\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})$, and the individual deviation condition distribution is $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell o})$ for each $k\in\mathcal{K}$ and $\ell\in\mathcal{L}$.
Applying the CSO framework leads to the following contextual stochastic order fulfillment problem (CSOFP):
(CSOFP) $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})]=\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\Pi_{k\in\mathcal{K},\ell\in\mathcal{L}}\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell o})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})].$ (18)
Effectively solving Problem (18) requires leveraging data from historical orders. As a specific instance of Problem (2), CSOFP also has an observational dataset. Let $\mathcal{O}$ be the set of historical orders. Let $\mathbf{z}_{o}=(z_{ik\ell o})_{i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}}$ be the vector of historical decisions made by the industrial partner’s sourcing engine for order $o$. There are numerous missing entries in the observed deviation vector $\mathbf{d}_{o}$ since actual observations depend on the decision vector $\mathbf{z}_{o}$. Let
$\mathbf{D}_{o}:=\{d_{k\ell o}\mid\sum_{i\in\mathcal{I}}z_{ik\ell o}>0,k\in\mathcal{K},\ell\in\mathcal{L}\}$
be the observable costs for order $o$. The dataset available to estimate $\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})$ is $\mathbb{D}=\{(\mathbf{s}_{o},\mathbf{D}_{o})\}_{o=1}^{|\mathcal{O}|}$.
$\mathbb{D}$ is partitioned into subsets to facilitate learning the individual condition distributions. For carrier $k$ and location $\ell$, the dataset for learning $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell o})$ is $\mathbb{D}_{k\ell}=\{(\mathbf{s}_{k\ell m},d_{k\ell m})\}_{m=1}^{M_{k\ell}}$, where $M_{k\ell}$ is the number of times the deviation has been observed for the carrier-location pair $(k,\ell)$ in $\mathbb{D}$.
4.4 Learning the Contextual Distribution of Delivery Deviations
This section describes how to learn Contextual Distribution of Delivery Deviations. It begins by selecting features to construct the covariate vector for each carrier-location pair. To predict the conditional distribution of deviations, two classes of machine learning methods are proposed: probabilistic multi-class classification, which leverages the discrete nature of deviations, and quantile regression, which is known for its effectiveness in distributional prediction (Salari et al. 2022). For simplicity, the index $k\ell$ is omitted in the notations used throughout this section. The data set can then be denoted by $\mathbb{D}=\{(\mathbf{s}_{m},d_{m})\}_{m\in[M]}$.
4.4.1 Contextual Features Selection
Carefully selecting contextual features is crucial for constructing informed covariates for delivery deviation predictions, as various stages in the order life cycle can contribute to deviations. To identify the most important features for predicting deviations, a feature importance analysis was conducted using a tree-based mean regression model. Two methods were examined: Mean Decrease in Impurity (MDI) and Shapley values (SHAP). MDI measures the importance of a feature by calculating the total reduction in impurity (e.g., Gini impurity or entropy) it brings when used in decision trees (Breiman 2001). SHAP values, on the other hand, provide a unified measure of feature importance based on cooperative game theory, attributing to each feature its contribution to the prediction by considering all possible feature combinations (Lundberg et al. 2020).
The features considered are categorized into three main classes, each corresponding to different stages of the order fulfillment process:
  1. 1.
    Order-level features: This category includes customer location, SKU weight and dimensions, SKU type, quantities, release hour, release day, and planned lead time (i.e., the difference between the promised delivery date and the order release time). These factors can influence deviations through logistical complexity, handling requirements, and time-sensitive elements affecting the order’s processing and delivery schedule.
  2. 2.
    Fulfillment center features: This includes coordinates, location type, on-hand capacity, and on-hand inventory. The geographical location and type of fulfillment center can affect the efficiency of order processing, while capacity and inventory levels determine the speed and reliability of fulfilling orders.
  3. 3.
    Carrier-related features: This includes the level of service, promised time in transit, shipping charges, and carrier zone. These factors impact the carrier ability to meet delivery deadlines, with service level and transit times directly affecting deviations, while shipping charges and zones reflect logistical challenges and potential bottlenecks.
4.4.2 Probabilistic Multi-Class Classification
Since deviations measured in days only take on discrete values in the dataset, they inherently follow a discrete probability distribution. Thus, the random variable $\tilde{d}$ has a finite discrete support. Let $C$ be the total number of possible values $\tilde{d}$ can take, and let $\xi_{c}$ be the $c$-th possible value for $c\in[C]$. By definition, $\tilde{d}\in\{\xi_{1},\xi_{2},\ldots,\xi_{C}\}$. This discrete nature allows the conditional distribution learning problem to be framed as a multi-class classification (MC-CLF) problem. Under MC-CLF, each class represents a possible delivery deviation value in number of days, and the output class probabilities correspond directly to the probability mass function of the discrete distribution. For a new sample with covariates vector $\mathbf{s}$, let the conditional probability of class $c$ be $p_{c}(\mathbf{s})=\mathbb{P}(d=\xi_{c}\mid\mathbf{s})$, with the corresponding predicted probability from MC-CLF denoted as $\hat{p}_{c}(\mathbf{s})$.
Multinomial logistic regression (MLR) and classification trees-based models are considered to solve MC-CLF. Specifically, MLR is a classic classifier that aims to estimate the class probabilities by maximizing the log-likelihood function, which is the negative of the log loss function. The log loss is defined as $-\sum_{m=1}^{M}\sum_{c=1}^{C}\mathbf{1}\{d_{m}=\xi_{c}\}\log(\hat{p}_{c}(\mathbf{s}_{m})).$ On the other hand, classification trees use a nonparametric approach that recursively partitions the data into subsets based on feature values in a top-down manner, resulting in a tree-like model of decisions.
Let $\mathcal{R}_{1},\ldots,\mathcal{R}_{R}$ be the regions partitioned by a classification tree. Given covariates $\mathbf{s}$, the predicted probability for class $c$ in a classification tree can be expressed as:
$\hat{p}_{c}(\mathbf{s})=\frac{\sum_{r=1}^{R}\mathbf{1}\{\mathbf{s}\in\mathcal{R}_{r}\}\cdot\mathbf{1}\{d=\xi_{c}\}}{\sum_{r=1}^{R}\mathbf{1}\{\mathbf{s}\in\mathcal{R}_{r}\}},$ (19)
To enhance predictive power, ensemble learning methods such as bagging and gradient boosting are employed, through the use of random forests (RF) and gradient boosted trees (GBT). The details for their calculation of the class probabilities are available in Appendix.
Moreover, while the proposed classification models excel at predicting class labels, their predicted probabilities may not be well calibrated (Niculescu-Mizil and Caruana 2005). Therefore, probability calibration is necessary to obtain accurate conditional distributional forecasts. To address this, a cross-validation approach is adopted to fit an isotonic regression calibrator (Zadrozny and Elkan 2002). In each split, the classification model is first trained on the training set, and then its predictions on the calibration set are used to fit an isotonic regression model. Let $\hat{\tilde{p}}(\cdot)$ be the vector of calibrated probabilities and $\hat{p}(\cdot)$ be the vector of uncalibrated probabilities for each class. Isotonic regression takes the vector of uncalibrated probabilities as input and solves the following problem to determine a vector of monotonically increasing calibrated probabilities:
$\min_{\hat{\tilde{p}}(\cdot)}$ $\sum_{m=1}^{M}(d_{m}-\hat{\tilde{p}}(\mathbf{s}_{m}))^{2}$ (20)
s.t. $\hat{\tilde{p}}(\mathbf{s}_{m})\geq\hat{\tilde{p}}(\mathbf{s}_{m^{\prime}})\quad\forall m^{\prime}\neq m:\hat{p}(\mathbf{s}_{m})\geq\hat{p}(\mathbf{s}_{m^{\prime}})$
To measure the accuracy of probabilistic predictions for MC-CLF model, the paper adopts the log loss and the Brier score. The latter is defined as the average squared error between predicted probabilities and actual outcomes $\sum_{m=1}^{M}\sum_{c=1}^{C}(\mathbf{1}\{d_{m}=\xi_{c}\}\hat{p}_{c}(\mathbf{s}_{m}))^{2}$.
For both metrics, lower values indicate better probabilistic prediction performance.
4.4.3 Quantile Regression
Unlike multi-class classification, quantile regression (QR) can directly estimate the cumulative distribution function (CDF) of the conditional distribution without assuming a discrete probability distribution. The CDF of the uncertain deviation $\tilde{d}$ is given by $F(u\mid\mathbf{s})=\mathbb{P}(d\leq u\mid\mathbf{s})$. Standard QR models aim at minimizing the pinball loss function, estimating different quantiles independently. To approximate the conditional distribution, these models typically calculate a series of conditional quantiles, forming a discretized approximation of the CDF. However, this approach can lead to the quantile crossing issue, where predicted quantiles do not maintain their proper order, resulting in an invalid distribution. Various methods exist to avoid quantile crossing, but tree-based quantile regression models have been shown to outperform standard QR models in distributional forecasting, particularly in delivery time performance settings (Salari et al. 2022). These tree-based methods do not suffer from the issue of quantile crossing. Therefore, this paper focuses on a regression tree-based model and a quantile random forest (QRF) model which leverages an ensemble of regression trees.
Similar to a classification tree, a regression tree partitions the feature space into multiple disjoint regions, but the split criterion is based on the mean squared error (MSE). While regression trees are typically used for generating conditional point predictions, they can be extended to provide distributional forecasts by preserving all observations within each partitioned region, rather than just summary statistics like the mean. Let $\mathcal{R}_{1},\ldots,\mathcal{R}_{R}$ be the regions partitioned by a regression tree. Given covariates $\mathbf{s}$, the region that $\mathbf{s}$ falls into is denoted as $\mathcal{R}(\mathbf{s})$. The conditional distribution of the new sample can be approximated by the empirical CDF:
$\hat{F}(u\mid\mathbf{s})=\sum_{m=1}^{M}\frac{\mathbf{1}\{\mathbf{s}_{m}\in\mathcal{R}(\mathbf{s})\}}{|\{j:\mathbf{s}_{j}\in\mathcal{R}(\mathbf{s})\}|}\mathbf{1}\{d_{m}\leq u\}.$ (21)
To refine this empirical CDF, a kernel density estimation method is applied to the observations in $\mathcal{R}(\mathbf{s})$ (Salari et al. 2022).
A similar approach can be extended to random forests, and this idea of QRF was first introduced by Meinshausen and Ridgeway (2006). In QRF, for a new sample with covariates $\mathbf{s}$, the empirical distribution is derived from all regions associated with $\mathbf{s}$ across all trees in the forest. Let $T$ be the total number of regression trees, and let $\mathcal{R}_{1}^{(t)},\ldots,\mathcal{R}_{R}^{(t)}$ be the partitioned regions of the $t$-th tree, where $t\in[T]$. The region that $\mathbf{s}$ falls into in the $t$-th tree is denoted as $\mathcal{R}^{(t)}(\mathbf{s})$. QRF approximates the conditional distribution using the empirical CDF:
$\hat{F}(u\mid\mathbf{s})=\frac{1}{T}\sum_{t=1}^{T}\sum_{m=1}^{M}\frac{\mathbf{1}\{\mathbf{s}_{m}\in\mathcal{R}^{(t)}(\mathbf{s})\}}{|\{j:\mathbf{s}_{j}\in\mathcal{R}^{(t)}(\mathbf{s})\}|}\mathbf{1}\{d_{m}\leq u\}.$ (22)
To measure the distributional forecasting performance of different learning models, the paper uses two metrics—the Continuous Ranked Probability Score (CRPS) and pinball loss. CRPS evaluates the closeness of the estimated CDF $\hat{F}$ given covariates $\mathbf{s}$ to the observed ground truth value $d$:
[…]
Let $\xi$ and $\xi^{\prime}$ be two independent random variables with distribution $\hat{F}(u\mid\mathbf{s})$. The CRPS can alternatively be expressed as:
$\text{CRPS}(\hat{F},d,\mathbf{s})=\mathbb{E}(|\xi-d|)-\frac{1}{2}\mathbb{E}(|\xi-\xi^{\prime}|).$ (24)
This equation allows for an approximation of CRPS by sampling from $\hat{F}(u\mid\mathbf{s})$. Pinball loss measures the accuracy of quantile predictions by penalizing errors asymmetrically, depending on whether the predictions are overestimates or underestimates. Let $q$ be a specific quantile. Then, the pinball loss between the observed value $d$ and the predicted value $\hat{d}$ is defined as:
$\text{pinball}(d,\hat{d})=q\max\{d-\hat{d},0\}+(1-q)\max\{\hat{d}-d,0\}.$ (25)
4.4.4 Applying the C-SAA and the C-RO
It is impractical to solve the CSOFP exactly with the learning models just proposed. Indeed, on the one hand, QR models learn continuous distributions, making it computationally intractable to expand the expectation directly. On the other hand, MC-CLF models leverage the fact that the uncertain delivery deviation vector $\tilde{\mathbf{d}}$ has a finite discrete support, which theoretically allows for expansion of the expectation term. However, MC-CLF models face the curse of dimensionality. The number of possible values of $\tilde{\mathbf{d}}$ grow exponentially in $K$ and $L$. This justifies the methodologies proposed in Section 3.
In the following, the C-SAA problems using samples from the conditional distribution estimates provided by the MC-CLF and QR models are referred to as C-SAA-CLF and C-SAA-QR, respectively. Similarly, the C-RO problems are referred to as C-RO-{B,D}-CLF and C-RO-{B,D}-QR, respectively.
5 Case Study
This section presents computational results that demonstrate the potential of the proposed CSO framework when evaluated on real-world industrial instances.
5.1 Data Overview
The industrial partner provides a month-long dataset in 2023, covering more than 1 million online orders for home delivery and over 20,000 unique stock-keeping units (SKUs). Each order is organized by SKUs, resulting in one or more order lines. For an incoming order, there are, on average, 250 to 1,000 possible fulfillment centers across the company’s network to source SKUs from, including mostly physical stores but also regional distribution centers and metro e-commerce centers. Each location is equipped with multiple eligible carriers that are capable of providing different levels of service (LOS), such as Ground, Same-Day, and Overnight shipping. The set of available carriers span from major shipping companies to crowdsourced gig carriers. There are 10 to 30 carrier-LOS pairs appeared in the dataset and throughout the paper, “carrier” refers to one of these specific carrier-LOS pairs.
Each row of the dataset contains transactional details for a single order line, along with the status of the fulfillment network at the time the order was placed. The order information includes the date and time of order placement, the type, dimensions, weight, and quantity of the SKUs, the fulfillment location from which each order line was sourced, the carrier that shipped the order, the customer location, the promised delivery date, and the actual arrival date. The network status data includes the on-hand inventory levels of the requested SKU and the remaining open capacity at various locations. Additionally, the dataset includes carrier-provided shipping costs and planned transit times.
5.1.1 Learning Results
For the evaluation of the contextual learning models, the dataset is split into two parts: orders from the first three weeks (1.3 million order lines) are used for training, while orders from the last week (0.4 million order lines) are reserved for testing and evaluation.
In Table 1, two metrics—log loss and Brier score—are reported to measure the accuracy of probabilistic predictions for the tuned classification models on the test set. Probability calibration is applied to the uncalibrated models that obtained the best metrics. It is notable that both the log loss and Brier scores show improvement after calibration. In addition, Figure 2 illustrates the performance of an example distributional forecast by multi-class classification against its true empirical distribution on the test set. The distribution of the majority class 0 (on-time deliveries) is predicted quite accurately, while the distributions of late and early deliveries are forecasted sufficiently close, albeit with small deviations.
Table 1: Average Log Loss and Brier Score on Test Set
Model MLR Classification Tree RF-CLF GBT-CLF Calibrated-CLF
Log loss 1.950 2.242 1.593 1.050 0.965
Brier score 0.073 0.072 0.054 0.034 0.031
Figure 2: Comparison of Calibrated Predicted Joint Distribution with True Empirical Distribution for an Example Carrier on Test Set.
Table 2 presents the pinball losses for different quantiles and the CRPS values of the tuned quantile regression models and classification models. The CRPS values are approximated by drawing 1,000 samples to compute Equation (24). It is observed that the MC-CLF models exhibit higher pinball losses compared to the two quantile regression models, likely due to the discrete nature of classification outputs. However, the CRPS of MC-CLF is the lowest among the three models, despite not being explicitly tuned for CRPS.
Additionally, Figure 3 displays the 95% prediction intervals created by the three conditional distribution learning methods for an example carrier. The visualization aligns with the pinball loss results, showing that QRF generally provides the narrowest interval widths. While the intervals generated by MC-CLF cover most of the observed values, their coverage is less effective at the two extremes compared to the two quantile regression methods. The inferior performance of MC-CLF at extreme quantiles is likely due to the discretization of the output space, which limits precision, and the cumulative probability method, which can cause jumps when selecting quantiles. In contrast, regression tree and QRF provide finer-grained predictions, leading to more accurate and narrower intervals at the extremes.
In summary, MC-CLF models excel at providing distributional forecasts, whereas QR models are better at generating prediction intervals for this dataset.
Table 2: Average Pinball Loss and CRPS on Test Set Across All Carriers.
Pinball loss CRPS
Model $q_{0.025}$ $q_{0.05}$ $q_{0.5}$ $q_{0.95}$ $q_{0.975}$
MC-CLF 0.199 0.198 0.676 0.526 0.304 0.386
Regression Tree 0.064 0.109 0.302 0.147 0.093 0.515
QRF 0.033 0.057 0.263 0.143 0.096 0.402
(a) Regression Tree Prediction Intervals
(b) MC-CLF Prediction Intervals
(c) QRF Prediction Intervals
Figure 3: Sorted Observed Deviation Values versus 95% Prediction Intervals for an Example Carrier.
5.2 Baseline Models
To compare with the proposed contextual optimization models, three baseline order fulfillment algorithms are introduced below.
  1. 1.
    Pure Cost-Driven Greedy Heuristic (Greedy): This greedy heuristic replicates the current practice employed by the industrial partner’s sourcing engine. The algorithm selects the location-carrier pair with the lowest fulfillment cost for each quantity of an SKU. For order $o$, the greedy problem is
    $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\{\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell o}^{ship})z_{ik\ell o}\}.$
  2. 2.
    Empirical-SAA: This method follows the approach described in Section 3.4 for C-SAA, with the key difference being that the deviation vector samples are drawn from the empirical distribution of deviations in the training set, without incorporating any contextual information. For each sample $n\in[N]$, let $\xi_{k\ell o}^{n}$ denote the $n$-th devaition value sampled from the empirical distribution $P^{E}_{k\ell}$ for carrier-location pair $(k,\ell)$. The corresponding deviation vector for sample $n$ is then expressed as $\bm{\xi}_{o}^{n}=(\xi_{k\ell o}^{n})_{k\in\mathcal{K},\ell\in\mathcal{L}}$. For order $o$, the empirical SAA problem is
    $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\{\frac{1}{N}\sum_{n=1}^{N}g(\mathbf{z}_{o},\bm{\xi}_{o}^{n})\}.$
  3. 3.
    Point Predict, Then Optimize (Point-Pred.-Opt.): Unlike the distributional forecasting models mentioned in Section 4.4, typical regression-based ML models aim at estimates only the conditional mean $\mathbb{E}[\tilde{d}_{k\ell}\mid\mathbf{s}_{k\ell o}]$. In this method, conditional mean regression models $h_{k\ell}$ are trained to map covariates to deviation for each $(k,\ell)\in\mathcal{K}\times\mathcal{L}$. This approach transforms CSOFP into a deterministic optimization problem by directly replacing the random parameter $\tilde{\mathbf{d}}_{o}$ with its point prediction $\hat{\mathbf{d}}_{o}=\{h_{k\ell}(\mathbf{s}_{k\ell o})\}_{k\in\mathcal{K},\ell\in\mathcal{L}}$ in the objective function of Problem (18). This method is effective only if the objective function $g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})$ is linear in $\tilde{\mathbf{d}}_{o}$, in which case estimating the conditional distribution simplifies to estimating the conditional mean. This is because, by linearity of expectation, $\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})]=g(\mathbf{z}_{o},\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})}[\tilde{\mathbf{d}}_{o}])$. However, this linearity does not hold in the proposed objective function due to the asymmetric penalty costs associated with the deviations. So a point prediction may be biased, potentially leading to suboptimal solutions. To facilitate the experiments, several different regression models are trained to minimize MSE, and the model with the smallest MSE in cross-validation is selected for each carrier. See Appendix for the average MSE of different regression models. For order $o$ and point prediction $\hat{\mathbf{d}}_{o}$, the optimization model is
    $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}g(\mathbf{z}_{o},\hat{\mathbf{d}}_{o}).$
5.3 The Experimental Setting
The optimization models in this study were implemented using Gurobi 11.0.0 and Python 3.9. All experiments were conducted on a server equipped with dual-socket Intel Xeon Gold 6226 CPUs, each featuring 24 cores running at 2.7 GHz. The regression and multi-class classification models were developed using the scikit-learn (Pedregosa et al. 2011), XGBoost (Chen and Guestrin 2016), and CatBoost libraries (Prokhorenkova et al. 2018). For quantile regression, the regression tree and quantile random forest models were implemented using the quantile-forest package (Johnson 2024).
For the following experiments, all ML models are trained on orders from the first three weeks, while 5,000 multi-item orders randomly sampled from the last week are used to evaluate the order fulfillment methods. During both training and inference, for multi-SKU orders, features are aggregated by retaining only the features of the SKU with the highest shipping charges and quantities.
One of the difficulties in this application is that the ground truth is not available for options that were not selected. To overcome this limitation, a simulation environment was built and generate the “ground truth” when an algorithm selects an option. The delivery deviations are drawn from a multinomial distribution whose probabilities were estimated by fitting a fine-tuned, calibrated multi-class classification model based on orders in the last week of the dataset. Obviously, this data is not available to the algorithm and only used by the simulator to replicate real-world conditions.
By default, a constant $\beta_{k}=0.5$ discount factor for item batching is considered since the exact values are not provided. Sensitivity analysis are perform on the impact of this factor. The default penalty for early delivery $\gamma^{-}$ is $0.1$ and that for late delivery $\gamma^{+}$ is $10$. The penalty can be understood as the opportunity cost (lost sale cost) for early (late) deliveries, and they are also analyzed with a range of different values. For methods relying on SAA, the default is to have $Q=20$ independent replications each with a sample size of $N_{1}=30$. The obtained solutions are then evaluated with a default sample size of $N_{2}=300$. The default sample size used for C-RO-D is $N=200$. As for C-RO-B, a default of $95\%$ interval is considered. Moreover, a scaling factor $\eta\in[0,1]$ (default of 0.5) determines the uncertainty budget, defined as $B=\eta(KL).$
5.4 Effectiveness of C-SAA
This section first compares the performance of the C-SAA method against the three baseline methods. Throughout the following sections, only normalized values of the metrics are reported due to the confidentiality of the data.
5.4.1 Overall Results
Table 3 compares the realized objective value of Problem (18) for C-SAA and the baseline methods. For C-SAA, only the metrics of the best prediction model are recorded. It can be observed that C-SAA consistently has the lowest objective value across different late delivery penalties. C-SAA maintains a lower objective value on average across all $\gamma^{+}$ levels (approximately 0.80), compared to Point-Pred.-Opt. (0.87) and Empirical-SAA (0.83), highlighting its stability and reliability.
A perfect balance between the fulfillment costs and delivery timeliness control needs to be tailored for the specific business needs. On the one hand, a lower fulfillment cost brings mores immediate profits. On the other hand, lower delivery timeliness can significantly increase customer satisfaction, thus retaining more loyal customers, benefiting the company in long-term operations. In CSOFP, the trade-off between these short-term and long-term profits is controlled by the key parameter $\gamma^{+}$, which determines the penalty of late deliveries in terms of costs. The trend indicates that C-SAA’s objective value decreases more sharply compared to other methods as $\gamma^{+}$ increases, making it particularly effective in scenarios where delivery timeliness is crucial. C-SAA outperforms Point-Pred.-Opt. by 2% (0.92 vs. 0.94) when $\gamma^{+}=2.5$ and by 14% (0.63 vs. 0.79) when $\gamma^{+}=40$, indicating that C-SAA’s advantage becomes more pronounced as the penalty increases. While the difference between C-SAA and Empirical SAA in terms of objective value is not huge, the subsequent section will further illustrate the performance gaps between these two methods, as Empirical SAA is more sensitive and could potentially perform worse with larger sample sizes.
Table 3: Simulated Realized Objective Values of C-SAA vs. the Three Baselines
Method $\gamma^{+}=2.5$ $\gamma^{+}=5$ $\gamma^{+}=10$ $\gamma^{+}=20$ $\gamma^{+}=40$
Greedy 1.0 (-) 1.0 (-) 1.0 (-) 1.0 (-) 1.0 (-)
Point-Pred.-Opt. 0.94 $(2.6\times 10^{-4})$ 0.92 $(3.2\times 10^{-3})$ 0.88 $(4.2\times 10^{-3})$ 0.83 $(5.5\times 10^{-3})$ 0.79 $(7.0\times 10^{-3})$
Empirical-SAA 0.93 $(0.9\times 10^{-4})$ 0.91 $(2.0\times 10^{-3})$ 0.86 $(3.1\times 10^{-3})$ 0.76 $(1.0\times 10^{-2})$ 0.64 $(1.2\times 10^{-2})$
C-SAA 0.92 $(1.3\times 10^{-4})$ 0.89 $(3.3\times 10^{-3})$ 0.82 $(5.5\times 10^{-3})$ 0.72 $(1.1\times 10^{-2})$ 0.63 $(1.3\times 10^{-2})$
  • Note: Values represent the normalized mean (standard error) across five different random seeds for the realized deviations simulation. The normalization is performed by setting the value for the Greedy method to 1.
5.4.2 Value of Contextual Information
To facilitate a closer examination of the performance of these SAA-based models, Figure 4 shows the impact of varying sample sizes. On the one hand, larger sample sizes should provide a more accurate approximation of the expected value in Problem (18), assuming the delivery deviation distributions are well-estimated. On the other hand, it is important to maintain a reasonably small sample size to ensure computational efficiency. For both C-SAA methods, the effect of increasing sample sizes is obvious. The objective values are reduced from 0.86 to 0.82 for C-SAA-CLF and 0.87 to 0.84 for C-SAA-QR on average, indicating their increasing alignment with the true risk-neutral objective function. This improvement is particularly evident in their enhanced effectiveness at avoiding cumulative lateness, defined to be the total number of late delivery days across all units in the orders. The cumulative lateness decreases by 21% for C-SAA-CLF and by 18% for C-SAA-QR.
In contrast, Empirical-SAA struggles to improve its performance in both the realized objective value and the cumulative lateness as the sample sizes grows beyond 10. This can be attributed to the fact that purely using the empirical distribution is not sufficient to capture the varying delivery deviation distributions that may be dependent upon contextual information unique upon each order. A biased distributional forecasts thus guide the model wrongly to create solutions with inferior quality. Based on these observations, both conditional distributional forecasting methods proposed in the paper are successfully in this application, demonstrating a stronger performance compared to using empirical distribution for delivery deviation forecasting. In summary, the value of contextual information is particularly important in order fulfillment decision-making, where the uncertainty of delivery deviation can vary drastically across different states of the environment.
(a) Performance in Realized Objective Value.
(b) Performance in Cumulative Lateness.
Figure 4: C-SAA vs. Empirical SAA Across Different Sample Sizes.
5.4.3 Value of Distributional Forecasts
This section examines the advantage of employing distributional forecasting models over the simpler and more common point prediction model for CSOFP. As supported by theory, simply using the conditional mean estimate of deviations to replace the expected value in CSOFP can introduce bias when approximating the true objective value. This approach may yield worse solutions compared to employing conditional distributional estimators. These performance differences become even more pronounced as late penalties increase, as illustrated in Figure 5. With higher late penalties, accurate forecasting of deviations becomes more crucial; the results reveal the inherent bias of Point-Pred.-Opt, leading to lower quality solutions. This is not only evident in the realized objective value (Figure 5(a)) but also in the cumulative lateness (Figure 5(b)). For example, the cumulative lateness for Point-Pred.-Opt. only improves by 7% from $\gamma^{+}=2.5$ to $\gamma^{+}=40$, whereas C-SAA-CLF and C-SAA-QR show significantly greater improvements of approximately 35% and 32%, respectively. This demonstrates that the C-SAA methods not only maintain lower cumulative lateness overall but also improve more significantly as the penalty for late deliveries increases, further highlighting the benefits of distributional forecasting.
(a) Performance in Realized Objective Value.
(b) Performance in Cumulative Lateness.
Figure 5: C-SAA vs. Point-Pred.-Opt. Across Different Late Delivery Penalties.
5.5 Evaluation of C-RO
This section evaluates the performance of the C-RO methods, whose results are shown in Figure 6. The C-RO methods generally provide more conservative solutions than C-SAA, with C-RO-B showing a performance closer to C-SAA, as shown by the realized objective values in Figure 6(a). Notably, C-RO-B outperforms C-RO-D by offering less conservative yet still robust solutions, while not requiring full distributional information. In real-time fulfillment scenarios, where sampling can be time-prohibitive, especially during peak hours, C-RO-B offers a quick and effective alternative, delivering reasonably high-quality solutions. Additionally, the effectiveness of C-RO methods in controlling late deliveries remains consistent as the penalty increases (see Figure 6(b)). This stability occurs because C-RO methods already account for worst-case scenarios, making them less sensitive to changes in the late delivery penalty, provided that the samples or uncertainty sets encompass these scenarios. As a result, C-RO methods outperform C-SAA in reducing cumulative lateness when the late delivery penalty is low but are outperformed by C-SAA when $\gamma^{+}>10$. However, the C-RO methods incur higher fulfillment costs—calculated as the total shipping costs and fixed non-parcel costs, minus item consolidation discounts across all orders—by up to 20% on average (see Figure 6(c)) This increase is driven by their risk-averse approach, which prioritizes avoiding extreme deviations by opting for more expensive fulfillment options.
(a) Performance in Realized Objective Value.
(b) Performance in Cumulative Lateness.
(c) Performance in Fulfillment costs.
Figure 6: Performance of C-RO Methods.
5.5.1 Flexibility of C-RO-B
The effectiveness of the C-RO models in controlling delivery deviations is evident since they can hedge against the worst-case deviation and potential misspecification of the prediction models. However, their robustness to extreme delivery deviations comes at the cost of more conservative solutions, which can result in higher fulfillment expenses. Fortunately, the C-RO-B models, beyond their distribution-free property, are equipped with the flexibility to control their level of conservativeness. Specifically, they are determined via the coverage of the prediction intervals and the robust uncertainty budget. Smaller prediction interval coverage and uncertainty budget may lead to less conservative solutions, making the models more beneficial in terms of addressing forecasting misspecifications rather than purely managing risk. Conversely, larger coverage and budget aim at minimizing the impact of worst deviations as much as possible with the sacrifice of overall solution quality.
The impact of different prediction interval coverages and robust uncertainty budgets on the quality of solutions obtained from C-RO-B is examined in Figures 7 and 8, respectively. As the prediction interval coverage increases from 50% to 95% in Figure 7(a) , both C-RO-B-CLF and C-RO-B-QR show a noticeable decline in the realized objective value, with the lowest values around 0.85 for C-RO-B-CLF and 0.86 for C-RO-B-QR at 95% and 90% coverage, respectively. This represents a reduction of approximately 6.5% for both C-RO-B methods compared to the objective values at 50% coverage. This suggests that increasing the coverage interval initially benefits the overall performance, up to a point. This is consistent with the general decline trend in cumulative lateness as prediction interval increases (see Figure 7(b)). However, at 99% coverage, there is a sharp increase in the objective value, This trend highlights that while conservative prediction intervals can reduce lateness, they also introduce higher fulfillment costs due to more conservative decisions, thus increasing the overall objective value. For C-RO-B-CLF, this increase in objective value is also due to a declined prediction performance with the largest prediction interval, as indicated by the final increase of cumulative lateness at 99% coverage.
(a) Performance in Realized Objective Value.
(b) Performance in Cumulative Lateness.
Figure 7: Performance of C-RO-B Across Different Prediction Interval Coverages.
A similar pattern is observed for varying robust budgets. As $\eta$ increases, the robust budget becomes larger, leading to more conservative solutions. For both C-RO-B methods, larger budget results in a more significant decrease in cumulative lateness (see Figure 8(b)), but the results are less clear when it comes to the objective value (see Figure 8(a)). The realized objective value decreases sharply as the robust budget scaling factor $\eta$ increases from 0.0 to 0.01 for both C-RO-B-CLF and C-RO-B-QR. However, after $\eta=0.01$, there is a noticeable jump in the objective value for C-RO-B-QR, while C-RO-B-CLF continues to stabilize. This jump suggests that C-RO-B-QR adopts more conservative decisions with higher robust budgets, which leads to an increase in fulfillment costs. Therefore, C-RO-B-QR hits its lowest objective value at $\eta=0.01$ whereas C-RO-B-CLF is able to obtain its minimum objective value at $\eta=1$. The analysis reveals a trade-off between increasing robustness and striking a balance between lower fulfillment costs and mitigating extreme deviation risks.
(a) Performance in Realized Objective Value.
(b) Performance in Cumulative Lateness.
Figure 8: Performance of C-RO-B Across Different Robust Uncertainty Budgets.
5.6 Impact of Consolidation Factor
This section provides a discussion on the impact of item consolidation. CSOFP relies on the coefficients $\beta_{k}$ to control the incentive to batch items. The industrial partner does not know the exact percentage of discount to be offered for item consolidation, which may depend on a variety of convoluted factors. Therefore, this section performs sensitivity analysis on a range of possible discount factors for the CSO models. The consolidation level, calculated as the percentage of units that are batched (i.e., sourced and shipped by the same location-carrier pair) in all orders, is introduced to measure the effectiveness of the methods in identifying item batching opportunities for cost discounts within each order.
Figure 9 illustrates the performance of C-SAA, C-RO-B, and C-RO-D under a series of different discount factors. As shown in Figure 9(a), C-SAA still consistently exhibits the lowest objective value across all discount factors, with C-RO-B comes next. As expected, all three methods exhibit lower objective values due to lower fulfillment costs as the discount factor increases (see Figure 9(b)). This aligns with the intuition that higher discounts encourage more cost-efficient fulfillment strategies.
The decrease in both C-RO methods can be explained by their higher consolidation levels as the discount factors increases. C-RO methods have higher fulfillment costs, so they benefit more from discounts and capitalize on this opportunity.
The trend in consolidation level for C-SAA is less straightforward. While it peaks at a discount factor of 0.25, the consolidation level remains lower than the baseline throughout. This suggests that, despite some consolidation opportunities, the C-SAA model does not exploit them to the same extent as the C-RO models. One possible explanation is that the benefits of consolidation in C-SAA might not be sufficient to outweigh the penalties associated with deviations. The peak at a discount factor of 0.25 could be an anomaly or indicative of specific conditions where consolidation was briefly more favorable, but overall, the model prioritizes minimizing deviation penalties over achieving higher consolidation.
(a) Performance in Realized Objective Value.
(b) Performance in Fulfillment Costs.
(c) Performance in Consolidation Level.
Figure 9: Performance of CSO Models Across Different Discount Factors.
6 Conclusion
This paper presented a data-driven contextual stochastic optimization framework, which was applied to an industrial omnichannel order fulfillment problem within the context of multi-courier logistics. Leveraging contextual optimization effectively addresses the inherent complexities of modern retail logistics, providing innovative solutions tailored to the dynamic e-commerce landscape. In particular, the proposed framework is capable of handling the availabilities of observational data only.
To achieve these results, the paper uses probabilistic multi-class classification and quantile regression models to estimate the conditional distribution of delivery time deviations. These machine learning models are then seamlessly integrated with SAA-based risk-neutral optimization models and robust optimization models under the proposed CSO framework. The paper evaluated the proposed methodology on a real dataset from one of the largest e-commerce company. The results, which significantly improve the baseline solution, showcase the ability of the CSO framework to improve service quality, customer satisfaction, and operational efficiency. They offer valuable insights and tools for practitioners aiming to optimize their order fulfillment strategies in the evolving e-commerce environment.
There are several potential directions for future work. The contextual optimization models could be extended to incorporate long-term inventory planning through multi-period order fulfillment optimization. This would shift the approach from a rather myopic perspective to a more comprehensive strategy that accounts for future inventory needs. The distributional forecasts of deviations could be improved by increasing the dataset to include factors such as seasonal trends, weather conditions, and carrier tracking details. Furthermore, integrating downstream optimization costs into the training of distributional forecasting models can be an interesting extension (Qi et al. 2021, Elmachtoub and Grigas 2022). However, it is crucial to address the observational nature of the dataset and the varying dimensions of different optimization problem instances. Another promising direction is to extend the current contextual optimization framework to an end-to-end optimization approach (Kotary et al. 2021) while adhering to all operational constraints .
This paper is relevant to the broader literature on uncertainty quantification, where conformal predictions offer a promising direction. Conformal predictions provide coverage guarantees for prediction intervals and could be particularly effective when combined with quantile regression (Romano et al. 2019). Moreover, adopting conformally calibrated predictive distributions (Manokhin 2017) and improving probability calibration methods could enhance model performance. The paper hints at this potential by adopting probability calibration in multi-class classification. The Venn-Abers calibrator was experimented, but it was too computationally intensive to be fitted for the dataset used in the paper.
[…]
This research was partly supported by the NSF AI Institute for Advances in Optimization (Award 2112533).

### v2

@@
 […]
-Problem definition: The paper studies a large-scale order fulfillment problem for a leading e-commerce company in the United States.
-The challenge involves selecting fulfillment centers and shipping carriers with observational data only to efficiently process orders from a vast network of physical stores and warehouses.
-The company’s current practice relies on heuristic rules that choose the cheapest fulfillment and shipping options for each unit, without considering opportunities for batching items or the reliability of carriers in meeting expected delivery dates.
-Methodology / results: The paper develops a data-driven Contextual Stochastic Optimization (CSO) framework that integrates distributional forecasts of delivery time deviations with stochastic and robust order fulfillment optimization models.
-The framework optimizes the selection of fulfillment centers and carriers, accounting for item consolidation and delivery time uncertainty.
-Validated on a real-world data set containing tens of thousands of products, each with hundreds of fulfillment options, the proposed CSO framework significantly enhances the accuracy of meeting customer-expected delivery dates compared to current practices.
-It provides a flexible balance between reducing fulfillment costs and managing delivery time deviation risks, emphasizing the importance of contextual information and distributional forecasts in order fulfillment.
-Managerial implications: This is the first paper that studies the omnichannel multi-courier order fulfillment problem with delivery time uncertainty through the lens of contextual optimization, fusing machine learning and optimization.
-The results provide actionable insights for e-commerce companies and online retailers to enhance service quality and customer satisfaction through efficient order fulfillment strategies that accounts for delivery time uncertainty.
-Adopting this framework can balance cost effectiveness with customer service, potentially leading to higher customer retention and profitability.
 […]
-logistics, order fulfillment, machine learning, contextual stochastic optimization, contextual robust optimization
 1 Introduction
@@
 Additionally, customers can have mixed sentiments over underpromised delivery times, as early deliveries may not align with their availability.
-Therefore, missing the exact promised delivery date can negatively impact short-term and long-term sales, both implicitly and explicitly (Chan et al.
-2018, Cui et al.
-2020, Salari et al.
-2022).
-Given the crucial role of delivery timeliness, the paper incorporates delivery time uncertainty into the omnichannel multi-courier order fulfillment optimization problem through a contextual stochastic optimization (CSO) framework.
-One of the key challenges of Omnichannel Multi-Courier Order Fulfillment Optimization comes from the nature of historical data: only observational data is available, i.e., only the delivery time deviations of the selected options are available in the data set.
-In addition, in real time, these deviations are only observable once the decisions have been taken.
-The availability of observation data only raises fundamental challenges that are typically not addressed in prior work.
+Therefore, missing the exact promised delivery date can negatively impact short-term and long-term sales, both implicitly and explicitly (Salari et al.
+2022, Cui et al.
+2024).
+Given the crucial role of delivery timeliness, the paper incorporates delivery time uncertainty into the omnichannel multi-courier order fulfillment optimization problem through a contextual stochastic optimization framework.
+One of the key challenges of Omnichannel Multi-Courier Order Fulfillment Optimization comes from the nature of historical data: only observational data—i.e., delivery time deviations for the fulfillment options actually chosen—are available.
+Counterfactual deviations for unchosen options are never observed, similar to the bandit feedback setting (Lattimore and Szepesvári 2020).
+The availability of observational data only, raises fundamental challenges that are typically not addressed in prior work.
 To address these challenges, this paper proposes a Contextual Stochastic Optimization (CSO) framework, encompassing both risk-neutral and robust approaches.
@@
 The key novelty in this paper is a CSO framework for an omnichannel multi-courier order fulfillment optimization with observational data only.
-Existing studies on CSO typically assume the availability of a dataset $\mathcal{D}=\{(\mathbf{s}_{1},\mathbf{c}_{1}),(\mathbf{s}_{2},\mathbf{c}_{2}),\ldots\}$, pairing observed covariates with complete uncertain parameter vectors.
+Existing studies on CSO typically assume the availability of a dataset $\mathcal{D}=\{(\mathbf{s}_{1},\mathbf{c}_{1}),(\mathbf{s}_{2},\mathbf{c}_{2}),\ldots\}$, pairing covariates with fully observed uncertain parameter vectors.
 However, real-world scenarios often involve high-dimensional uncertain parameters where individual data points capture only partial information, particularly when uncertainty measurement depends on the chosen solution, which is the case in the application considered in this paper.
@@
   1. 1.
-    To the best of the authors’ knowledge, this paper is the first study of the order fulfillment problem in the context of omnichannel and multi-courier logistics through the lens of data-driven contextual stochastic optimization.
+    Generic CSO framework for observational data: The paper presents a novel, distribution‑agnostic contextual stochastic optimization (CSO) framework that converts problems with missing‑counterfactual observational data into tractable problems.
+A contextual distribution oracle is learned from partial observations, and the resulting problem is solved via two scalable solution paradigms—Contextual Sample Average Approximation (C‑SAA) and Contextual Robust Optimization (C‑RO).
   2. 2.
-    A mixed-integer linear program (MILP) is developed for a novel order fulfillment optimization problem that jointly considers the locations of fulfillment centers and the selection of shipping carriers for online order fulfillment.
-This MILP accounts for the consolidation of items within each order.
+    Tailored contextual distribution learning for e-commerce delivery time deviations: The paper adapts calibrated probabilistic multi‐class classification and tree‐based quantile regression to exploit the discrete, ordinal nature of delivery time deviations, yielding well‑calibrated contextual distributions for the CSO framework.
   3. 3.
-    Two classes of machine learning models are proposed to learn the delivery time deviations: probabilistic multi-class classification models and quantile regression models that provide distributional forecasts of delivery time deviations.
+    Consolidation‑ and timeliness‑aware MILP for omnichannel, multi‑courier fulfillment: The paper proposes the first data‑driven mixed‑integer linear program (MILP) that jointly selects fulfillment centers and carriers and captures item‑consolidation discounts while embedding learned contextual deviation distributions in a tractable way.
   4. 4.
-    The proposed MILP and machine learning models are integrated into a generalized CSO framework.
-The framework integrates the conditional distributions of delivery time deviations with stochastic and robust optimization models via sample average approximation (SAA) and tractable reformulations.
-  5. 5.
-    The paper applies the CSO framework to a large-scale real-world instance, demonstrating its impact on balancing reduced fulfillment costs and increased on-time order deliveries.
-The results validate the practical effectiveness of contextual optimization in an e-commerce setting, bringing significant benefits compared to the baselines.
+    Industry‐scale real‐world validation: The paper includes an extensive case study at a major e‐commerce retailer to (1) validate the efficacy of the proposed approaches in reducing expected fulfillment costs and (2) show significantly improvements in on‐time delivery rates compared to standard heuristics and baseline models.
+These results demonstrate the potential for substantial real‑world impact.
 1.2 Structure of the Paper
-The remaining sections of the paper are organized as follows.
-Section 2 reviews the related literature.
-Section 3 formalizes the generic CSO problem addressed in this paper and proposes SAA and robust optimization methods.
-The methods are applied to an unique order fulfillment problem in Section 4, which describes the MILP model and the machine learning models.
-Section 5 describes the case study, provides experimental results, and discusses actionable insights obtained from the experiments based on a real dataset.
-Section 6 concludes the paper.
+The remainder of this paper is organized as follows.
+Section 2 reviews related literature.
+Section 3 formulates the omnichannel multi‑courier order‑fulfillment problem under delivery‑time deviation uncertainty.
+Section 4 develops a generic contextual stochastic optimization framework for observational data and introduces two tractable solution paradigms—Contextual Sample Average Approximation and Contextual Robust Optimization.
+Section 5 describes the machine‑learning methods (multi‑class classification and tree‑based quantile regression) used to implement the contextual distribution oracle.
+Section 6 presents the case study, computational results, and actionable insights from a real dataset.
+Finally, Section 7 concludes and outlines directions for future research.
 2 Related Literature
@@
 (2023).
-Among the body of research on CSO, the most relevant to this paper are those that focus on learning conditional distributions of uncertain parameters and incorporating these forecasts into optimization models.
+Among the body of research on CSO, the most relevant to this paper are those that focus on learning contextual distributions of uncertain parameters and incorporating these forecasts into optimization models.
 Data-driven approaches typically build these forecasts from empirical residual errors (Deng and Sen 2022, Kannan et al.
@@
 Recent research has also explored integrating downstream optimization effects into ML model training.
-Elmachtoub and Grigas (2022) introduce the Smart Predict then Optimize (SPO) loss and its convex surrogate for point prediction models, with subsequent extensions to conditional distributional estimation (Qi et al.
+Elmachtoub and Grigas (2022) introduce the Smart Predict then Optimize (SPO) loss and its convex surrogate for point prediction models, with subsequent extensions to contextual distributional estimation (Qi et al.
 2021, Kallus and Mao 2023).
@@
 (2019) and Jo et al.
-(2021) using optimal prescriptive trees for prescriptive analytics in personalized treatment.
-While these methods effectively learn policies mapping features to actions, they cannot be directly applied to this fulfillment problem considered in this paper due to the presence of instance-specific hard constraints on decisions.
+(2021), using optimal prescriptive trees for prescriptive analytics in personalized treatment.
+While these methods effectively learn policies mapping features to actions, they cannot be directly applied to the fulfillment problem considered in this paper due to the presence of instance-specific hard constraints on decisions.
 The proposed framework addresses these limitations, offering a novel approach to CSO problems with observational data and constrained decision spaces.
@@
 (2022) focus on a single-carrier scenario.
-3 Problem and Methodology
-This section describes a generic contextual stochastic optimization problem (denoted by ${\cal P}$) for fulfillment applications where only an observational dataset is available.
-It then outlines the overall approach taken to address the problem.
-3.1 Problem Description
-Problem ${\cal P}$ consists in fulfilling a set of incoming requests sequentially, each of which contains orders for a set of products.
-Each request $o$ has an associated quantity $q_{io}$ for some product $i\in\mathcal{I}$.
-These products can be fulfilled from a set $\mathcal{V}$ of selection options.
-When a request $o$ arrives and $o$ contains a subset of products in $\mathcal{I}$, the goal is to determine the quantity of product $i$ assigned to selection option $v$.
-The decision variables can thus be captured by a vector $\mathbf{z}_{o}=(z_{ivo})_{i\in\mathcal{I},v\in\mathcal{V}}$, where each $z_{ivo}$ indicates the quantity of product $i$ assigned to selection option $v$ in request $o$.
-Each such assignment must satisfy operational constraints and the set of feasible assignments is denoted by $\mathcal{Z}_{o}$.
-For each product $i$, each selection option $v$ is associated with an uncertain cost $\tilde{c}_{ivo}\in\mathbb{R}$.
-The full vector of uncertain costs, governed by a probability distribution $\mathbb{P}$, is denoted by $\mathbf{\tilde{c}}_{o}=(\tilde{c}_{ivo})_{i\in\mathcal{I},v\in\mathcal{V}}$.
-The objective function $g(\mathbf{z}_{o},\mathbf{\tilde{c}_{o}})$ of the problem depends on these uncertain costs and the decision variables.
-Problem ${\cal P}$ is thus a risk-neutral stochastic optimization which can be formulated as follows for request $o$:
-$\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\mathbf{\tilde{c}}_{o}\sim\mathbb{P}}[g(\mathbf{z}_{o},\mathbf{\tilde{c}}_{o})],$ (1)
-where $\mathbb{E}_{\mathbf{\tilde{c}}_{o}\sim\mathbb{P}}[\cdot]$ represents the expectation of the objective function over the probability distribution $\mathbb{P}$.
-3.2 Contextual Stochastic Optimization (CSO)
-In Problem (1), the distribution $\mathbb{P}$ is not readily available.
-However, side information is typically available for product $i$ and selection option $v$.
-Let the vector of covariates for request be $\mathbf{s}_{ivo}$.
-In the context of an omnichannel order fulfillment problem, these covariates may represent the state of the fulfillment network, the weather, and the current demand for products.
-All these covariate vectors are aggregated into state vector $\mathbf{s}_{o}=(\mathbf{s}_{ivo})_{i\in\mathcal{I},v\in\mathcal{V}}$, which encompasses all contextual information available to the request at the time of the assignment decision.
-The distribution $\mathbb{P}$ is then approximated by the conditional distribution $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ of the cost vector given the observed state vector.
-This leads to the following Contextual Stochastic Optimization (CSO) problem:
-$\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\tilde{\mathbf{c}}_{o}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}[g(\mathbf{z}_{o},\mathbf{\tilde{c}}_{o})],$ (2)
-where $\mathbb{E}_{\tilde{\mathbf{c}}_{o}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}[\cdot]$ is the conditional expectation of the objective function given the state vector $\mathbf{s}_{o}$.
-Let the uncertain cost vector and the aggregated covariates for selection option $v$ be $\tilde{\mathbf{c}}_{vo}=(\tilde{c}_{ivo})_{i\in\mathcal{I}}$ and $\mathbf{s}_{vo}=(\mathbf{s}_{ivo})_{i\in\mathcal{I}}$, respectively.
-This paper assumes that option costs are independent, i.e.,
-$\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})=\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo}),$
-where $\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo})$ is the conditional distribution of the uncertain cost $\tilde{\mathbf{c}}_{vo}$ given the covariates $\mathbf{s}_{vo}$.
-This assumption is reasonable as the uncertain cost for a specific option depends solely on the contextual features of that option, not on the contextual features of other options.
-Note that, although these conditional distributions are considered independent, selection options that share common covariates (e.g., weather) and thus their distributions will tend to be similar.
-Problem (2) can then be rewritten as
-$\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\tilde{\mathbf{c}}_{o}\sim\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo})}[g(\mathbf{z}_{o},\mathbf{\tilde{c}}_{o})].$ (3)
-3.3 Learning the Contextual Distribution
-The conditional distribution $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ is also generally unknown but, in practice, it can often be learned from historical data available.
-In many real-world settings however, each data point is observational, i.e., only those costs involved in the actual decision can be observed.
-As a result, the dataset available to estimate $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ is of the form $\mathbb{D}=\{(\mathbf{s}_{o},\mathbf{C}_{o})\}_{o=1}^{|\mathbb{D}|}$, where $\mathbf{C}_{o}$ are the observable costs for order $o$.
-The learning task consists in finding a machine learning $\mathcal{M}$ that approximates $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ in a supervised manner using the dataset $\mathbb{D}$: $\mathcal{M}$ receives a context $\mathbf{s}$ as input and returns a probability distribution $\mathcal{M}(\mathbf{s})$ of costs.
-Since learning $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ amounts to learning $\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo})$ for each option $v$, the dataset $\mathbb{D}$ is partitioned into subsets, with $\mathbb{D}_{v}=\{(\mathbf{s}_{v1},\mathbf{c}_{v1}),(\mathbf{s}_{v2},\mathbf{c}_{v2}),\ldots\}$ representing the dataset for option $v$.
-A machine learning model $\mathcal{M}_{v}$ is associated with each option $v$ and $\mathcal{M}(\mathbf{\mathbf{s}})=(\mathcal{M}_{1}(\mathbf{s_{1}}),\ldots,\mathcal{M}_{|\mathcal{V}|}(\mathbf{s}_{|\mathcal{V}|}))$.
-3.4 Contextual Sample Average Approximation (C-SAA)
-Problem (3) is approximated by generalizing the Sample Average Approximation (SAA) method Kleywegt et al.
-(2002) to contextual stochastic optimization.
-The core of the Contextual Sample Average Approximation (C-SAA) for a request $o$ with covariates $\mathbf{s}_{o}$ amounts to
-  1. 1.
-    sampling the distribution $\mathcal{M}(\mathbf{s}_{o})$ $N$ times to obtain a set $\bm{\xi}_{o}^{n}=(\bm{\xi}_{vo}^{n})_{v\in\mathcal{V}}$ of cost vectors $(n\in[N])$;
-  2. 2.
-    solving the C-SAA problem
-    (C-SAA) $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\frac{1}{N}\sum_{n=1}^{N}g(\mathbf{z}_{o},\bm{\xi}_{o}^{n}).$ (4)
-Following Kleywegt et al. (2002), the C-SAA procedure can be outlined as follows:
-  1. 1.
-    Solve $Q$ independent C-SAA problems, each of sample size $N=N_{1}$, to produce candidate solutions $\mathbf{z}_{1},\ldots,\mathbf{z}_{Q}$ with corresponding objective values $\underaccent{\bar}{f}_{1},\ldots\underaccent{\bar}{f}_{Q}$.
-Calculate the average objective value as $\underaccent{\bar}{f}=\frac{1}{Q}\sum_{i=1}^{Q}\underaccent{\bar}{f}_{i}$.
-Let $f^{*}$ be the true optimal objective value of Problem (2).
-Since $\underaccent{\bar}{f}$ is an unbiased estimator of $\mathbb{E}_{{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}}[\underaccent{\bar}{f}]$, and because $\mathbb{E}_{{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}}[\underaccent{\bar}{f}]\leq f^{*}$, $\underaccent{\bar}{f}$ serves as a statistical lower bound on $f^{*}$ (Kleywegt et al.
-2002).
-  2. 2.
-    Then, re-evaluate the solutions $\mathbf{z}_{1},\ldots,\mathbf{z}_{Q}$ using a larger sample size $N=N_{2}\gg N_{1}$, yielding new objective values $\bar{f}(\mathbf{z}_{1}),\ldots,\bar{f}(\mathbf{z}_{Q})$.
-Since each candidate solution is feasible, its re-estimated objective values $\bar{f}(\cdot)$ provides an upper bound on $f^{*}$.
-Therefore, the approximated upper bound on the optimality gap for each candidate solution is $\bar{f}(\cdot)-\underaccent{\bar}{f}$.
-Finally, select the best solution $\mathbf{z}^{*}$ from the candidates, which is the one with the smallest optimality gap.
-It is interesting to compare C-SAA with the approach proposed in Bertsimas and Kallus (2020).
-A key idea in Bertsimas and Kallus (2020) is to use the historial data (the dataset $\mathbb{D}$ in this paper, as the samples for the stochastic optimization and to weight them properly.
-Unfortunately, this is not possible in this problem $\mathcal{P}$ considered here, because only observational data is available, i.e., only the costs of the selected options are revealed.
-One way to proceed would be to define the sample set for the cost vector as all possible combinations of data points across the selection options.
-However, as the number of selection options and historical data points increase, this approach suffers from the curse of dimensionality and becomes intractable for large dataset with big selection space, which is the case in the application considered in this paper.
-3.5 Contextual Robust Optimization (C-RO)
-The performance of C-SAA largely hinges on the accuracy of the estimated conditional distribution.
-To mitigate the risk of model misspecification in forecasting, robust optimization can be employed to account for worst-case scenarios, making it a viable approach for contextual stochastic optimization.
-Let $\mathcal{U}(\mathbf{s}_{o})$ represent the contextual uncertainty set for the cost vector $\tilde{\mathbf{c}}_{o}$.
-The Contextual Robust Optimization (C-RO) problem for (3) is then formulated as follows:
-(C-RO) $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\max_{\bm{\xi}_{o}\in\mathcal{U}(\mathbf{s}_{o})}g(\mathbf{z}_{o},\bm{\xi}_{o}).$ (5)
-To effectively leverage the conditional distributional forecasts and create tractable reformulations of Problem (5), two types of uncertainty sets are considered: a discrete uncertainty set $\mathcal{U}_{d}(\mathbf{s}_{o})$ and a budgeted interval uncertainty set $\mathcal{U}_{b}(\mathbf{s}_{o})$.
-3.5.1 Contextual Discrete Uncertainty Set
-With a discrete uncertainty set $\mathcal{U}_{d}(\mathbf{s}_{o})$, Problem (5) can be reformulated into a single-level problem by enumerating all possible values of $\tilde{\mathbf{c}}_{o}$ within $\mathcal{U}_{d}(\mathbf{s}_{o})$.
-This is a common technique in robust optimization models (Goerigk and Khosravi 2024):
-[…]
-s.t. $t\geq g(\mathbf{z}_{o},\bm{\xi}_{o}),\quad\forall\bm{\xi}_{o}\in\mathcal{U}_{d}(\mathbf{s}_{o}).$
-A discrete distribution can be generated as mentioned before, by combining all the datapoints for the selection options.
-However, this will make the problem intractable as mentioned earlier.
-For comparison purposes, the results will provide a sampling counterpart to that approach in the same spirit as C-SAA.
-Since it is not quite satisfactory, another uncertainty set is considered next.
-3.5.2 Contextual Budgeted Interval Uncertainty Set
-A commonly used uncertainty set is the budgeted interval set $\mathcal{U}_{b}(\mathbf{s}_{o})$, which is defined by an uncertainty budget $B$.
-This budget limits the extent to which all components of the uncertain parameter $\tilde{\mathbf{c}}_{o}$ can simultaneously deviate to their upper bounds.
-This section assumes that the uncertain cost $\tilde{\mathbf{c}}_{o}$ is element-wise positive and $g(\mathbf{z}_{o},\tilde{\mathbf{c}}_{o})$ is a piecewise linear function of the following form:
-$g(\mathbf{z}_{o},\tilde{\mathbf{c}}_{o})=(\tilde{\mathbf{c}}_{o}+\mathbf{a}_{o}^{j})^{T}\mathbf{z}_{o}\quad\text{if}~{}\mathbf{z}_{o}\in[\mathbf{b}^{j-1}_{o},\mathbf{b}^{j}_{o}],~{}j=1,\ldots n,$ (7)
-where $\{\mathbf{b}^{0}_{o},\mathbf{b}^{1}_{o},\ldots,\mathbf{b}^{n}_{o}\}$ are the breakpoints of the piecewise linear function, and $\mathbf{a}_{o}^{j}=(a_{ivo}^{j})_{i\in\mathcal{I},v\in\mathcal{V}}$ is a deterministic cost vector corresponding to the segment $[\mathbf{b}^{j-1}_{o},\mathbf{b}^{j}_{o}]=[(b_{ivo}^{j-1})_{i\in\mathcal{I},v\in\mathcal{V}},(b_{ivo}^{j})_{i\in\mathcal{I},v\in\mathcal{V}}]$.
-Given these assumptions, the typical approach to ensure tractability when using a budgeted interval set is to dualize the inner maximization problem, resulting in the robust counterpart of the original problem (Bertsimas and Sim 2003, Bayram et al.
-2022, Goerigk and Khosravi 2024).
-Let the conditional prediction interval of $\tilde{\mathbf{c}}_{vo}$ be $[\bm{\underline{\xi}}_{vo},\bm{\underline{\xi}}_{vo}+\hat{\bm{\xi}}_{vo}]$, where $\bm{\underline{\xi}}_{vo}=(\underline{\xi}_{ivo})_{i\in\mathcal{I}}$ and $\bm{\underline{\xi}}_{vo}+\hat{\bm{\xi}}_{vo}=(\underline{\xi}_{ivo}+\hat{\xi}_{ivo})_{i\in\mathcal{I}}$ represent the lower and upper estimated conditional quantiles of the cost vector, respectively.
-The budgeted interval set $\mathcal{U}_{b}(\mathbf{s}_{o})$ can then be expressed as follows:
-$\mathcal{U}_{b}(\mathbf{s}_{o},B)=\{\bm{\xi}_{o}\in\mathbb{R}^{|\mathcal{I}||\mathcal{V}|}:\xi_{ivo}=\underline{\xi}_{ivo}+\hat{\xi}_{ivo}\delta_{iv},\sum_{i\in\mathcal{I},v\in\mathcal{V}}\delta_{iv}\leq B,\delta_{iv}\in[0,1]\}.$ (8)
-For a budgeted uncertainty set, it should be noted that learning the entire conditional distribution is unnecessary.
-Instead, learning the lower and upper conditional quantiles is sufficient to construct the conditional prediction interval.
-For instance, to create a 95% conditional prediction interval, only the 2.5% and 97.5% conditional quantiles are required.
-Let $q^{low}$ and $q^{upp}$ be the lower and upper quantiles for the conditional prediction interval of selection option $v$, respectively.
-Then, the lower bound $\bm{\underline{\xi}}_{vo}$ is defined as $\inf\{\mathbf{u}:\mathbb{P}_{v}(\mathbf{c}_{v}\leq\mathbf{u}\mid{\mathbf{s}_{vo}})\geq q^{low}\}$, and the upper bound $(\bm{\underline{\xi}}_{vo}+\hat{\bm{\xi}}_{vo})$ is defined as $\inf\{\mathbf{u}:\mathbb{P}_{v}(\mathbf{c}_{v}\leq\mathbf{u}\mid{\mathbf{s}_{vo}})\geq q^{upp}\}$.
-With a fixed $\mathbf{z}$, the inner maximization problem of Problem (5) is expressed as follows:
-$\max\quad$ $\sum_{i\in\mathcal{I}}\sum_{v\in\mathcal{V}}\hat{\xi}_{ivo}z_{ivo}\delta_{iv}$ (9)
-s.t. $\sum_{i\in\mathcal{I},v\in\mathcal{V}}\delta_{iv}\leq B$ (10)
-$\delta_{iv}\in[0,1],\quad\forall i\in\mathcal{I},v\in\mathcal{V}.$ (11)
-By taking the dual of the inner maximization problem and introducing auxiliary variables, a tractable reformulation of the robust counterpart of Problem (5) is obtained as follows:
-[…]
-s.t. $\pi+\lambda_{iv}\geq\hat{\xi}_{ivo}z_{ivo},\quad\forall i\in\mathcal{I},v\in\mathcal{V}$
-$\pi\geq 0$
-$\lambda_{iv}\geq 0,\quad\forall i\in\mathcal{I},v\in\mathcal{V}$
-$z_{ivo}\leq b^{j}_{ivo}+M(1-y_{j}),\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
-$z_{ivo}\geq b^{j-1}_{ivo}-M(1-y_{j}),\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
-$w_{ivo}^{j}\geq(\underline{\xi}_{ivo}+a_{ivo}^{j})z_{ivo}-M(1-y_{j}),\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
-$y_{j}\in\{0,1\},\quad\forall j\in[n]$
-$w_{ivo}^{j}\geq 0,\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
-where $\pi$ and $(\lambda_{iv})_{i\in\mathcal{I},v\in\mathcal{V}}$ are the dual variables corresponding to Constraints (10) and (11), respectively, and $M$ is a reasonably large constant.
-4 Application: The Ominichannel Multi-Courier Order Fulfillment Problem
-This sections introduces a real-world order fulfillment problem encountered by the industrial partner.
-The proposed CSO framework is then applied to this problem, demonstrating significant advantages over the current practice and common baseline methods.
+3 The Omnichannel Multi-Courier Order Fulfillment Problem
+This section introduces a real-world order fulfillment problem encountered by the industrial partner.
 3.1 High-Level Problem Description
@@
 These centers include primarily physical stores, as well as regional distribution centers and metro e-commerce centers.
-Each location is equipped with multiple eligible carriers capable of providing different levels of service (LOS), such as Ground, Same-Day, and Overnight shipping.
+Each location is equipped with multiple eligible carriers capable of providing different levels of service, such as Ground, Same-Day, and Overnight shipping.
+Throughout this paper, “carrier” refers to one of these specific carrier-service pairs.
 The set of available carriers spans from major shipping companies to crowdsourced gig carriers.
@@
 3.2 The Nominal Optimization Problem
-This section introduces the nominal MILP model developed for optimizing the omnichannel multi-courier order fulfillment problem, which is a special case of the generic problem described in Section 3.1.
+This section introduces the nominal MILP model developed for optimizing the omnichannel multi-courier order fulfillment problem.
 A comprehensive list of the notations presented in this section is provided in Appendix for reference.
@@
 Let $\mathcal{I}$ be the set of SKUs and $\mathcal{V}$ be the set of fulfillment options.
-Here, $\mathcal{V}$ is defined as the Cartesian product $\mathcal{K}\times\mathcal{L}$, where $\mathcal{K}=\{1,\ldots,K\}$ denotes the set of carriers, and =$\mathcal{L}=\{1,\ldots,L\}$ denotes the set of locations.
-Each order $o$ needs to solve an independent optimization problem to decide the selections of carrier-location pairs for each SKU.
-The primary variables are denoted by the vector $\mathbf{z_{o}}=(z_{ik\ell o})_{i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}}$, where each $z_{ik\ell o}\in\mathbb{Z}$ indicates the quantity of SKU $i$ sourced from location $\ell$ and shipped by carrier $k$ in order $o$.
+Here, $\mathcal{V}$ is defined as the Cartesian product $\mathcal{K}\times\mathcal{L}$, where $\mathcal{K}$ denotes the set of carriers, and $\mathcal{L}$ denotes the set of locations.
+Each order needs to solve an independent optimization problem to decide the selection of carrier-location pairs for each SKU.
+The primary decision variables are denoted by the vector $\mathbf{z}=(z_{ik\ell})_{i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}}$, where each $z_{ik\ell}\in\mathbb{Z}$ indicates the quantity of SKU $i$ sourced from location $\ell$ and shipped by carrier $k$.
 3.2.2 Objective Function
-The objective of the optimization problem for order $o$ is to minimize the overall fulfillment costs while ensuring timely delivery.
+The objective of the optimization problem for an order is to minimize the overall fulfillment costs while ensuring timely delivery.
 The fulfillment costs for an order include the fixed sourcing (non-parcel) costs and the shipping (parcel) costs, with consolidation discounts applied when multiple units are assigned to the same carrier-location pair.
@@
 The objective function is formally defined as follows.
-Let $c_{ik\ell o}^{ship}>0$ be the per-unit shipping cost when using location $\ell$ and carrier $k$ to fulfill SKU $i$ in order $o$.
+Let $c_{ik\ell}^{ship}>0$ be the per-unit shipping cost when using location $\ell$ and carrier $k$ to fulfill SKU $i$.
 For each location $\ell$, let $c_{\ell}^{fixed}>0$ be the per-unit fixed sourcing cost.
@@
 Let $\gamma^{+}\geq 0$ and $\gamma^{-}\geq 0$ be constants that convert late and early delivery penalties into per-unit costs, respectively.
-In an ideal setting where all carrier-location deviations are known at the time of fulfillment, let the vector $\mathbf{d}_{o}=(d_{k\ell o})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ represent the realized uncertain delivery deviations for order $o$, where each $d_{k\ell o}\in\mathbb{R}$ denotes the number of days that the carrier-location pair $(k,\ell)$ deviates from desired delivery date.
+In an ideal setting where all carrier-location deviations are known at the time of fulfillment, let the vector $\mathbf{d}=(d_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ represent the realized uncertain delivery deviations, where each $d_{k\ell}$ denotes the number of days that the carrier-location pair $(k,\ell)$ deviates from desired delivery date.
 A positive deviation indicates late delivery, a zero deviation indicates on-time delivery, and a negative deviation indicates early delivery.
-The weighted objective function for order $o$ can be constructed as follows:
-$g(\mathbf{z}_{o},\mathbf{d}_{o})=\begin{cases}\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}[c_{\ell}^{fixed}+(1-\beta_{k})c_{ik\ell o}^{ship}+\gamma^{+}d_{k\ell o}^{+}+\gamma^{-}d_{k\ell o}^{-}]z_{ik\ell o}&\text{if }\sum_{i\in\mathcal{I}}z_{ik\ell o}\geq 2\\ \sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell o}^{ship}+\gamma^{+}d_{k\ell o}^{+}+\gamma^{-}d_{k\ell o}^{-})z_{ik\ell o}&\text{otherwise},\end{cases}$ (13)
-where $d_{k\ell o}^{+}=\max\{0,d_{k\ell o}\}$ and $d_{k\ell o}^{-}=\max\{0,-d_{k\ell o}\}$.
-This objective function encourages item consolidation within an order by offering a carrier-specific discount when applicable.
-In addition, it leverages an asymmetric penalty cost function for delivery deviations, with $\gamma^{+}\gg\gamma^{-}$ to primarily discourage late deliveries while also disincentivizing early deliveries to a lesser extent.
-The objection function is piecewise linear and can be linearized using standard rewritings.
+The weighted objective function for an order can be constructed as follows:
+$g(\mathbf{z},\mathbf{d})=\begin{cases}\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}[c_{\ell}^{fixed}+(1-\beta_{k})c_{ik\ell}^{ship}+\gamma^{+}d_{k\ell}^{+}+\gamma^{-}d_{k\ell}^{-}]z_{ik\ell}&\text{if }\sum_{i\in\mathcal{I}}z_{ik\ell}\geq 2,\\ \sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell}^{ship}+\gamma^{+}d_{k\ell}^{+}+\gamma^{-}d_{k\ell}^{-})z_{ik\ell}&\text{otherwise},\end{cases}$ (1)
+where $d_{k\ell}^{+}=\max\{0,d_{k\ell}\}$ and $d_{k\ell}^{-}=\max\{0,-d_{k\ell}\}$.
+This objective function encourages item consolidation within an order by applying a carrier-specific discount if at least two units are sourced from the same location and shipped by the same carrier.
+In addition, it leverages an asymmetric penalty cost function for delivery deviations, where $\gamma^{+}\gg\gamma^{-}$.
+This asymmetry strongly penalizes late deliveries to discourage delays while also applying a smaller penalty to disincentivize early deliveries.
+Specifically, the penalties can be interpreted as the opportunity cost associated with early deliveries and the lost-sale cost associated with late deliveries.
+The objective function is piecewise linear and can be linearized using standard rewritings.
 3.2.3 Operational Constraints
 Several operational constraints ensure that the selection of fulfillment center locations and carriers is feasible in practice.
-For location $\ell\in\mathcal{L}$, let $\operatorname{inv}_{i\ell o}\geq 0$ be the inventory level of SKU $i$ at location $\ell$ and $\operatorname{cap}_{\ell o}\geq 0$ be the available capacity at location $\ell$ when order $o$ is placed.
-For SKU $i$ in order $o$, let $q_{io}\geq 0$ be the quantity of SKU $i$ in order $o$, and let $e_{ik\ell o}\in\{0,1\}$ be a binary indicator of whether the location-carrier pair $(\ell,k)$ is an eligible fulfillment option.
+For location $\ell\in\mathcal{L}$, let $\operatorname{inv}_{i\ell}\geq 0$ be the inventory level of SKU $i$ at location $\ell$ and $\operatorname{cap}_{\ell}\geq 0$ be the available capacity at location $\ell$ when the order is placed.
+Let $q_{i}\geq 0$ be the quantity of SKU $i$ in the order, and let $e_{ik\ell}\in\{0,1\}$ be a binary indicator of whether the location-carrier pair $(\ell,k)$ is an eligible fulfillment option for SKU $i$.
+An assignment $\mathbf{z}$ in the feasible set $\mathcal{Z}$ must obey the following constraints.
   • Each unit of an SKU must be sourced from exactly one location and shipped by exactly one carrier.
-    $\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}z_{ik\ell o}=q_{io},\quad\forall i\in\mathcal{I}$ (14)
+    $\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}z_{ik\ell}=q_{i},\quad\forall i\in\mathcal{I}$ (2)
   • Each SKU can only be fulfilled by eligible candidate location-carrier pairs, which may vary depending on the carrier’s coverage and package requirements.
-    $z_{ik\ell o}\leq q_{io}e_{ik\ell o},\quad\forall i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}$ (15)
+    $z_{ik\ell}\leq q_{i}e_{ik\ell},\quad\forall i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}$ (3)
   • Only locations with available inventory of the requested SKUs can be chosen.
-    $\sum_{k\in\mathcal{K}}z_{ik\ell o}\leq\operatorname{inv}_{i\ell o},\quad\forall i\in\mathcal{I},\ell\in\mathcal{L}$ (16)
+    $\sum_{k\in\mathcal{K}}z_{ik\ell}\leq\operatorname{inv}_{i\ell},\quad\forall i\in\mathcal{I},\ell\in\mathcal{L}$ (4)
   • Each location has a daily capacity limit on the total number of units it can process.
-    $\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}z_{ik\ell o}\leq\operatorname{cap}_{\ell o},\quad\forall\ell\in\mathcal{L}$ (17)
-4.3 Applying the CSO Framework to the Problem
+    $\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}z_{ik\ell}\leq\operatorname{cap}_{\ell},\quad\forall\ell\in\mathcal{L}$ (5)
+3.3 The Contextual Stochastic Order Fulfillment Problem
 In practice, the delivery deviations of each carrier-location pair cannot be perfectly estimated in advance.
-To account for this uncertainty for each order $o$, this section models the vector of uncertain delivery deviations $\tilde{\mathbf{d}}_{o}=(\tilde{d}_{k\ell o})_{k\in\mathcal{K},\ell\in\mathcal{L}}\sim\mathbb{P}$ as a $(KL)$-dimensional random variable governed by the probability distribution $\mathbb{P}$.
-In addition, let $\mathbf{s}_{o}=(\mathbf{s}_{k\ell o})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ denote the state vector observable for order $o$ before the fulfillment decision is made, where each $\mathbf{s}_{k\ell o}$ is a vector of covariates encompassing information about the fulfillment network status and order details related to carrier-location pair $(k,\ell)$.
-The conditional distribution of the deviation vector, upon observing $\mathbf{s}_{o}$, is denoted by $\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})$, and the individual deviation condition distribution is $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell o})$ for each $k\in\mathcal{K}$ and $\ell\in\mathcal{L}$.
-Applying the CSO framework leads to the following contextual stochastic order fulfillment problem (CSOFP):
-(CSOFP) $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})]=\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\Pi_{k\in\mathcal{K},\ell\in\mathcal{L}}\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell o})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})].$ (18)
-Effectively solving Problem (18) requires leveraging data from historical orders.
-As a specific instance of Problem (2), CSOFP also has an observational dataset.
+To account for this uncertainty for each order, this section models the vector of uncertain delivery deviations $\tilde{\mathbf{d}}=(\tilde{d}_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}\sim\mathbb{P}$ as a $(|\mathcal{K}||\mathcal{L}|)$-dimensional random vector governed by a joint probability distribution $\mathbb{P}$.
+Since deviations measured in days only take on discrete values in the dataset, each $\tilde{d}_{k\ell}$ inherently follows a discrete probability distribution.
+It is further assumed that every $\tilde{d}_{k\ell}$ shares the same finite, ordered support.
+To formalize this, let $C$ be the total number of distinct deviation values and denote them by $\xi_{1}<\xi_{2}<\ldots<\xi_{C}$.
+By definition, any realization $d_{k\ell}$ of $\tilde{d}_{k\ell}$ must lie in this set: $d_{k\ell}\in\{\xi_{1},\xi_{2},\ldots,\xi_{C}\}.$
+In addition, let $\mathbf{s}=(\mathbf{s}_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ denote the state vector observable to the order before the fulfillment decision is made, where each $\mathbf{s}_{k\ell}$ is a vector of covariates encompassing information about order details and fulfillment network status related to carrier-location pair $(k,\ell)$.
+The contextual distribution of the deviation vector, upon observing $\mathbf{s}$, is denoted by $\mathbb{P}(\mathbf{d}\mid\mathbf{s})$, and the individual deviation contextual distribution is $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell})$ for each $k\in\mathcal{K}$ and $\ell\in\mathcal{L}$.
+This leads to the following contextual stochastic order fulfillment problem (CSOFP):
+(CSOFP) $\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s})}[g(\mathbf{z},\tilde{\mathbf{d}})].$ (6)
+Solving Problem (6) in practice requires leveraging data from past orders.
+Unlike classical CSO formulations where the entire random cost vector is observed, CSOFP is based on an observational dataset of past decisions and their realized costs.
 Let $\mathcal{O}$ be the set of historical orders.
-Let $\mathbf{z}_{o}=(z_{ik\ell o})_{i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}}$ be the vector of historical decisions made by the industrial partner’s sourcing engine for order $o$.
-There are numerous missing entries in the observed deviation vector $\mathbf{d}_{o}$ since actual observations depend on the decision vector $\mathbf{z}_{o}$.
-Let
-$\mathbf{D}_{o}:=\{d_{k\ell o}\mid\sum_{i\in\mathcal{I}}z_{ik\ell o}>0,k\in\mathcal{K},\ell\in\mathcal{L}\}$
-be the observable costs for order $o$. The dataset available to estimate $\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})$ is $\mathbb{D}=\{(\mathbf{s}_{o},\mathbf{D}_{o})\}_{o=1}^{|\mathcal{O}|}$.
-$\mathbb{D}$ is partitioned into subsets to facilitate learning the individual condition distributions.
-For carrier $k$ and location $\ell$, the dataset for learning $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell o})$ is $\mathbb{D}_{k\ell}=\{(\mathbf{s}_{k\ell m},d_{k\ell m})\}_{m=1}^{M_{k\ell}}$, where $M_{k\ell}$ is the number of times the deviation has been observed for the carrier-location pair $(k,\ell)$ in $\mathbb{D}$.
-4.4 Learning the Contextual Distribution of Delivery Deviations
-This section describes how to learn Contextual Distribution of Delivery Deviations.
-It begins by selecting features to construct the covariate vector for each carrier-location pair.
-To predict the conditional distribution of deviations, two classes of machine learning methods are proposed: probabilistic multi-class classification, which leverages the discrete nature of deviations, and quantile regression, which is known for its effectiveness in distributional prediction (Salari et al.
+For each $o\in\mathcal{O},$ define the decision vector $\mathbf{z}_{o}=(z_{ik\ell,o})_{i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}}.$ Since deviations were observed only for chosen $(i,k,\ell)$ triples, the full deviation vector $\mathbf{d}_{o}=(d_{k\ell,o})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ contains many unobserved entries.
+The subset of realized deviations for order $o$ is:
+$\mathbf{D}_{o}:=\{d_{k\ell,o}\mid\sum_{i\in\mathcal{I}}z_{ik\ell,o}>0,k\in\mathcal{K},\ell\in\mathcal{L}\}.$
+The resulting dataset available for estimating the contextual distribution $\mathbb{P}(\mathbf{d}\mid\mathbf{s})$ is $\mathbb{D}=\{(\mathbf{s}_{o},\mathbf{D}_{o})\}_{o=1}^{|\mathcal{O}|}$.
+4 Methodology: A Generic CSO Framework for Observational Data
+Solving CSOFP under observational data requires: (1) learning a contextual distributional oracle from partial realizations, and (2) optimizing with that oracle in a tractable way.
+This section lays out a generic contextual stochastic optimization (CSO) framework that does exactly that.
+In Section 4.1, the abstract CSO problem (denoted by ${\cal P}$) is defined.
+Section 4.2 shows how to construct a contextual distribution oracle $\mathcal{M}(\mathbf{s})$ from partial observations.
+Sections 4.3 and 4.4 introduce two solution paradigms—Contextual Sample Average Approximation and Contextual Robust Optimization—which use $\mathcal{M}(\mathbf{s})$ as input.
+Finally, Section 4.6 maps all of the above back to the omnichannel multi-courier CSOFP introduced in Section 3.
+4.1 Generic Contextual Stochastic Optimization (CSO) Problem
+Problem ${\cal P}$ consists of fulfilling a set of incoming requests sequentially, each of which specifies demands for a subset of products.
+Let $\mathcal{I}$ be the set of products, $\mathcal{V}$ the set of selection options (e.g., sourcing or transport modes), and $\mathcal{Z}\subseteq\mathbb{Z}_{+}^{|\mathcal{I}|\times|\mathcal{V}|}$ the feasible assignment set.
+Each request arrives with known demand quantities $\{q_{i}:i\in\mathcal{I}\}$ at decision time.
+The decision variable $\mathbf{z}=(z_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}\in\mathcal{Z}$ assigns quantity $z_{iv}$ of product $i$ to option $v$.
+Let $\tilde{\mathbf{c}}=(\tilde{c}_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}\sim\mathbb{P}$ denote the random cost vector.
+Assume that $\tilde{\mathbf{c}}$ is element-wise positive and the cost function $g(\mathbf{z},\tilde{\mathbf{c}})$ is piecewise linear in $\mathbf{z}$.
+Introduce, for $j=0,1,\dots,n$, the breakpoint vectors
+[…]
+so that each segment is defined as
+$[\mathbf{b}^{j-1},\,\mathbf{b}^{j}]=\bigl{\{}\mathbf{z}:b^{\,j-1}_{iv}\leq z_{iv}\leq b^{j}_{iv}\;\forall\,i\in\mathcal{I},\;v\in\mathcal{V}\bigr{\}}.$
+Let $\mathbf{a}^{j}=(a^{j}_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}$ be the marginal‐cost adjustment on segment $j$. Then
+$g(\mathbf{z},\tilde{\mathbf{c}})=(\tilde{\mathbf{c}}+\mathbf{a}^{j})^{\mathsf{T}}\mathbf{z}\quad\text{if}\quad\mathbf{z}\in[\mathbf{b}^{j-1},\mathbf{b}^{j}],\;j=1,\dots,n,$ (7)
+thereby capturing volume‐dependent pricing behaviors.
+For each request, the fulfillment decision solves the risk‐neutral stochastic optimization problem
+$\min_{\mathbf{z}\in\mathcal{Z}}\;\mathbb{E}_{\tilde{\mathbf{c}}\sim\mathbb{P}}\bigl{[}g(\mathbf{z},\tilde{\mathbf{c}})\bigr{]}.$ (8)
+The piecewise‐linear cost in (7) partitions the decision space into shipment‐size tiers determined by the breakpoints $\{\mathbf{b}^{j}\}$.
+Within tier $j$, the marginal cost is $\tilde{\mathbf{c}}+\mathbf{a}^{j}$, which shifts to $\tilde{\mathbf{c}}+\mathbf{a}^{j+1}$ upon entering tier $j+1$.
+This formulation accommodates economies of scale (declining unit costs at higher volumes), quantity discounts (step‐down pricing beyond thresholds), capacity surcharges (unit‐cost premiums when exceeding lower‐cost bands), and other tiered‐pricing schemes.
+In Problem (8), the true distribution $\mathbb{P}$ is not readily available.
+However, side information often exists for each product-option pair.
+Let $\mathbf{s}_{iv}$ denote the covariate vector for product $i$ under selection option $v$ (e.g., network state, weather, current demand).
+All these covariate vectors are aggregated into the state vector $\mathbf{s}=(\mathbf{s}_{iv})_{i\in\mathcal{I},v\in\mathcal{V}}$, which encompasses all contextual information available at the time of decision.
+The distribution $\mathbb{P}$ is then approximated by the conditional distribution $\mathbb{P}(\mathbf{c}\mid\mathbf{s})$, yielding the Contextual Stochastic Optimization (CSO) problem
+$\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s})}[g(\mathbf{z},\mathbf{\tilde{c}})].$ (9)
+Let $\tilde{\mathbf{c}}_{v}=(\tilde{c}_{iv})_{i\in\mathcal{I}}$ and $\mathbf{s}_{v}=(\mathbf{s}_{iv})_{i\in\mathcal{I}}$ for each $v\in\mathcal{V}$.
+Under the assumption that, conditional on its own covariates, each component $c_{v}$ is independent across options $v\in\mathcal{V}$, one has
+$\mathbb{P}(\mathbf{c}\mid\mathbf{s})=\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{v}),$
+where $\mathbb{P}_{v}(\cdot\mid\mathbf{s}_{v})$ is the contextual distribution for option $v$.
+This assumption is reasonable when each option’s cost depends primarily on its own covariates.
+Shared features, such as weather, create similarity but do not induce direct dependence.
+Consequently, Problem (9) can be written as
+$\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{c}}\sim\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{v})}[g(\mathbf{z},\mathbf{\tilde{c}})].$ (10)
+4.2 Constructing a Contextual Distribution Oracle from Observational Data
+The contextual distribution $\mathbb{P}(\mathbf{c}\mid\mathbf{s})$ is also generally unknown but can often be inferred from historical observations.
+In many real-world COS applications (including CSOFP) however, data are observational, i.e., only the costs corresponding to decisions actually taken were recorded.
+As a result, the dataset available to estimate $\mathbb{P}(\mathbf{c}\mid\mathbf{s})$ takes the form $\mathbb{D}=\{(\mathbf{s}_{o},\mathbf{C}_{o})\}_{o=1}^{|\mathbb{D}|}$, where $\mathbf{s}_{o}=(s_{iv,o})_{i\in\mathcal{I},\;v\in\mathcal{V}}~{}\text{and}~{}\mathbf{C}_{o}=(\mathbf{c}_{v,o})_{v\in\mathcal{V}}$ collect the context and realized costs, with $\mathbf{c}_{v,o}$ observed only if option $v$ was selected in request $o$.
+The learning task consists in finding a full‑vector forecasting oracle $\mathcal{M}$ that approximates $\mathbb{P}(\mathbf{c}\mid\mathbf{s})$ in a supervised manner using the dataset $\mathbb{D}$: $\mathcal{M}$ receives a context $\mathbf{s}$ as input and returns a probability distribution $\mathcal{M}(\mathbf{s})$ on the cost vector $\tilde{\mathbf{c}}$.
+Under the conditional independence assumption from the previous section, learning $\mathbb{P}(\mathbf{c}\mid\mathbf{s})$ amounts to learning $\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{v})$ for each selection option $v$ separately.
+Partition the full dataset $\mathbb{D}$ into subsets $\mathbb{D}_{v}=\{(\mathbf{s}_{v,1},\mathbf{c}_{v,1}),(\mathbf{s}_{v,2},\mathbf{c}_{v,2}),\ldots\},$ representing the dataset for option $v$.
+A machine learning model $\mathcal{M}_{v}$ is then trained on $\mathbb{D}_{v}$ to approximate each $\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{v})$.
+Assemble the oracle as the product $\mathcal{M}(\mathbf{s})=\bigl{(}\mathcal{M}_{v}(\mathbf{s}_{v})\bigr{)}_{v\in\mathcal{V}},$ which generates joint forecasts by treating each component independently.
+Note that if every $\mathbf{s}_{v}$ shares the same structure and dimension, a single learning architecture can serve as $\mathcal{M}_{v}$ for all $v\in\mathcal{V}$, trained on the union $\cup_{v\in\mathcal{V}}\mathbb{D}_{v}$.
+By reducing the high‑dimensional joint learning task to $|\mathcal{V}|$ manageable subproblems, this decomposition turns observational data into a black‑box contextual distribution oracle $\mathcal{M}(\mathbf{s})$, which will serve as the input for the solution methods developed in Section 4.3 (C-SAA) and Section 4.4 (C-RO).
+4.3 Contextual Sample Average Approximation (C-SAA)
+Problem (10) is approximated by generalizing the Sample Average Approximation (SAA) method (Kleywegt et al.
+2002) to contextual stochastic optimization.
+Given a request with covariates $\mathbf{s}$, the Contextual Sample Average Approximation (C-SAA) proceeds in two stages:
+  1. 1.
+    Candidate generation.
+    1. (a)
+      Draw $N_{1}$ independent cost vector samples $\{\bm{\xi}^{n}\}_{n=1}^{N_{1}}\sim\mathcal{M}(\mathbf{s})$, where $\bm{\xi}^{n}=(\xi^{n}_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}$.
+    2. (b)
+      Solve
+      $\mathbf{z}_{i}\;=\;\arg\min_{\mathbf{z}\in\mathcal{Z}}\;\frac{1}{N_{1}}\sum_{n=1}^{N_{1}}g(\mathbf{z},\bm{\xi}^{n}),\quad i=1,\dots,Q,$
+      independently $Q$ times to obtain candidate solutions $\{\mathbf{z}_{1},\dots,\mathbf{z}_{Q}\}$ with sample‐average objectives $\{f_{1},\dots,f_{Q}\}$.
+    3. (c)
+      Compute the average objective $\underaccent{\bar}{f}=\tfrac{1}{Q}\sum_{i=1}^{Q}f_{i}$, which is a statistical lower bound on the true optimal value $f^{*}$ because $\underaccent{\bar}{f}$ is an unbiased estimator of $\mathbb{E}_{{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s})}}[\underaccent{\bar}{f}]$ and $\mathbb{E}_{{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s})}}[\underaccent{\bar}{f}]\leq f^{*}$.
+  2. 2.
+    Candidate evaluation.
+    1. (a)
+      For each $\mathbf{z}_{i}$, draw a larger set of evaluation samples $\{\bm{\xi}^{n}\}_{n=1}^{N_{2}}\sim\mathcal{M}(\mathbf{s})$, with $N_{2}\gg N_{1}$.
+    2. (b)
+      Compute the re‐estimated objectives $\bar{f}(\mathbf{z}_{i})=\tfrac{1}{N_{2}}\sum_{n=1}^{N_{2}}g(\mathbf{z}_{i},\bm{\xi}^{n})$.
+    3. (c)
+      Each $\bar{f}(\mathbf{z}_{i})$ is an upper bound on $f^{*}$ since each candidate solution is feasible; the difference $\bar{f}(\mathbf{z}_{i})-\underaccent{\bar}{f}$ approximates the optimality gap.
+    4. (d)
+      Select the candidate $\mathbf{z}^{*}$ with the smallest estimated gap.
+Discussion.
+Unlike the reweighting approach of Bertsimas and Kallus (2020), which leverages complete historical samples for each fully-observed cost vector, C–SAA relies solely on observational data (i.e., only the costs of the selected options are available).
+Attempting to form all possible cost vector combinations from partial observations would lead to a combinatorial explosion as $|\mathcal{V}|$ and $|\mathbb{D}|$ grow, making such method intractable in large, high‐dimensional selection spaces.
+4.4 Contextual Robust Optimization (C-RO)
+The performance of C-SAA largely hinges on the accuracy of the estimated contextual distribution.
+Robust optimization provides a hedge against misspecification of the contextual distribution by guarding against worst–case scenarios, making it a viable approach for contextual stochastic optimization.
+Let $\mathcal{U}(\mathbf{s})\subset\mathbb{R}^{|\mathcal{I}|\times|\mathcal{V}|}$ represent a contextual uncertainty set for the cost vector $\tilde{\mathbf{c}}$.
+The Contextual Robust Optimization (C-RO) problem is
+$\min_{\mathbf{z}\in\mathcal{Z}}\max_{\bm{\xi}\in\mathcal{U}(\mathbf{s})}g(\mathbf{z},\bm{\xi}).$ (11)
+Two constructions of $\mathcal{U}(\mathbf{s})$ are considered: a discrete sampling–based set $\mathcal{U}_{d}(\mathbf{s})$ and a budgeted interval set $\mathcal{U}_{b}(\mathbf{s},B)$.
+4.4.1 Contextual Discrete Uncertainty Set
+Problem (11) can be reformulated into a single-level problem by enumerating every possible realization of the random vector $\tilde{\mathbf{c}}$, a standard technique in robust optimization.
+In principle, a discrete distribution can be generated as mentioned before, by combining all combinations of the observed data points.
+However, this will make the problem intractable as discussed.
+To retain tractability, a sampling-based discrete uncertainty set in the same spirit of C-SAA is introduced.
+Let $\mathcal{U}_{d}(\mathbf{s})=\{\bm{\xi}^{1},\ldots,\bm{\xi}^{N}\}$ denote a collection of $N$ deviation scenarios drawn from $\mathcal{M}(\mathbf{s})$.
+The robust counterpart of (11) becomes
+[…]
+s.t. $t\geq g(\mathbf{z},\bm{\xi}^{n}),\quad n=1,\dots,N.$
+4.4.2 Contextual Budgeted Interval Uncertainty Set
+An alternative uncertainty set is the budgeted interval set $\mathcal{U}_{b}(\mathbf{s},B)$ (Bertsimas and Sim 2004), which is defined by a robust budget $B$.
+This budget limits the extent to which limits how many components of the uncertain parameter $\tilde{\mathbf{c}}$ can simultaneously deviate to their upper bounds.
+Rather than estimating the full joint distribution, only the conditional lower and upper quantiles at levels $q_{\rm low}$ and $q_{\rm up}$ are needed to define a $(q_{\rm low},q_{\rm up})$ prediction interval for each option.
+For each option $v\in\mathcal{V}$, let $\underline{\bm{\xi}}_{v}=\bigl{(}\underline{\xi}_{iv}\bigr{)}_{i\in\mathcal{I}}$ and $\overline{\bm{\xi}}_{v}=\bigl{(}\overline{\xi}_{iv}\bigr{)}_{i\in\mathcal{I}}$ denote the estimated conditional $q_{\rm low}$- and $q_{\rm up}$-quantiles extracted from $\mathcal{M}_{v}(\mathbf{s}_{v})$:
+$\underline{\bm{\xi}}_{v}=\inf\{\mathbf{u}:\mathbb{P}_{v}(\mathbf{c}_{v}\leq\mathbf{u}\mid\mathbf{s}_{v})\geq q_{\rm low}\},\quad\overline{\bm{\xi}}_{v}=\inf\{\mathbf{u}:\mathbb{P}_{v}(\mathbf{c}_{v}\leq\mathbf{u}\mid\mathbf{s}_{v})\geq q_{\rm up}\}.$
+Given a budget $B$, the budgeted interval uncertainty set is
+$\mathcal{U}_{b}(\mathbf{s},B)=\Bigl{\{}\bm{\xi}:\;\xi_{iv}=\underline{\xi}_{iv}+\delta_{iv}\,(\overline{\xi}_{iv}-\underline{\xi}_{iv}),\;\sum_{i,v}\delta_{iv}\leq B,\;\delta_{iv}\in[0,1]\Bigr{\}}.$
+With a fixed $\mathbf{z}$, the inner maximization problem in (11) is
+$\max_{\delta_{iv}}\quad$ $\sum_{i,v}(\overline{\xi}_{iv}-\underline{\xi}_{iv})z_{iv}\delta_{iv}$ (13)
+s.t. $\sum_{i,v}\delta_{iv}\leq B$ (14)
+$\delta_{iv}\in[0,1],\quad\forall i,v.$ (15)
+Let $y_{j}\in\{0,1\}$ and $w_{iv}^{j}\geq 0$ be auxiliary variables.
+Denote by $\pi\geq 0$ and $\lambda_{iv}\geq 0$ the dual variables corresponding to Constraints (14) and (15), respectively.
+Finally, let $M$ be a sufficiently large constant.
+A valid choice for the “big‑M” constant is $M\;=\;\max_{i,v,j}\bigl{\{}\,b_{iv}^{n}-b_{iv}^{1},\;b_{iv}^{\,n-1},\;(\underline{\xi}_{iv}+a_{iv}^{j})\,b_{iv}^{n}\bigr{\}}.$ By taking the dual of the inner maximization problem and introducing $y_{j}$ and $w_{iv}^{j}$, a robust counterpart of (11) is obtained as
+[…]
+s.t. $\pi+\lambda_{iv}\geq(\overline{\xi}_{iv}-\underline{\xi}_{iv})z_{iv},\quad\forall i,v$
+$z_{iv}\leq b^{j}_{iv}+M(1-y_{j}),\quad\forall i,v,j$
+$z_{iv}\geq b^{j-1}_{iv}-M(1-y_{j}),\quad\forall i,v,j$
+$w_{iv}^{j}\geq(\underline{\xi}_{iv}+a_{iv}^{j})z_{iv}-M(1-y_{j}),\quad\forall i,v,j$
+$y_{j}\in\{0,1\},~{}w_{iv}^{j}\geq 0,~{}\pi\geq 0,~{}\lambda_{iv}\geq 0.$
+4.5 Performance Guarantees on C-SAA & C-RO
+C–SAA replaces the unknown cost distribution with Monte-Carlo samples generated by a learned contextual oracle and minimizes the resulting empirical objective.
+Under standard regularity conditions, the SAA objective is statistically consistent: its optimal value converges exponentially fast to the true optimum as $N_{1}$ grows (Kleywegt et al.
+2002, Shapiro et al.
+2021).
+When the oracle is well-calibrated and $N_{1}$ is large, C–SAA thus attains the smallest expected cost.
+However, it remains sensitive to distributional misspecification and finite-sample error, which can expose the solution to rare but severe tail events.
+C–RO mitigates this risk by optimizing the worst-case cost over a prescribed uncertainty set.
+Let a feasible solution $\mathbf{z}$ satisfy $\max_{\bm{\xi}\in\mathcal{U}(\mathbf{s})}g(\mathbf{z},\bm{\xi})\leq t.$ For a discrete uncertainty set built from $N$ samples (C-RO-D), one obtains with probability $1-\delta$ that
+$\Pr\bigl{(}g(\mathbf{z},\tilde{\mathbf{c}})>t\bigr{)}\;\leq\;\epsilon$
+provided
+[…]
+samples are drawn (Calafiore and Campi 2006, Campi and Garatti 2008, Bertsimas et al.
+2021).
+Although this bound grows linearly in the dimension of the uncertainty $|\mathcal{V}||\mathcal{I}|$, it too can degrade under distributional misspecification.
+Under a budgeted-interval uncertainty set (C-RO-B), the violation probability instead satisfies
+$\Pr\bigl{(}g(\mathbf{z},\tilde{\mathbf{c}})>t\bigr{)}\;\leq\;\exp\!\Bigl{(}-\frac{B^{2}}{2\,|\mathcal{V}||\mathcal{I}|}\Bigr{)},$
+so that the budget parameter $B$ directly controls the trade-off between average-case performance and robustness (Bertsimas and Sim 2004).
+This tunable flexibility makes C–RO-B an attractive middle ground between C–SAA and C–RO-D.
+Extending these guarantees to bound worst-case regret or post-decision surprise under partial-feedback settings—where counterfactual costs are unobserved—remains an important direction for future work.
+4.6 Application to the Omnichannel Multi-Courier CSOFP
+The generic CSO framework applies directly to CSOFP via the following identifications:
+$\mathcal{I}\;\mapsto\;\{\text{SKUs}\},$ $\mathcal{V}\;\mapsto\;\mathcal{K}\times\mathcal{L},$
+$z_{iv}\;\mapsto\;z_{ik\ell},$ $\mathbf{s}_{v}\;\mapsto\;\mathbf{s}_{k\ell},$
+$\tilde{\mathbf{c}}=(\tilde{c}_{v})_{v\in\mathcal{V}}\;\mapsto\;\tilde{\mathbf{d}}=(\tilde{d}_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}.$ $\mathbb{P}_{v}\;\mapsto\;\mathbb{P}_{k\ell}.$
+Under the conditional independence assumption,
+$\mathbb{P}({\mathbf{d}}\mid\mathbf{s})=\prod_{k\in\mathcal{K}}\prod_{\ell\in\mathcal{L}}\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell}).$
+Hence CSOFP (Problem 6) can be rewritten as
+$\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{d}}\sim\Pi_{k,\ell}\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell})}[g(\mathbf{z},\tilde{\mathbf{d}})].$ (17)
+To learn each marginal distribution $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell})$, partition the historical dataset into
+$\mathbb{D}_{k\ell}=\bigl{\{}(\mathbf{s}_{k\ell,m},\,d_{k\ell,m})\bigr{\}}_{m=1}^{M_{k\ell}},$
+and train a per‑carrier model $\mathcal{M}_{k}$ on $\mathbb{D}_{k}=\bigcup_{\ell\in\mathcal{L}}\mathbb{D}_{k\ell}$.
+The full contextual distribution oracle is $\mathcal{M}(\mathbf{s})=\bigl{(}\mathcal{M}_{k}(\mathbf{s}_{k\ell})\bigr{)}_{k\in\mathcal{K},\,\ell\in\mathcal{L}}.$
+This per‑carrier decomposition simplifies learning, generalizes to new locations, and outperforms a single global model according to preliminary experiments.
+4.6.1 Applying C-SAA and C-RO
+Computing the expectation in $\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathcal{M}(\mathbf{s})}\bigl{[}g(\mathbf{z},\tilde{\mathbf{d}})\bigr{]}$ directly would require either enumerating all $C^{|\mathcal{K}|\times|\mathcal{L}|}$ deviation vectors or numerically integrating a continuous CDF—both computationally intractable.
+Instead, the two solution paradigms use the contextual distribution oracle $\mathcal{M}(\mathbf{s})$ as follows:
+  • C-SAA replaces the expectation by a sample average: draw $N_{1}$ full‑vector scenarios $\{\tilde{\mathbf{d}}^{n}\}\sim\mathcal{M}(\mathbf{s})$ and solve
+    $\min_{\mathbf{z}\in\mathcal{Z}}\;\frac{1}{N_{1}}\sum_{n=1}^{N_{1}}g\bigl{(}\mathbf{z},\tilde{\mathbf{d}}^{n}\bigr{)}.$
+  • C-RO builds an uncertainty set from oracle outputs—either the discrete scenario set $\{\tilde{\mathbf{d}}^{n}\}_{n\in[N]}$ or a budgeted‑interval set from marginal quantiles—and then solves
+    $\min_{\mathbf{z}\in\mathcal{Z}}\;\max_{\mathbf{d}\in\mathcal{U}(\mathbf{s})}g(\mathbf{z},\mathbf{d}).$
+Both approaches avoid the exponential support growth while remaining agnostic to the internal structure of $\mathcal{M}$. A full C‑RO‑B reformulation is provided in the Appendix.
+5 Learning the Contextual Distribution of Delivery Deviations for CSOFP
+Having defined in Section 4 a generic contextual distribution oracle $\mathcal{M}(\mathbf{s})$, this section describes its concrete implementation for CSOFP.
+It begins by selecting features to construct the covariate vector.
+Two classes of machine learning methods are then proposed to estimate the contextual distribution of deviations: (1) Probabilistic multi-class classification, which leverages the discrete support of deviations; and (2) Tree-based quantile regression, which directly estimates the conditional CDF and avoids quantile crossing (Salari et al.
 2022).
-For simplicity, the index $k\ell$ is omitted in the notations used throughout this section.
-The data set can then be denoted by $\mathbb{D}=\{(\mathbf{s}_{m},d_{m})\}_{m\in[M]}$.
-4.4.1 Contextual Features Selection
+For simplicity, the carrier index $k$ is omitted in the notation, and the dataset for training the per-carrier model is written as $\mathbb{D}=\{(\mathbf{s}_{m},d_{m})\}_{m\in[M]}$.
+5.1 Contextual Feature Selection
 Carefully selecting contextual features is crucial for constructing informed covariates for delivery deviation predictions, as various stages in the order life cycle can contribute to deviations.
-To identify the most important features for predicting deviations, a feature importance analysis was conducted using a tree-based mean regression model.
-Two methods were examined: Mean Decrease in Impurity (MDI) and Shapley values (SHAP).
-MDI measures the importance of a feature by calculating the total reduction in impurity (e.g., Gini impurity or entropy) it brings when used in decision trees (Breiman 2001).
-SHAP values, on the other hand, provide a unified measure of feature importance based on cooperative game theory, attributing to each feature its contribution to the prediction by considering all possible feature combinations (Lundberg et al.
-2020).
-The features considered are categorized into three main classes, each corresponding to different stages of the order fulfillment process:
+Candidate features fall into three categories:
   1. 1.
-    Order-level features: This category includes customer location, SKU weight and dimensions, SKU type, quantities, release hour, release day, and planned lead time (i.e., the difference between the promised delivery date and the order release time).
+    Order-level: This includes customer location, SKU weight and dimensions, SKU type, quantities, release hour, release day, and planned lead time (i.e., the difference between the promised delivery date and the order release time).
 These factors can influence deviations through logistical complexity, handling requirements, and time-sensitive elements affecting the order’s processing and delivery schedule.
   2. 2.
-    Fulfillment center features: This includes coordinates, location type, on-hand capacity, and on-hand inventory.
+    Fulfillment center: This includes coordinates, location type, on-hand capacity, and on-hand inventory.
 The geographical location and type of fulfillment center can affect the efficiency of order processing, while capacity and inventory levels determine the speed and reliability of fulfilling orders.
   3. 3.
-    Carrier-related features: This includes the level of service, promised time in transit, shipping charges, and carrier zone.
+    Carrier-related: This includes service level, promised transit time, shipping charges, and carrier zone.
 These factors impact the carrier ability to meet delivery deadlines, with service level and transit times directly affecting deviations, while shipping charges and zones reflect logistical challenges and potential bottlenecks.
-4.4.2 Probabilistic Multi-Class Classification
-Since deviations measured in days only take on discrete values in the dataset, they inherently follow a discrete probability distribution.
-Thus, the random variable $\tilde{d}$ has a finite discrete support.
-Let $C$ be the total number of possible values $\tilde{d}$ can take, and let $\xi_{c}$ be the $c$-th possible value for $c\in[C]$.
-By definition, $\tilde{d}\in\{\xi_{1},\xi_{2},\ldots,\xi_{C}\}$.
-This discrete nature allows the conditional distribution learning problem to be framed as a multi-class classification (MC-CLF) problem.
+5.2 Probabilistic Multi-Class Classification
+This discrete nature allows the contextual distribution learning problem to be framed as a multi-class classification (MC-CLF) problem.
 Under MC-CLF, each class represents a possible delivery deviation value in number of days, and the output class probabilities correspond directly to the probability mass function of the discrete distribution.
 For a new sample with covariates vector $\mathbf{s}$, let the conditional probability of class $c$ be $p_{c}(\mathbf{s})=\mathbb{P}(d=\xi_{c}\mid\mathbf{s})$, with the corresponding predicted probability from MC-CLF denoted as $\hat{p}_{c}(\mathbf{s})$.
-Multinomial logistic regression (MLR) and classification trees-based models are considered to solve MC-CLF.
-Specifically, MLR is a classic classifier that aims to estimate the class probabilities by maximizing the log-likelihood function, which is the negative of the log loss function.
+Multinomial logistic regression (MLR) and classification tree-based models are considered to solve MC-CLF.
+Specifically, MLR is a classic classifier that aims to estimate the class probabilities by maximizing the log-likelihood function, the negative of the log loss function.
 The log loss is defined as $-\sum_{m=1}^{M}\sum_{c=1}^{C}\mathbf{1}\{d_{m}=\xi_{c}\}\log(\hat{p}_{c}(\mathbf{s}_{m})).$ On the other hand, classification trees use a nonparametric approach that recursively partitions the data into subsets based on feature values in a top-down manner, resulting in a tree-like model of decisions.
-Let $\mathcal{R}_{1},\ldots,\mathcal{R}_{R}$ be the regions partitioned by a classification tree.
-Given covariates $\mathbf{s}$, the predicted probability for class $c$ in a classification tree can be expressed as:
+Let $\mathcal{R}_{1},\ldots,\mathcal{R}_{R}$ be the regions partitioned by a classification tree. Classification tree estimates
 $\hat{p}_{c}(\mathbf{s})=\frac{\sum_{r=1}^{R}\mathbf{1}\{\mathbf{s}\in\mathcal{R}_{r}\}\cdot\mathbf{1}\{d=\xi_{c}\}}{\sum_{r=1}^{R}\mathbf{1}\{\mathbf{s}\in\mathcal{R}_{r}\}},$ (18)
@@
 The details for their calculation of the class probabilities are available in Appendix.
-Moreover, while the proposed classification models excel at predicting class labels, their predicted probabilities may not be well calibrated (Niculescu-Mizil and Caruana 2005).
-Therefore, probability calibration is necessary to obtain accurate conditional distributional forecasts.
-To address this, a cross-validation approach is adopted to fit an isotonic regression calibrator (Zadrozny and Elkan 2002).
-In each split, the classification model is first trained on the training set, and then its predictions on the calibration set are used to fit an isotonic regression model.
-Let $\hat{\tilde{p}}(\cdot)$ be the vector of calibrated probabilities and $\hat{p}(\cdot)$ be the vector of uncalibrated probabilities for each class.
-Isotonic regression takes the vector of uncalibrated probabilities as input and solves the following problem to determine a vector of monotonically increasing calibrated probabilities:
-$\min_{\hat{\tilde{p}}(\cdot)}$ $\sum_{m=1}^{M}(d_{m}-\hat{\tilde{p}}(\mathbf{s}_{m}))^{2}$ (20)
-s.t. $\hat{\tilde{p}}(\mathbf{s}_{m})\geq\hat{\tilde{p}}(\mathbf{s}_{m^{\prime}})\quad\forall m^{\prime}\neq m:\hat{p}(\mathbf{s}_{m})\geq\hat{p}(\mathbf{s}_{m^{\prime}})$
-To measure the accuracy of probabilistic predictions for MC-CLF model, the paper adopts the log loss and the Brier score.
+The performance of the MC-CLF models are evaluated by log loss and Brier score.
 The latter is defined as the average squared error between predicted probabilities and actual outcomes $\sum_{m=1}^{M}\sum_{c=1}^{C}(\mathbf{1}\{d_{m}=\xi_{c}\}\hat{p}_{c}(\mathbf{s}_{m}))^{2}$.
 For both metrics, lower values indicate better probabilistic prediction performance.
-4.4.3 Quantile Regression
-Unlike multi-class classification, quantile regression (QR) can directly estimate the cumulative distribution function (CDF) of the conditional distribution without assuming a discrete probability distribution.
-The CDF of the uncertain deviation $\tilde{d}$ is given by $F(u\mid\mathbf{s})=\mathbb{P}(d\leq u\mid\mathbf{s})$.
-Standard QR models aim at minimizing the pinball loss function, estimating different quantiles independently.
-To approximate the conditional distribution, these models typically calculate a series of conditional quantiles, forming a discretized approximation of the CDF.
-However, this approach can lead to the quantile crossing issue, where predicted quantiles do not maintain their proper order, resulting in an invalid distribution.
-Various methods exist to avoid quantile crossing, but tree-based quantile regression models have been shown to outperform standard QR models in distributional forecasting, particularly in delivery time performance settings (Salari et al.
-2022).
-These tree-based methods do not suffer from the issue of quantile crossing.
-Therefore, this paper focuses on a regression tree-based model and a quantile random forest (QRF) model which leverages an ensemble of regression trees.
-Similar to a classification tree, a regression tree partitions the feature space into multiple disjoint regions, but the split criterion is based on the mean squared error (MSE).
-While regression trees are typically used for generating conditional point predictions, they can be extended to provide distributional forecasts by preserving all observations within each partitioned region, rather than just summary statistics like the mean.
-Let $\mathcal{R}_{1},\ldots,\mathcal{R}_{R}$ be the regions partitioned by a regression tree.
-Given covariates $\mathbf{s}$, the region that $\mathbf{s}$ falls into is denoted as $\mathcal{R}(\mathbf{s})$.
-The conditional distribution of the new sample can be approximated by the empirical CDF:
+5.2.1 Probability Calibration
+Raw class‐probability estimates $\hat{p}_{c}(\mathbf{s})$ from classifiers often exhibit mis‐calibration: a predicted probability $\alpha$ does not correspond to an empirical frequency $\alpha$ (Niculescu-Mizil and Caruana 2005).
+Good calibration means that, over all samples with $\hat{p}_{c}(\mathbf{s})=\alpha$, the true fraction belonging to class $c$ is approximately $\alpha$.
+Formally, let $f(\alpha)=\mathbb{P}\bigl{(}d=\xi_{c}\mid\hat{p}_{c}(\mathbf{s})=\alpha\bigr{)}$ be the calibration function.
+Perfect calibration means $f(\alpha)=\alpha$ for all $\alpha\in[0,1]$.
+Isotonic regression (Zadrozny and Elkan 2002) learns a non‐decreasing mapping $g:[0,1]\to[0,1]$ that best aligns raw scores to observed frequencies, preserving the rank order of $\hat{p}_{c}$.
+Given a calibration set $\mathbb{D}_{cal}=\{(\hat{p}_{c}(\mathbf{s}_{m}),\,y_{c,m})\}_{m=1}^{M_{cal}}$ with $y_{c,m}=\mathbf{1}\{d_{m}=\xi_{c}\}$, isotonic regression solves
+$\min_{g}\;\sum_{m=1}^{M_{cal}}\bigl{(}y_{c,m}-g(\hat{p}_{c}(\mathbf{s}_{m}))\bigr{)}^{2}\quad\text{s.t.}\quad g(\alpha)\leq g(\alpha^{\prime})\;\text{whenever}\;\alpha\leq\alpha^{\prime}.$
+The calibrated probabilities are $\tilde{p}_{c}(\mathbf{s})=g(\hat{p}_{c}(\mathbf{s}))$.
+By enforcing monotonicity, isotonic calibration ensures that higher raw scores remain higher after mapping, and empirical frequencies within each score bin match the calibrated values.
+In this paper, a $k$-fold cross‐validation procedure is employed: for each fold $i$, the classifier is trained on its training split, raw scores are obtained on the held‐out calibration split, and an isotonic regression model $g^{(i)}$ is fit to those scores.
+On unseen test instance with covariates $\mathbf{s}$, each fold’s classifier and corresponding $g^{(i)}$ yield calibrated probabilities $\tilde{p}_{c}^{(i)}(\mathbf{s})$, which are then averaged across all $k$ folds to produce the final estimate $\tilde{p}_{c}(\mathbf{s})=\frac{1}{k}\sum_{i=1}^{k}\tilde{p}_{c}^{(i)}(\mathbf{s})\,.$
+5.3 Tree-Based Quantile Regression
+Unlike multi‑class classification methods that treat each deviation value as an unordered category, quantile regression (QR) exploits the inherent ordering of deviations by framing the task as ordinal regression and directly estimating the conditional cumulative distribution function (CDF) $F(u\mid\mathbf{s})=\mathbb{P}(d\leq u\mid\mathbf{s})$.
+Standard QR models minimizes, for each quantile level $q$, the pinball loss
+$\rho_{q}(d,\hat{d})\;=\;q\,\max(d-\hat{d},0)\;+\;(1-q)\,\max(\hat{d}-d,0),$ (19)
+where $\hat{d}$ denotes the model’s predicted $q$‑th quantile.
+Collecting these quantile estimates across $q$ yields a stepwise approximation of the CDF.
+However, estimating each quantile independently can violate the ordering $\hat{d}_{(q_{1})}\leq\hat{d}_{(q_{2})}$ for $q_{1}<q_{2}$, resulting in a non‑monotonic—and thus invalid—CDF.
+Tree‐based QR methods—namely regression trees extended to CDF estimation, and their ensemble version, quantile random forests (QRF) (Meinshausen and Ridgeway 2006)—avoid this “quantile crossing” problem by extracting quantiles from the empirical distribution of training sample in each leaf region.
+A regression tree partitions the covariate space into disjoint regions to minimize mean squared error (MSE).
+Although regression trees are often used for point forecasts (e.g., predicting the mean response in each leaf), they can be adapted for distributional forecasting by preserving all observations within each partitioned region.
+For a new sample with covariates $\mathbf{s}$, let $\mathcal{R}(\mathbf{s})$ be the leaf region into which $\mathbf{s}$ falls.
+The conditional CDF of the deviation can then be approximated by the empirical distribution of the training samples in that leaf:
 $\hat{F}(u\mid\mathbf{s})=\sum_{m=1}^{M}\frac{\mathbf{1}\{\mathbf{s}_{m}\in\mathcal{R}(\mathbf{s})\}}{|\{j:\mathbf{s}_{j}\in\mathcal{R}(\mathbf{s})\}|}\mathbf{1}\{d_{m}\leq u\}.$ (20)
-To refine this empirical CDF, a kernel density estimation method is applied to the observations in $\mathcal{R}(\mathbf{s})$ (Salari et al. 2022).
-A similar approach can be extended to random forests, and this idea of QRF was first introduced by Meinshausen and Ridgeway (2006).
-In QRF, for a new sample with covariates $\mathbf{s}$, the empirical distribution is derived from all regions associated with $\mathbf{s}$ across all trees in the forest.
-Let $T$ be the total number of regression trees, and let $\mathcal{R}_{1}^{(t)},\ldots,\mathcal{R}_{R}^{(t)}$ be the partitioned regions of the $t$-th tree, where $t\in[T]$.
-The region that $\mathbf{s}$ falls into in the $t$-th tree is denoted as $\mathcal{R}^{(t)}(\mathbf{s})$.
-QRF approximates the conditional distribution using the empirical CDF:
+A subsequent kernel density estimation applied to $\{d_{m}:\mathbf{s}_{m}\in\mathcal{R}(\mathbf{s})\}$ can further smooth this step‐function CDF (Salari et al. 2022).
+QRF extends this approach by pooling the leaf‐level samples across an ensemble of $T$ trees.
+For each tree $t$, let $\mathcal{R}^{(t)}(\mathbf{s})$ be leaf region containing $\mathbf{s}.$ The QRF estimate of the conditional CDF is
 $\hat{F}(u\mid\mathbf{s})=\frac{1}{T}\sum_{t=1}^{T}\sum_{m=1}^{M}\frac{\mathbf{1}\{\mathbf{s}_{m}\in\mathcal{R}^{(t)}(\mathbf{s})\}}{|\{j:\mathbf{s}_{j}\in\mathcal{R}^{(t)}(\mathbf{s})\}|}\mathbf{1}\{d_{m}\leq u\}.$ (21)
-To measure the distributional forecasting performance of different learning models, the paper uses two metrics—the Continuous Ranked Probability Score (CRPS) and pinball loss.
-CRPS evaluates the closeness of the estimated CDF $\hat{F}$ given covariates $\mathbf{s}$ to the observed ground truth value $d$:
+Because this procedure pools ordered observations directly, it naturally preserves the ordinal support and never produces crossing quantiles.
+To measure the distributional forecasting performance, the paper uses two metrics—the Continuous Ranked Probability Score (CRPS) and pinball loss.
+CRPS evaluates the closeness of the estimated CDF $\hat{F}$ given covariates $\mathbf{s}$ to the observed ground truth value $d$
 […]
-Let $\xi$ and $\xi^{\prime}$ be two independent random variables with distribution $\hat{F}(u\mid\mathbf{s})$. The CRPS can alternatively be expressed as:
+Let $\xi$ and $\xi^{\prime}$ be two independent random variables with distribution $\hat{F}(u\mid\mathbf{s})$. The CRPS can alternatively be expressed as
 $\text{CRPS}(\hat{F},d,\mathbf{s})=\mathbb{E}(|\xi-d|)-\frac{1}{2}\mathbb{E}(|\xi-\xi^{\prime}|).$ (23)
@@
 Pinball loss measures the accuracy of quantile predictions by penalizing errors asymmetrically, depending on whether the predictions are overestimates or underestimates.
-Let $q$ be a specific quantile.
-Then, the pinball loss between the observed value $d$ and the predicted value $\hat{d}$ is defined as:
-$\text{pinball}(d,\hat{d})=q\max\{d-\hat{d},0\}+(1-q)\max\{\hat{d}-d,0\}.$ (25)
-4.4.4 Applying the C-SAA and the C-RO
-It is impractical to solve the CSOFP exactly with the learning models just proposed.
-Indeed, on the one hand, QR models learn continuous distributions, making it computationally intractable to expand the expectation directly.
-On the other hand, MC-CLF models leverage the fact that the uncertain delivery deviation vector $\tilde{\mathbf{d}}$ has a finite discrete support, which theoretically allows for expansion of the expectation term.
-However, MC-CLF models face the curse of dimensionality.
-The number of possible values of $\tilde{\mathbf{d}}$ grow exponentially in $K$ and $L$.
-This justifies the methodologies proposed in Section 3.
-In the following, the C-SAA problems using samples from the conditional distribution estimates provided by the MC-CLF and QR models are referred to as C-SAA-CLF and C-SAA-QR, respectively.
-Similarly, the C-RO problems are referred to as C-RO-{B,D}-CLF and C-RO-{B,D}-QR, respectively.
 6 Case Study
@@
 6.1 Data Overview
-The industrial partner provides a month-long dataset in 2023, covering more than 1 million online orders for home delivery and over 20,000 unique stock-keeping units (SKUs).
-Each order is organized by SKUs, resulting in one or more order lines.
-For an incoming order, there are, on average, 250 to 1,000 possible fulfillment centers across the company’s network to source SKUs from, including mostly physical stores but also regional distribution centers and metro e-commerce centers.
-Each location is equipped with multiple eligible carriers that are capable of providing different levels of service (LOS), such as Ground, Same-Day, and Overnight shipping.
-The set of available carriers span from major shipping companies to crowdsourced gig carriers.
-There are 10 to 30 carrier-LOS pairs appeared in the dataset and throughout the paper, “carrier” refers to one of these specific carrier-LOS pairs.
-Each row of the dataset contains transactional details for a single order line, along with the status of the fulfillment network at the time the order was placed.
-The order information includes the date and time of order placement, the type, dimensions, weight, and quantity of the SKUs, the fulfillment location from which each order line was sourced, the carrier that shipped the order, the customer location, the promised delivery date, and the actual arrival date.
-The network status data includes the on-hand inventory levels of the requested SKU and the remaining open capacity at various locations.
-Additionally, the dataset includes carrier-provided shipping costs and planned transit times.
-5.1.1 Learning Results
-For the evaluation of the contextual learning models, the dataset is split into two parts: orders from the first three weeks (1.3 million order lines) are used for training, while orders from the last week (0.4 million order lines) are reserved for testing and evaluation.
-In Table 1, two metrics—log loss and Brier score—are reported to measure the accuracy of probabilistic predictions for the tuned classification models on the test set.
-Probability calibration is applied to the uncalibrated models that obtained the best metrics.
-It is notable that both the log loss and Brier scores show improvement after calibration.
-In addition, Figure 2 illustrates the performance of an example distributional forecast by multi-class classification against its true empirical distribution on the test set.
-The distribution of the majority class 0 (on-time deliveries) is predicted quite accurately, while the distributions of late and early deliveries are forecasted sufficiently close, albeit with small deviations.
-Table 1: Average Log Loss and Brier Score on Test Set
+The industrial partner provides a month-long dataset in 2023, covering more than 1 million online orders for home delivery and over 20,000 unique SKUs.
+Each order can be sourced from roughly 250 to 1,000 possible fulfillment centers across the company’s network and shipped via one of 10 to 30 distinct carriers.
+The target variable in the machine learning problem, delivery deviation, is defined as the difference between the actual delivered date and the promised delivery date.
+To enhance the quality of the analysis, order lines with extreme deviations were excluded—specifically, those with absolute values greater than 10 days.
+Such outliers often reflect exceptional circumstances, including returns to sender, lost shipments, or atypical business orders, which can introduce noise and bias into the model.
+6.2 Evaluation of Contextual Distributional Learning of Delivery Deviations
+This section evaluates the contextual distributional learning models presented in Section 5 for the case study.
+6.2.1 Machine Learning Model Settings
+For the evaluation of the learning models, the dataset was partitioned into a training set comprising orders from the first three weeks (approximately 1.3 million order lines) and a test set consisting of orders from the final week (approximately 0.4 million order lines).
+For multi-SKU orders, features were aggregated by retaining only those of the SKU with the highest shipping charge and quantity.
+The MC-CLF models were implemented using scikit-learn (Pedregosa et al.
+2011) and CatBoost (Prokhorenkova et al.
+2018), with probability calibration applied to the best‐performing uncalibrated models based on cross‐validated metrics.
+Quantile regression models, including regression tree and QRF, were implemented via the quantile-forest package (Johnson 2024).
+All hyperparameters were tuned using three‑fold cross‑validation in Optuna (Akiba et al.
+2019): MC‑CLF models were tuned to minimize log loss, whereas quantile regression models were tuned to minimize CRPS.
+6.2.2 Learning Performance Analysis
+In Table 1, two standard metrics—log loss and Brier score—are reported to measure the accuracy of probabilistic predictions for each tuned MC-CLF model on the test set.
+Calibration yields a clear improvement: the calibrated classifier achieves the lowest log loss (0.965) and Brier score (0.031), down from 1.050 and 0.034 for the uncalibrated gradient‑boosted tree (GBT‑CLF), and substantially outperforming both the random‑forest classifier (RF‑CLF; 1.593, 0.054) and the simpler baselines (MLR and classification tree).
+Figure 2 presents the joint predicted distributions of (a) the calibrated‑CLF, (b) the QRF, and (c) the discretized regression tree, each plotted against the true empirical distribution for orders fulfilled by a representative carrier.
+The calibrated‑CLF closely reproduces the empirical mass at zero deviation (on‑time deliveries) and tightly follows the tails corresponding to early and late deliveries.
+The QRF exhibits slightly greater discrepancies, particularly in the majority classes (-1 to 1), while the discretized regression tree shows the largest misalignment from the empirical distribution.
+Table 1: Out-of-Sample Average Log Loss and Brier Score Across All Carriers.
 Model MLR Classification Tree RF-CLF GBT-CLF Calibrated-CLF
@@
 Brier score 0.073 0.072 0.054 0.034 0.031
-Figure 2: Comparison of Calibrated Predicted Joint Distribution with True Empirical Distribution for an Example Carrier on Test Set.
-Table 2 presents the pinball losses for different quantiles and the CRPS values of the tuned quantile regression models and classification models.
-The CRPS values are approximated by drawing 1,000 samples to compute Equation (24).
-It is observed that the MC-CLF models exhibit higher pinball losses compared to the two quantile regression models, likely due to the discrete nature of classification outputs.
-However, the CRPS of MC-CLF is the lowest among the three models, despite not being explicitly tuned for CRPS.
-Additionally, Figure 3 displays the 95% prediction intervals created by the three conditional distribution learning methods for an example carrier.
+(a) Calibrated-CLF
+(b) QRF
+(c) Regression Tree (discretized)
+Figure 2: Out‐of‐Sample Comparison of the Joint Predicted Distribution and True Empirical Distribution for an Example Carrier.
+Table 2 reports pinball losses at selected quantiles and CRPS values for the tuned quantile regression and calibrated-CLF models.
+CRPS was approximated by drawing 1,000 samples to compute Equation (23).
+The calibrated-CLF model exhibit higher pinball losses than the two quantile regression models, reflecting its inherently discrete output, but nonetheless achieves the lowest CRPS overall, even though it was not directly tuned for this metric.
+This superior CRPS performance likely results from calibration producing sharper probability estimates, especially for the majority deviation classes.
+Additionally, Figure 3 displays the 95% prediction intervals created by the three contextual distribution learning methods for an example carrier.
 The visualization aligns with the pinball loss results, showing that QRF generally provides the narrowest interval widths.
-While the intervals generated by MC-CLF cover most of the observed values, their coverage is less effective at the two extremes compared to the two quantile regression methods.
-The inferior performance of MC-CLF at extreme quantiles is likely due to the discretization of the output space, which limits precision, and the cumulative probability method, which can cause jumps when selecting quantiles.
+While the intervals generated by calibrated-CLF cover most of the observed values, their coverage is less effective at the two extremes compared to the two QR methods.
+The inferior performance of calibrated-CLF at extreme quantiles is likely due to the discretization of the output space, which limits precision, and the cumulative probability method, which can cause jumps when selecting quantiles.
 In contrast, regression tree and QRF provide finer-grained predictions, leading to more accurate and narrower intervals at the extremes.
-In summary, MC-CLF models excel at providing distributional forecasts, whereas QR models are better at generating prediction intervals for this dataset.
-Table 2: Average Pinball Loss and CRPS on Test Set Across All Carriers.
+In summary, MC-CLF models excel at providing distributional forecasts, whereas QR models are better suited to generating prediction intervals on this dataset.
+Since QRF benefits from ensembling multiple regression trees to reduce variance and improve accuracy and consistently outperform a single regression tree, only QRF is kept for QR in the subsequent CSO evaluations.
+Table 2: Out-of-Sample Average Pinball Loss and CRPS Across All Carriers.
 Pinball Loss CRPS
 Model $q_{0.025}$ $q_{0.05}$ $q_{0.5}$ $q_{0.95}$ $q_{0.975}$
-MC-CLF 0.199 0.198 0.676 0.526 0.304 0.386
+Calibrated-CLF 0.199 0.198 0.676 0.526 0.304 0.386
+QRF 0.033 0.057 0.263 0.143 0.096 0.402
 Regression Tree 0.064 0.109 0.302 0.147 0.093 0.515
-QRF 0.033 0.057 0.263 0.143 0.096 0.402
-(a) Regression Tree Prediction Intervals
-(b) MC-CLF Prediction Intervals
-(c) QRF Prediction Intervals
-Figure 3: Sorted Observed Deviation Values versus 95% Prediction Intervals for an Example Carrier.
+(a) Calibrated-CLF Prediction Intervals
+(b) QRF Prediction Intervals
+(c) Regression Tree Prediction Intervals
+Figure 3: Sorted Observed Deviation Values versus 95% Prediction Intervals for an Example Carrier on Out‑of‑Sample Data.
 6.3 Baseline Models
-To compare with the proposed contextual optimization models, three baseline order fulfillment algorithms are introduced below.
+To compare with the proposed contextual optimization models, four baseline order-fulfillment algorithms are introduced below.
   1. 1.
-    Pure Cost-Driven Greedy Heuristic (Greedy): This greedy heuristic replicates the current practice employed by the industrial partner’s sourcing engine.
-The algorithm selects the location-carrier pair with the lowest fulfillment cost for each quantity of an SKU.
-For order $o$, the greedy problem is
-    $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\{\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell o}^{ship})z_{ik\ell o}\}.$
+    Pure Cost-Driven Greedy Heuristic (Greedy): Mimics the industrial partner’s sourcing engine by selecting, for each unit of an SKU, the carrier–location pair that minimizes the fulfillment cost.
+Formally, for each order it solves $\min_{\mathbf{z}\in\mathcal{Z}}\{\sum_{i,k,\ell}(c_{\ell}^{fixed}+c_{ik\ell}^{ship})z_{ik\ell}\}.$
   2. 2.
-    Empirical-SAA: This method follows the approach described in Section 3.4 for C-SAA, with the key difference being that the deviation vector samples are drawn from the empirical distribution of deviations in the training set, without incorporating any contextual information.
-For each sample $n\in[N]$, let $\xi_{k\ell o}^{n}$ denote the $n$-th devaition value sampled from the empirical distribution $P^{E}_{k\ell}$ for carrier-location pair $(k,\ell)$.
-The corresponding deviation vector for sample $n$ is then expressed as $\bm{\xi}_{o}^{n}=(\xi_{k\ell o}^{n})_{k\in\mathcal{K},\ell\in\mathcal{L}}$.
-For order $o$, the empirical SAA problem is
-    $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\{\frac{1}{N}\sum_{n=1}^{N}g(\mathbf{z}_{o},\bm{\xi}_{o}^{n})\}.$
+    Empirical‑SAA: Follows the SAA framework of Section 4.3, except that each deviation sample is drawn from the overall empirical distribution $P^{E}$ of the training set, without using any contextual information.
+For each $n\in[N_{1}]$ and every $(k,\ell)\in\mathcal{K}\times\mathcal{L}$, ${}\xi_{k\ell}^{n}\sim P^{E},~{}\bm{\xi}^{n}=\bigl{(}\xi_{k\ell}^{n}\bigr{)}_{k\in\mathcal{K},\,\ell\in\mathcal{L}}.$ The empirical SAA problem is $\min_{\bm{z}\in\mathcal{Z}}\;\frac{1}{N_{1}}\sum_{n=1}^{N_{1}}g\bigl{(}\bm{z},\bm{\xi}^{n}\bigr{)}.$
   3. 3.
-    Point Predict, Then Optimize (Point-Pred.-Opt.): Unlike the distributional forecasting models mentioned in Section 4.4, typical regression-based ML models aim at estimates only the conditional mean $\mathbb{E}[\tilde{d}_{k\ell}\mid\mathbf{s}_{k\ell o}]$.
-In this method, conditional mean regression models $h_{k\ell}$ are trained to map covariates to deviation for each $(k,\ell)\in\mathcal{K}\times\mathcal{L}$.
-This approach transforms CSOFP into a deterministic optimization problem by directly replacing the random parameter $\tilde{\mathbf{d}}_{o}$ with its point prediction $\hat{\mathbf{d}}_{o}=\{h_{k\ell}(\mathbf{s}_{k\ell o})\}_{k\in\mathcal{K},\ell\in\mathcal{L}}$ in the objective function of Problem (18).
-This method is effective only if the objective function $g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})$ is linear in $\tilde{\mathbf{d}}_{o}$, in which case estimating the conditional distribution simplifies to estimating the conditional mean.
-This is because, by linearity of expectation, $\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})]=g(\mathbf{z}_{o},\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})}[\tilde{\mathbf{d}}_{o}])$.
-However, this linearity does not hold in the proposed objective function due to the asymmetric penalty costs associated with the deviations.
-So a point prediction may be biased, potentially leading to suboptimal solutions.
-To facilitate the experiments, several different regression models are trained to minimize MSE, and the model with the smallest MSE in cross-validation is selected for each carrier.
-See Appendix for the average MSE of different regression models.
-For order $o$ and point prediction $\hat{\mathbf{d}}_{o}$, the optimization model is
-    $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}g(\mathbf{z}_{o},\hat{\mathbf{d}}_{o}).$
-5.3 The Experimental Setting
-The optimization models in this study were implemented using Gurobi 11.0.0 and Python 3.9.
+    C-Empirical-SAA: A slightly smarter SAA baseline that maintains separate empirical distributions $P^{E}_{k}$ for each carrier.
+For each $n$ and $(k,\ell)$, $\xi_{k\ell}^{n}\sim P^{E}_{k},~{}\bm{\xi}^{n}=\bigl{(}\xi_{k\ell}^{n}\bigr{)}_{k\in\mathcal{K},\,\ell\in\mathcal{L}},$ and it solves the same SAA problem as above.
+  4. 4.
+    Point Predict–Then–Optimize (PTO): Trains per-carrier regression models $h_{k}(\cdot)$ to estimate the conditional mean deviations for each $(k,\ell)\in\mathcal{K}\times\mathcal{L}$, producing the point‐prediction vector $\hat{\mathbf{d}}=\{h_{k}(\mathbf{s}_{k\ell})\}_{k\in\mathcal{K},\ell\in\mathcal{L}}$.
+It then solves the deterministic problem $\min_{\bm{z}\in\mathcal{Z}}g\bigl{(}\bm{z},\hat{\mathbf{d}}\bigr{)}.$
+    This method is effective only if the objective function $g(\mathbf{z},\tilde{\mathbf{d}})$ is linear in $\tilde{\mathbf{d}}$, in which case estimating the conditional distribution simplifies to estimating the conditional mean.
+This is because, by linearity of expectation, $\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}})]=g(\mathbf{z},\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s})}[\tilde{\mathbf{d}}])$.
+However, $g(\bm{z},\tilde{\mathbf{d}})$ is nonlinear due to asymmetric penalty costs of the deviations, hence replacing the distribution by its mean may bias the solution.
+For each carrier, the regression model with the lowest cross‐validated mean squared error is selected (see Appendix for details).
+6.4 Experimental Setting
+This section outlines the computational environment, simulation framework, and parameter configurations used in the experiments.
+6.4.1 Implementation Environment
+The optimization models in this study were implemented in Python 3.9 using Gurobi 11.0.0, with eight threads allocated to each order instance.
+A termination criterion of a 1% MIP gap or a 5-minute time limit was applied.
 All experiments were conducted on a server equipped with dual-socket Intel Xeon Gold 6226 CPUs, each featuring 24 cores running at 2.7 GHz.
-The regression and multi-class classification models were developed using the scikit-learn (Pedregosa et al.
-2011), XGBoost (Chen and Guestrin 2016), and CatBoost libraries (Prokhorenkova et al.
-2018).
-For quantile regression, the regression tree and quantile random forest models were implemented using the quantile-forest package (Johnson 2024).
-For the following experiments, all ML models are trained on orders from the first three weeks, while 5,000 multi-item orders randomly sampled from the last week are used to evaluate the order fulfillment methods.
-During both training and inference, for multi-SKU orders, features are aggregated by retaining only the features of the SKU with the highest shipping charges and quantities.
+6.4.2 Simulation Framework
 One of the difficulties in this application is that the ground truth is not available for options that were not selected.
-To overcome this limitation, a simulation environment was built and generate the “ground truth” when an algorithm selects an option.
-The delivery deviations are drawn from a multinomial distribution whose probabilities were estimated by fitting a fine-tuned, calibrated multi-class classification model based on orders in the last week of the dataset.
-Obviously, this data is not available to the algorithm and only used by the simulator to replicate real-world conditions.
-By default, a constant $\beta_{k}=0.5$ discount factor for item batching is considered since the exact values are not provided.
-Sensitivity analysis are perform on the impact of this factor.
-The default penalty for early delivery $\gamma^{-}$ is $0.1$ and that for late delivery $\gamma^{+}$ is $10$.
-The penalty can be understood as the opportunity cost (lost sale cost) for early (late) deliveries, and they are also analyzed with a range of different values.
-For methods relying on SAA, the default is to have $Q=20$ independent replications each with a sample size of $N_{1}=30$.
-The obtained solutions are then evaluated with a default sample size of $N_{2}=300$.
-The default sample size used for C-RO-D is $N=200$.
-As for C-RO-B, a default of $95\%$ interval is considered.
-Moreover, a scaling factor $\eta\in[0,1]$ (default of 0.5) determines the uncertainty budget, defined as $B=\eta(KL).$
+To overcome this limitation, a simulation environment was built to generate synthetic “ground truth” when an algorithm selects an option.
+Specifically, delivery deviations were drawn from a multinomial distribution, with probabilities estimated by fitting a fine-tuned, calibrated multi-class classification model based on orders from the last week of the dataset (testing period).
+Obviously, this data was not accessible to the algorithm and only used by the simulator to replicate real-world conditions.
+The training and calibration procedures of the simulator closely followed the methodology outlined in Section 5.2.
+An evaluation of the simulator is provided in Appendix.
+For the following experiments, as mentioned in Section 6.2, all ML models were trained on order data from the first three weeks.
+Evaluation was conducted on 5,000 multi-item orders, randomly sampled from the final week.
+To account for variability, solution metrics were computed across 50 random realizations.
+All reported metrics were normalized by setting the Greedy baseline to 1 for each random seed; true performance values were omitted to due to data confidentiality.
+6.4.3 Model Names and Parameters
+In the following, the C-SAA problems using samples from the contextual distribution estimates provided by the MC-CLF and QRF models are referred to as C-SAA-CLF and C-SAA-QRF, respectively.
+Similarly, the C-RO problems are referred to as C-RO-{B,D}-CLF and C-RO-{B,D}-QRF, respectively.
+By default, a constant $\beta_{k}=0.5$ consolidation discount factor for all $k\in\mathcal{K}$ was considered.
+The default penalty for early delivery $\gamma^{-}$ was $0.1$ and that for late delivery $\gamma^{+}$ was $10$.
+For SAA-based methods, the default was to have $Q=20$ independent replications each with a sample size of $N_{1}=30$.
+The obtained solutions were then evaluated with a default sample size of $N_{2}=300$.
+The default sample size used for C-RO-D was $N=200$.
+As for C-RO-B, a default of $95\%$ interval was considered.
+Moreover, a scaling factor $\eta\in[0,1]$ (default of 0.5) determined the uncertainty budget in C-RO-B, defined as $B=\eta\,|\mathcal{K}|\,|\mathcal{L}|.$
+6.5 Overall Results
+Table 3 reports the normalized average and worst‐case objective values across 50 random seeds and all order instances. Based on these results, the following observations are drawn.
+Average Performance
+The C-SAA variants and C-RO-B-CLF method lead to more reduction in the expected objective than other methods.
+C‑SAA‑CLF achieves the greatest reduction, with a normalized mean of 0.820 $\pm$ 0.003 (–18.0% versus Greedy), followed by C‑SAA‑QRF at 0.844 $\pm$ 0.003 (–15.6%) and C‑RO‑B‑CLF at 0.842 $\pm$ 0.003 (–15.8%).
+By comparison, Empirical‑SAA (0.962 $\pm$ 0.002) and PTO (0.878 $\pm$ 0.003) offer smaller improvements, and C‑Empirical‑SAA (0.870 $\pm$ 0.003) performs only marginally better than PTO.
+Worst-Case Performance
+The C-RO-D methods deliver the strongest guarantees under the worst simulated outcome: C‑RO‑D‑CLF attains 0.779 (–22.1%), C‑RO‑D‑QRF 0.787 (–21.3%), and C‑RO‑B‑CLF 0.811 (–18.9%).
+These three methods uniformly dominate the baseline and other approaches in the worst‐case metric.
+Overall, the C-SAA approaches excel in minimizing the expected objective, while the C‑RO‑D variants excel in worst‐case robustness.
+Notably, C‑RO‑B‑CLF strikes a strong balance between average efficiency and tail‐risk protection.
+Table 3: Average and Worst-Case Realized Objective Values Across Different Order Fulfillment Methods.
+Method Average Objective (95% CI) Worst-Case Objective
+Greedy 1.000 $\pm$ 0.000 1.000
+PTO 0.878 $\pm$ 0.003 0.919
+Empirical-SAA 0.962 $\pm$ 0.002 1.049
+C-Empirical-SAA 0.870 $\pm$ 0.003 0.896
+C-SAA-CLF 0.820 $\pm$ 0.003 0.823
+C-SAA-QRF 0.844 $\pm$ 0.003 0.842
+C-RO-B-CLF 0.842 $\pm$ 0.003 0.811
+C-RO-B-QRF 0.861 $\pm$ 0.003 0.824
+C-RO-D-CLF 0.928 $\pm$ 0.003 0.779
+C-RO-D-QRF 0.914 $\pm$ 0.003 0.787
+  • Note: Top 3 methods for each metric are highlighted in bold.
 6.6 Effectiveness of C-SAA
-This section first compares the performance of the C-SAA method against the three baseline methods.
-Throughout the following sections, only normalized values of the metrics are reported due to the confidentiality of the data.
-5.4.1 Overall Results
-Table 3 compares the realized objective value of Problem (18) for C-SAA and the baseline methods.
-For C-SAA, only the metrics of the best prediction model are recorded.
-It can be observed that C-SAA consistently has the lowest objective value across different late delivery penalties.
-C-SAA maintains a lower objective value on average across all $\gamma^{+}$ levels (approximately 0.80), compared to Point-Pred.-Opt.
-(0.87) and Empirical-SAA (0.83), highlighting its stability and reliability.
-A perfect balance between the fulfillment costs and delivery timeliness control needs to be tailored for the specific business needs.
-On the one hand, a lower fulfillment cost brings mores immediate profits.
-On the other hand, lower delivery timeliness can significantly increase customer satisfaction, thus retaining more loyal customers, benefiting the company in long-term operations.
-In CSOFP, the trade-off between these short-term and long-term profits is controlled by the key parameter $\gamma^{+}$, which determines the penalty of late deliveries in terms of costs.
-The trend indicates that C-SAA’s objective value decreases more sharply compared to other methods as $\gamma^{+}$ increases, making it particularly effective in scenarios where delivery timeliness is crucial.
-C-SAA outperforms Point-Pred.-Opt.
-by 2% (0.92 vs.
-0.94) when $\gamma^{+}=2.5$ and by 14% (0.63 vs.
-0.79) when $\gamma^{+}=40$, indicating that C-SAA’s advantage becomes more pronounced as the penalty increases.
-While the difference between C-SAA and Empirical SAA in terms of objective value is not huge, the subsequent section will further illustrate the performance gaps between these two methods, as Empirical SAA is more sensitive and could potentially perform worse with larger sample sizes.
-Table 3: Simulated Realized Objective Values of C-SAA vs. the Three Baselines
-Method $\gamma^{+}=2.5$ $\gamma^{+}=5$ $\gamma^{+}=10$ $\gamma^{+}=20$ $\gamma^{+}=40$
-Greedy 1.0 (-) 1.0 (-) 1.0 (-) 1.0 (-) 1.0 (-)
-Point-Pred.-Opt. 0.94 $(2.6\times 10^{-4})$ 0.92 $(3.2\times 10^{-3})$ 0.88 $(4.2\times 10^{-3})$ 0.83 $(5.5\times 10^{-3})$ 0.79 $(7.0\times 10^{-3})$
-Empirical-SAA 0.93 $(0.9\times 10^{-4})$ 0.91 $(2.0\times 10^{-3})$ 0.86 $(3.1\times 10^{-3})$ 0.76 $(1.0\times 10^{-2})$ 0.64 $(1.2\times 10^{-2})$
-C-SAA 0.92 $(1.3\times 10^{-4})$ 0.89 $(3.3\times 10^{-3})$ 0.82 $(5.5\times 10^{-3})$ 0.72 $(1.1\times 10^{-2})$ 0.63 $(1.3\times 10^{-2})$
-  • Note: Values represent the normalized mean (standard error) across five different random seeds for the realized deviations simulation.
-The normalization is performed by setting the value for the Greedy method to 1.
+This section further compares the performance of the C-SAA methods with the two Empirical-SAA baselines and the PTO.
 6.6.1 Value of Contextual Information
-To facilitate a closer examination of the performance of these SAA-based models, Figure 4 shows the impact of varying sample sizes.
-On the one hand, larger sample sizes should provide a more accurate approximation of the expected value in Problem (18), assuming the delivery deviation distributions are well-estimated.
+To facilitate a closer examination of the performance of the SAA-based models, Figure 4 shows the impact of varying sample sizes ($N_{1}$).
+On the one hand, larger sample sizes should provide a more accurate approximation of the expected value in Problem (6), assuming the delivery deviation distributions were well-estimated.
 On the other hand, it is important to maintain a reasonably small sample size to ensure computational efficiency.
 For both C-SAA methods, the effect of increasing sample sizes is obvious.
-The objective values are reduced from 0.86 to 0.82 for C-SAA-CLF and 0.87 to 0.84 for C-SAA-QR on average, indicating their increasing alignment with the true risk-neutral objective function.
+The average objectives are reduced from 0.85 to 0.82 for C-SAA-CLF and 0.88 to 0.84 for C-SAA-QRF, indicating their increasing alignment with the true risk-neutral objective function.
 This improvement is particularly evident in their enhanced effectiveness at avoiding cumulative lateness, defined to be the total number of late delivery days across all units in the orders.
-The cumulative lateness decreases by 21% for C-SAA-CLF and by 18% for C-SAA-QR.
-In contrast, Empirical-SAA struggles to improve its performance in both the realized objective value and the cumulative lateness as the sample sizes grows beyond 10.
-This can be attributed to the fact that purely using the empirical distribution is not sufficient to capture the varying delivery deviation distributions that may be dependent upon contextual information unique upon each order.
+The cumulative lateness decreases by 14% for C-SAA-CLF and by 12% for C-SAA-QRF.
+In contrast, both Empirical-SAA and C-Empirical-SAA struggles to improve there performance in both the average realized objective value and the cumulative lateness as the sample size grows.
+This is because relying solely on the empirical distribution fails to capture the variability in delivery deviation distributions, which may depend on contextual information specific to each order.
 A biased distributional forecasts thus guide the model wrongly to create solutions with inferior quality.
-Based on these observations, both conditional distributional forecasting methods proposed in the paper are successfully in this application, demonstrating a stronger performance compared to using empirical distribution for delivery deviation forecasting.
+Based on these observations, both contextual distributional forecasting methods proposed in the paper are successful in this application, demonstrating a stronger performance compared to using empirical distribution for delivery deviation forecasting.
 In summary, the value of contextual information is particularly important in order fulfillment decision-making, where the uncertainty of delivery deviation can vary drastically across different states of the environment.
-(a) Performance in Realized Objective Value.
+(a) Performance in Average Realized Objective Value.
 (b) Performance in Cumulative Lateness.
@@
 6.6.2 Value of Distributional Forecasts
-This section examines the advantage of employing distributional forecasting models over the simpler and more common point prediction model for CSOFP.
-As supported by theory, simply using the conditional mean estimate of deviations to replace the expected value in CSOFP can introduce bias when approximating the true objective value.
-This approach may yield worse solutions compared to employing conditional distributional estimators.
-These performance differences become even more pronounced as late penalties increase, as illustrated in Figure 5.
-With higher late penalties, accurate forecasting of deviations becomes more crucial; the results reveal the inherent bias of Point-Pred.-Opt, leading to lower quality solutions.
-This is not only evident in the realized objective value (Figure 5(a)) but also in the cumulative lateness (Figure 5(b)).
-For example, the cumulative lateness for Point-Pred.-Opt.
-only improves by 7% from $\gamma^{+}=2.5$ to $\gamma^{+}=40$, whereas C-SAA-CLF and C-SAA-QR show significantly greater improvements of approximately 35% and 32%, respectively.
-This demonstrates that the C-SAA methods not only maintain lower cumulative lateness overall but also improve more significantly as the penalty for late deliveries increases, further highlighting the benefits of distributional forecasting.
-(a) Performance in Realized Objective Value.
+This section examines the advantage of employing distributional forecasting models over the simpler and more common point prediction model within the CSO framework.
+As supported by theory, simply using the conditional mean estimate of deviations to replace the expected value in CSOFP can introduce bias when approximating the true objective value, thus degrading solution quality.
+Figure 5 confirms that, under changing $\gamma^{+}$, CSO methods with distributional estimator (C‑SAA‑CLF and C‑SAA‑QRF) consistently outperform PTO in both average realized objective value (Figure 5(a)) and cumulative lateness (Figure 5(b)).
+In fact, the performance differences become more pronounced as late delivery penalty ($\gamma^{+}$) increases.
+When $\gamma^{+}$ raises from 2.5 to 40, PTO achieves only a 7% reduction in cumulative lateness, whereas C-SAA-CLF and C-SAA-QRF realize reductions of approximately 22% and 20%, respectively.
+These results demonstrate that distributional forecasting not only enables lower overall lateness but also leverages higher penalties more effectively, yielding a sharper decline in average objective and making C‑SAA particularly well suited to settings where delivery timeliness is critical.
+(a) Performance in Average Realized Objective Value.
 (b) Performance in Cumulative Lateness.
-Figure 5: C-SAA vs. Point-Pred.-Opt. Across Different Late Delivery Penalties.
-5.5 Evaluation of C-RO
-This section evaluates the performance of the C-RO methods, whose results are shown in Figure 6.
-The C-RO methods generally provide more conservative solutions than C-SAA, with C-RO-B showing a performance closer to C-SAA, as shown by the realized objective values in Figure 6(a).
-Notably, C-RO-B outperforms C-RO-D by offering less conservative yet still robust solutions, while not requiring full distributional information.
-In real-time fulfillment scenarios, where sampling can be time-prohibitive, especially during peak hours, C-RO-B offers a quick and effective alternative, delivering reasonably high-quality solutions.
-Additionally, the effectiveness of C-RO methods in controlling late deliveries remains consistent as the penalty increases (see Figure 6(b)).
-This stability occurs because C-RO methods already account for worst-case scenarios, making them less sensitive to changes in the late delivery penalty, provided that the samples or uncertainty sets encompass these scenarios.
-As a result, C-RO methods outperform C-SAA in reducing cumulative lateness when the late delivery penalty is low but are outperformed by C-SAA when $\gamma^{+}>10$.
-However, the C-RO methods incur higher fulfillment costs—calculated as the total shipping costs and fixed non-parcel costs, minus item consolidation discounts across all orders—by up to 20% on average (see Figure 6(c)) This increase is driven by their risk-averse approach, which prioritizes avoiding extreme deviations by opting for more expensive fulfillment options.
-(a) Performance in Realized Objective Value.
-(b) Performance in Cumulative Lateness.
-(c) Performance in Fulfillment costs.
-Figure 6: Performance of C-RO Methods.
+Figure 5: C-SAA vs. PTO Across Different Late Delivery Penalties.
+6.7 Cost of Robustness
+To quantify the trade‑off between robustness and efficiency, C‑RO‑B‑CLF and C‑RO‑D‑CLF were compared with C‑SAA‑CLF as the late‑delivery penalty ($\gamma^{+}$) increases (Figure 6).
+As depicted in Figure 6(a), at $\gamma^{+}=$ 2.5 , average objectives are approximately 0.92 for C‑SAA‑CLF, 0.95 for C‑RO‑B‑CLF, and 1.00 for C‑RO‑D‑CLF.
+As $\gamma^{+}$ rises to 40, these values decline in parallel—C‑SAA‑CLF to 0.63, C‑RO‑B‑CLF to 0.67, and C‑RO‑D‑CLF to 0.73.
+Throughout this range, C‑SAA‑CLF consistently achieves the lowest average objective, followed by C‑RO‑B‑CLF and then C‑RO‑D‑CLF.
+Worst‑case objectives (Figure 6(b)) underscore the robustness of the C‑RO methods: C‑RO‑D‑CLF consistently yields the strongest tail protection, dropping from 0.88 at $\gamma^{+}=$ 2.5 to 0.72 at $\gamma^{+}=$ 40.
+C‑RO‑B‑CLF falls from 0.885 to 0.775, while C‑SAA‑CLF declines more steeply from 0.925 to 0.73.
+Notably, at $\gamma^{+}=$ 40, C‑SAA‑CLF nearly matches C‑RO‑D‑CLF, demonstrating that C-SAA can approach robust performance under extreme penalty scenarios.
+This robustness manifests in cumulative lateness (Figure 6(c)).
+Both C‑RO variants tightly control lateness with minimal sensitivity to $\gamma^{+}$, whereas C‑SAA‑CLF’s lateness drops substantially from 0.67 at $\gamma^{+}=$ 2.5 to 0.46 at $\gamma^{+}=$ 40.
+These robustness gains incur higher expected fulfillment costs (Figure 6(d)).
+C‑RO‑D‑CLF’s average cost increases from 1.07 to 1.20, and C‑RO‑B‑CLF’s from 1.01 to 1.05, and C‑SAA‑CLF from 0.96 to 1.05.
+Thus, while all methods incur higher costs under stricter penalties, C‑RO‑B‑CLF exhibits the smallest relative cost increase, C‑SAA‑CLF and C‑RO‑D‑CLF have more aggressive increases in costs due to more conservative decisions.
+A final strategic consideration lies in selecting $\gamma^{+}$ to reflect firm priorities.
+Lower $\gamma^{+}$ emphasizes immediate cost savings—maximizing short‑term profits—whereas higher $\gamma^{+}$ prioritizes on‑time delivery, fostering customer satisfaction, loyalty, and long‑term value.
+The CSOFP framework encodes this trade‑off directly via $\gamma^{+}$; robust formulations (C‑RO‑B‑CLF and C‑RO‑D‑CLF) further reduce the need for precise $\gamma^{+}$ tuning by delivering stable, worst‑case performance across penalty levels, while C‑SAA‑CLF can leverage high $\gamma^{+}$ most aggressively when strict timeliness is paramount.
+(a) Performance in Average Realized Objective Value.
+(b) Performance in Worst-Case Realized Objective Value.
+(c) Performance in Cumulative Lateness.
+(d) Performance in Fulfillment costs.
+Figure 6: C-RO vs. C-SAA Across Different Late Delivery Penalties.
 6.7.1 Flexibility of C-RO-B
-The effectiveness of the C-RO models in controlling delivery deviations is evident since they can hedge against the worst-case deviation and potential misspecification of the prediction models.
-However, their robustness to extreme delivery deviations comes at the cost of more conservative solutions, which can result in higher fulfillment expenses.
-Fortunately, the C-RO-B models, beyond their distribution-free property, are equipped with the flexibility to control their level of conservativeness.
-Specifically, they are determined via the coverage of the prediction intervals and the robust uncertainty budget.
-Smaller prediction interval coverage and uncertainty budget may lead to less conservative solutions, making the models more beneficial in terms of addressing forecasting misspecifications rather than purely managing risk.
-Conversely, larger coverage and budget aim at minimizing the impact of worst deviations as much as possible with the sacrifice of overall solution quality.
-The impact of different prediction interval coverages and robust uncertainty budgets on the quality of solutions obtained from C-RO-B is examined in Figures 7 and 8, respectively.
-As the prediction interval coverage increases from 50% to 95% in Figure 7(a) , both C-RO-B-CLF and C-RO-B-QR show a noticeable decline in the realized objective value, with the lowest values around 0.85 for C-RO-B-CLF and 0.86 for C-RO-B-QR at 95% and 90% coverage, respectively.
-This represents a reduction of approximately 6.5% for both C-RO-B methods compared to the objective values at 50% coverage.
-This suggests that increasing the coverage interval initially benefits the overall performance, up to a point.
-This is consistent with the general decline trend in cumulative lateness as prediction interval increases (see Figure 7(b)).
-However, at 99% coverage, there is a sharp increase in the objective value, This trend highlights that while conservative prediction intervals can reduce lateness, they also introduce higher fulfillment costs due to more conservative decisions, thus increasing the overall objective value.
-For C-RO-B-CLF, this increase in objective value is also due to a declined prediction performance with the largest prediction interval, as indicated by the final increase of cumulative lateness at 99% coverage.
-(a) Performance in Realized Objective Value.
-(b) Performance in Cumulative Lateness.
+The C‑RO models effectively hedge against worst‐case delivery deviations and potential misspecifications in the prediction models.
+However, this robustness incurs more conservative decisions and higher fulfillment costs.
+The C‑RO‑B variants retain the distribution‐free guarantees while introducing two tuning parameters—prediction‐interval coverage and robust uncertainty budget—that directly control conservativeness.
+Lower coverage levels and smaller budgets yield less conservative, more cost‐efficient solutions (at the expense of risk protection), whereas higher values emphasize hedging against extreme deviations by sacrificing average performance.
+Figures 7 and 8 summarize how these parameters affect solution quality under both C‑RO‑B‑CLF and C‑RO‑B‑QRF.
+As the prediction interval coverage increases from 70% to 90%.
+In Figure 7(a), increasing coverage from 70% to 90% reduces the average objective to about 0.84 (C‑RO‑B‑CLF) and 0.86 (C‑RO‑B‑QRF), an approximately 4% improvement relative to 70% coverage.
+This is consistent with the concurrent decline in cumulative lateness shown in Figure 7(c).
+However, pushing coverage beyond 95% reverses this trend: at 99% coverage the average objective rises sharply due to the cost of extreme conservatism and, for C‑RO‑B‑CLF, degraded prediction accuracy (reflected in a spike in lateness).
+By contrast, C‑RO‑B‑QRF continues to reduce lateness at 99%, illustrating its superior tail modeling.
+In terms of worst‐case performance (Figure 7(b)), both methods improve with coverage, but C‑RO‑B‑CLF plateaus after 95%, whereas C‑RO‑B‑QRF continues to improve at 99%.
+The effect of the robust budget, controlled by the scaling factor $\eta$, appears in Figure 8.
+Increasing $\eta$ from zero to 0.01 leads to a steep drop in the average objective for both methods (Figure 8(a)).
+Both method reach their minimum values at $\eta=0.01$.
+At the same time, the worst-case objective follows a similar pattern (Figure 8(b)): it falls sharply up to $\eta=0.1$, then edges back up for both methods as the budget grows.
+This trend is also seen in cumulative lateness, which drops dramatically until $\eta=0.1$ and then levels off (Figure 8(c)).
+Together, these trends show that moderate budgets capture most of the gains in both average and worst case performance, while further increases deliver diminishing returns.
+(a) Performance in Avg. Obj.
+(b) Performance in Worst-Case Obj.
+(c) Performance in Cumulative Lateness.
 Figure 7: Performance of C-RO-B Across Different Prediction Interval Coverages.
-A similar pattern is observed for varying robust budgets.
-As $\eta$ increases, the robust budget becomes larger, leading to more conservative solutions.
-For both C-RO-B methods, larger budget results in a more significant decrease in cumulative lateness (see Figure 8(b)), but the results are less clear when it comes to the objective value (see Figure 8(a)).
-The realized objective value decreases sharply as the robust budget scaling factor $\eta$ increases from 0.0 to 0.01 for both C-RO-B-CLF and C-RO-B-QR.
-However, after $\eta=0.01$, there is a noticeable jump in the objective value for C-RO-B-QR, while C-RO-B-CLF continues to stabilize.
-This jump suggests that C-RO-B-QR adopts more conservative decisions with higher robust budgets, which leads to an increase in fulfillment costs.
-Therefore, C-RO-B-QR hits its lowest objective value at $\eta=0.01$ whereas C-RO-B-CLF is able to obtain its minimum objective value at $\eta=1$.
-The analysis reveals a trade-off between increasing robustness and striking a balance between lower fulfillment costs and mitigating extreme deviation risks.
-(a) Performance in Realized Objective Value.
-(b) Performance in Cumulative Lateness.
+(a) Performance in Avg. Obj.
+(b) Performance in Worst-Case Obj.
+(c) Performance in Cumulative Lateness.
 Figure 8: Performance of C-RO-B Across Different Robust Uncertainty Budgets.
@@
 The industrial partner does not know the exact percentage of discount to be offered for item consolidation, which may depend on a variety of convoluted factors.
-Therefore, this section performs sensitivity analysis on a range of possible discount factors for the CSO models.
-The consolidation level, calculated as the percentage of units that are batched (i.e., sourced and shipped by the same location-carrier pair) in all orders, is introduced to measure the effectiveness of the methods in identifying item batching opportunities for cost discounts within each order.
-Figure 9 illustrates the performance of C-SAA, C-RO-B, and C-RO-D under a series of different discount factors.
-As shown in Figure 9(a), C-SAA still consistently exhibits the lowest objective value across all discount factors, with C-RO-B comes next.
-As expected, all three methods exhibit lower objective values due to lower fulfillment costs as the discount factor increases (see Figure 9(b)).
+Therefore, this section performs sensitivity analysis on a range of possible discount factors (from 0.1 to 0.7).
+Figure 9 compares the performance of three leading methods—C-SAA-CLF, C-RO-B-CLF, and C-RO-D-CLF across average and worst-case realized objective, fulfillment cost, and consolidation level.
+Consolidation level is defined as the percentage of units batched (sourced and shipped by the same location–carrier pair) in each order.
+As shown in Figure 9(a), deeper discounts lead each method to exploit batching incentives more aggressively.
+C-SAA-CLF consistently attains the lowest average objective across all discount factors, declining from approximately 0.875 at $\beta=$0.1 to about 0.785 at $\beta=$0.7.
+C‑RO‑B‑CLF follows closely, from 0.895 down to 0.805, while C‑RO‑D‑CLF remains the most conservative, decreasing from 0.955 to 0.905.
+The worst‑case objective exhibits a more differentiating pattern (Figure 9(a)).
+Both C‑RO‑D‑CLF and C‑RO‑B‑CLF are less susceptible to change in the discount, whereas C‑SAA‑CLF starts at 0.842 and falls to 0.802 under high discounts.
+As expected, all three methods exhibit lower average objective values due to lower fulfillment costs as the discount factor increases (see Figure 9(c)).
 This aligns with the intuition that higher discounts encourage more cost-efficient fulfillment strategies.
-The decrease in both C-RO methods can be explained by their higher consolidation levels as the discount factors increases.
-C-RO methods have higher fulfillment costs, so they benefit more from discounts and capitalize on this opportunity.
-The trend in consolidation level for C-SAA is less straightforward.
-While it peaks at a discount factor of 0.25, the consolidation level remains lower than the baseline throughout.
-This suggests that, despite some consolidation opportunities, the C-SAA model does not exploit them to the same extent as the C-RO models.
-One possible explanation is that the benefits of consolidation in C-SAA might not be sufficient to outweigh the penalties associated with deviations.
-The peak at a discount factor of 0.25 could be an anomaly or indicative of specific conditions where consolidation was briefly more favorable, but overall, the model prioritizes minimizing deviation penalties over achieving higher consolidation.
-(a) Performance in Realized Objective Value.
-(b) Performance in Fulfillment Costs.
-(c) Performance in Consolidation Level.
-Figure 9: Performance of CSO Models Across Different Discount Factors.
+The relative decrease in fulfillment costs across the methods can be attributed to their higher consolidation levels as the discount factor increases, as shown in Figure 9(d).
+Notably, C-SAA-CLF responds most sharply, with its consolidation rate rising the most rapidly as the discount grows, which is consistent with its pronounced improvement in worst‑case objective performance.
+(a) Performance in Average Realized Objective Value.
+(b) Performance in Worst-Case Realized Objective Value.
+(c) Performance in Fulfillment Costs.
+(d) Performance in Consolidation Level.
+Figure 9: Performance of CSO Methods Across Different Consolidation Discount Factors.
+6.9 Perturbation Analysis under Varying Simulation Environments
+To assess the robustness of the proposed CSO methods, this section evaluates performance under varying simulation environments by introducing systematic perturbations to the simulated delivery time deviation distributions.
+These perturbations represent moderate to adversarial divergence from the test distribution, providing insight into the methods’ stability under distributional shifts.
+Two types of perturbations are considered: random uniform noise and delay-specific bias applied to the simulation probability distributions.
+Random Uniform Noise Perturbation
+Let the original simulation distribution for each test instance be denoted by $P^{sim}=\{p_{i}\}_{i=1}^{C}$.
+For each class $i\in[C]$, a random noise term $\delta_{i}\sim\mathcal{U}(-\epsilon,\epsilon)$ is added to the probability $p_{i}$, resulting in: $\tilde{p}_{i}=p_{i}+\delta_{i}.$ Negative values are clipped: $\tilde{p}_{i}=\max(\tilde{p}_{i},0)$, and the resulting vector is normalized to form a valid probability distribution: $\tilde{p}_{i}=\tilde{p}_{i}/\sum_{j}\tilde{p}_{j}$.
+Figure 10 reports the average and worst‐case realized objectives as the noise level $\epsilon$ increases.
+In the average case (Figure 10(a)), both C‑RO variants degrade more gradually than C‑SAA, with C‑RO‑B‑CLF becoming the top performer beyond $\epsilon=10$.
+In the worst‐case scenario (Figure 10(b)), while all methods experience performance degrade, the ranking remains unchanged: C‑RO‑D achieves the lowest realized objective, followed by C‑RO‑B, and then C‑SAA.
+These results indicate the superior robustness of the C‑RO methods under uniform noise.
+Delay-Specific Bias Perturbation
+Define the set of late‐delay classes $C_{\mathrm{late}}=\{\,i:\xi_{i}>0\}$.
+A multiplicative bias $\eta$ is applied to those classes: $\tilde{p}_{i}=p_{i}\cdot(1+\eta).$ The probabilities for classes not in $C_{\mathrm{late}}$ remain unchanged.
+The resulting vector is clipped to the interval $[0,1]:$ $\tilde{p}_{i}=\max(\min(\tilde{p}_{i},1),0)$, followed by normalization.
+Figure 11 shows performance metrics for bias levels up to 250%.
+In the average case (Figure 11(a)), all CSO methods demonstrate improved performance relative to the Greedy baseline, which deteriorates rapidly.
+This trend is evidenced by the overall decrease in normalized objective values as the bias parameter $\eta$ increases.
+Notably, C‑RO‑B overtakes C‑SAA once $\eta$ exceeds 150%, highlighting its robustness under adversarial delay distortions.
+In the worst‐case objective (Figure 11(b)), C‑RO‑D consistently achieves the best performance, followed by C‑RO‑B‑CLF.
+Overall, the C‑RO methods exhibit lower variability in performance compared to C‑SAA, indicating enhanced stability under adversarial delay conditions.
+(a) Change in Average Objective.
+(b) Change in Worst-Case Objective.
+Figure 10: Performance Across Different Random Uniform Noise Perturbations.
+(a) Change in Average Objective.
+(b) Change in Worst-Case Objective.
+Figure 11: Performance Across Different Delay-Specific Bias Perturbations.
+6.10 Computational Scalability of the CSO Methods
+The computational scalability of the proposed framework was evaluated by measuring both the change in computation time (model construction + solve time) and solution quality as instance size grows.
+Each CSOFP instance is characterized by the number of items $|\mathcal{I}|$ and the number of location–carrier pairs (fulfillment options) $|\mathcal{K}|\times|\mathcal{L}|$.
+In practice, the eligibility constraint (Constraint (3)) is used to prune infeasible location–carrier pairs in advance.
+Then, $(\mathcal{K}^{\prime},\mathcal{L}^{\prime})\;=\;\bigl{\{}(k,\ell)\in\mathcal{K}\times\mathcal{L}\,\bigm{|}\,\exists\,i\in\mathcal{I}:\;e_{ik\ell}=1\bigr{\}}$ yields the final set of eligible fulfillment options and thus the effective problem size.
+Figure 12(a) shows computation time as the number of eligible fulfillment options increases.
+Only the MC‑CLF variants of C‑RO and C‑SAA are displayed, since their QRF counterparts exhibit nearly identical runtimes.
+All three methods achieve optimality on every instance within minutes, confirming their practical applicability.
+In particular, C‑RO‑B solves most instances in under one second (never exceeding three seconds), and C‑RO‑D solves most in under ten seconds.
+These timing differences reflect each method’s formulation.
+Compared to the nominal problem, the robust counterpart in C‑RO‑B introduces only continuous variables whose count grows linearly with $|\mathcal{K}^{\prime}|$ and $|\mathcal{L}^{\prime}|$.
+By contrast, C‑RO‑D adds $N$ extra constraints from sampling, and C‑SAA requires solving $Q$ independent problems whose objective function complexity scales with the sample size $N_{1}$.
+Parallel computation of the $Q$ subproblems can further reduce C‑SAA’s runtime.
+Since single-item orders account for the majority of order volume, multi-item orders (what the methods are designed for) can be routed to a dedicated sourcing engine without disrupting standard operations.
+Under normal load, any of the three methods may be applied; under peak loads—such as holiday surges—C‑RO‑B is recommended to preserve fast solve times while maintaining high solution quality.
+Concurrent batches of orders can be processed in parallel when computational resources permit.
+Figure 12(b) reports average and worst-case objective values for small ($\leq 400$ eligible options), medium ($\leq 1000$ eligible options), and large ($\leq 3000$ eligible options) instances.
+Average performance remains uniformly high across all size categories, whereas worst-case deviations grow with instance size.
+Collectively, these results demonstrate near-linear scalability in computation time, consistent high solution quality, and robustness across problem sizes, providing empirical evidence of the framework’s practical scalability.
+(a) Change in the Computation Time vs. the Number of Eligible Fulfillment Options.
+(b) Objective Value by Instance Categories.
+Figure 12: Computational Scalability Analysis.
+6.11 Managerial and Operational Implications
+The proposed fulfillment framework offers actionable guidance for practitioners seeking to align sourcing strategies with demand variability, service-level requirements, and computational constraints.
+Depending on the operating context, different methods are preferable.
+When demand is stable and the goal is to minimize average fulfillment cost, C-SAA-CLF provides the most cost-effective solution.
+Under moderate runtime constraints, C-RO-B-CLF offers a balanced trade-off between efficiency and robustness.
+In contrast, during high-stakes periods such as holiday promotions or flash sales—where the cost of service-level violations is high—C-RO-D-CLF and C-RO-B-QRF (with 99% coverage) are better suited to mitigate tail risks and ensure reliability.
+This scenario-driven approach allows e-commerce platforms to dynamically adapt fulfillment strategies to shifting conditions while leveraging a unified optimization infrastructure.
+As retailers increasingly face variable demand patterns and rising customer expectations, the ability to flexibly match fulfillment strategy to context becomes a key operational advantage.
+Empirical evaluations indicate that the proposed methods can reduce combined fulfillment cost and delivery-timeliness penalties by an average of up to 18% per order under the default settings.
+To illustrate the potential impact, consider a retailer operating at a scale comparable to the industrial partner, with per-order costs ranging from \$5 to \$30 and monthly volumes of 0.1 to 1 million multi-item orders.
+Under these scenarios, the corresponding annual savings could range from several million to tens of millions of dollars.
+Moreover, this framework can be readily extended to single-item orders, which could yield further savings with minimal incremental investment.
 7 Conclusion
-This paper presented a data-driven contextual stochastic optimization framework, which was applied to an industrial omnichannel order fulfillment problem within the context of multi-courier logistics.
-Leveraging contextual optimization effectively addresses the inherent complexities of modern retail logistics, providing innovative solutions tailored to the dynamic e-commerce landscape.
-In particular, the proposed framework is capable of handling the availabilities of observational data only.
-To achieve these results, the paper uses probabilistic multi-class classification and quantile regression models to estimate the conditional distribution of delivery time deviations.
-These machine learning models are then seamlessly integrated with SAA-based risk-neutral optimization models and robust optimization models under the proposed CSO framework.
-The paper evaluated the proposed methodology on a real dataset from one of the largest e-commerce company.
-The results, which significantly improve the baseline solution, showcase the ability of the CSO framework to improve service quality, customer satisfaction, and operational efficiency.
-They offer valuable insights and tools for practitioners aiming to optimize their order fulfillment strategies in the evolving e-commerce environment.
-There are several potential directions for future work.
-The contextual optimization models could be extended to incorporate long-term inventory planning through multi-period order fulfillment optimization.
-This would shift the approach from a rather myopic perspective to a more comprehensive strategy that accounts for future inventory needs.
-The distributional forecasts of deviations could be improved by increasing the dataset to include factors such as seasonal trends, weather conditions, and carrier tracking details.
-Furthermore, integrating downstream optimization costs into the training of distributional forecasting models can be an interesting extension (Qi et al.
-2021, Elmachtoub and Grigas 2022).
-However, it is crucial to address the observational nature of the dataset and the varying dimensions of different optimization problem instances.
-Another promising direction is to extend the current contextual optimization framework to an end-to-end optimization approach (Kotary et al.
-2021) while adhering to all operational constraints .
-This paper is relevant to the broader literature on uncertainty quantification, where conformal predictions offer a promising direction.
-Conformal predictions provide coverage guarantees for prediction intervals and could be particularly effective when combined with quantile regression (Romano et al.
-2019).
-Moreover, adopting conformally calibrated predictive distributions (Manokhin 2017) and improving probability calibration methods could enhance model performance.
-The paper hints at this potential by adopting probability calibration in multi-class classification.
-The Venn-Abers calibrator was experimented, but it was too computationally intensive to be fitted for the dataset used in the paper.
-[…]
-This research was partly supported by the NSF AI Institute for Advances in Optimization (Award 2112533).
+This paper introduces a generic contextual stochastic optimization (CSO) framework for data‑driven decision making when only observational data—hence missing counterfactual costs—are available.
+The framework is deliberately modular.
+First, any algorithm capable of returning a full predictive distribution can serve as the contextual distribution oracle.
+In the empirical study, calibrated multi‑class classifiers and tree‑based quantile‑regression forests were selected to capture the discrete, ordered nature of delivery‑time deviations.
+Second, the oracle feeds directly into two tractable optimization engines: (i) a Contextual Sample‑Average Approximation (C‑SAA) for risk‑neutral objectives and (ii) a Contextual Robust Optimization (C‑RO) formulation for risk‑averse objectives.
+Because the learning and optimization layers communicate only through the oracle, the template can be readily translated to other operations‐management settings.
+When applied to an omnichannel, multi‑courier order‑fulfillment problem, the CSO framework reduced the combined cost–service objective by up 18% on average and up to 22% in the simulated worst case on an industry‑scale data set, implying annual savings from several to tens of millions of dollars.
+This framework allows practitioners to tune robustness—raising it during promotional peaks and relaxing it in steadier periods—while still exploiting item‑consolidation discounts and heterogeneous carrier performance in the dynamic e-commerce landscape.
+While the paper delivers an end‑to‑end prototype with demonstrable impact, several avenues remain open:
+  • Performance guarantees under partial feedback: Establish theoretical performance guarantees, such as worst-case regret and post-decision surprise, for CSO methods under the partial‑feedback setting.
+Linking causal‑inference techniques with SAA and RO theory can be a promising direction, thereby quantifying the value of additional data when counterfactual costs remain unobserved.
+  • Hybrid distributional forecasters: Develop distributional models that fuse the calibration strength of classification with the sharp tails of quantile regression, yielding tighter and more informative uncertainty sets.
+  • End‑to‑end predict‑then‑prescribe learning for observational data: Incorporate downstream optimization loss directly into training, extending recent “integrated predict‑then‑optimize” approaches (Qi et al.
+2021, Elmachtoub and Grigas 2022) to settings where data are purely observational.
+  • Multi‑period, inventory‑aware CSO: Extend the framework to jointly optimize sourcing, shipping, and stocking over rolling horizons, capturing inventory replenishment spillovers and elevate the model from a myopic to a long-term strategic planning tool.

### v3

@@
-Contextual Stochastic Optimization for Order Fulfillment Optimization
+MSOM-2024-1328.R2
 […]
+MSOM-2024-1328.R2
+[…]
+Ye et al.
 […]
@@
 […]
-Contextual Stochastic Optimization for Omnichannel Multi-Courier Order Fulfillment Under Delivery Time Uncertainty
+Contextual Stochastic Optimization for
+Omnichannel Multi-Courier Order Fulfillment Under Delivery Time Uncertainty
 […]
@@
 […]
-NSF AI Institute for Advances in Optimization (AI4OPT), Georgia Institute of Technology, Atlanta, GA
+[…]
 […]
@@
 This application demonstrates the framework capabilities and its potential impact on large-scale, real-world optimization challenges.
+Based on the significant performance improvements demonstrated in this study, the industrial partner is actively considering a pilot program to integrate key elements of the proposed framework into their production fulfillment optimization engine.
 The contributions of the paper can be summarized as follows.
   1. 1.
-    Generic CSO framework for observational data: The paper presents a novel, distribution‑agnostic contextual stochastic optimization (CSO) framework that converts problems with missing‑counterfactual observational data into tractable problems.
+    A modular and scalable CSO framework for observational data: The paper presents a novel, distribution‑agnostic contextual stochastic optimization (CSO) framework that converts problems with missing‑counterfactual observational data into tractable problems.
 A contextual distribution oracle is learned from partial observations, and the resulting problem is solved via two scalable solution paradigms—Contextual Sample Average Approximation (C‑SAA) and Contextual Robust Optimization (C‑RO).
+This features a tractable formulation for assignment-type problems with piecewise linear objectives.
   2. 2.
@@
 3.1.1 Overview of the Current Practice
-Currently, the industrial partner uses a sourcing engine that employs a straightforward greedy algorithm.
-For each unit in an order, the sourcing engine sorts the eligible location-carrier pairs by the estimated shipping costs and transit times.
+Currently, the industrial partner uses a fulfillment optimization engine that employs a straightforward greedy algorithm.
+For each unit in an order, the fulfillment optimization engine sorts the eligible location-carrier pairs by the estimated shipping costs and transit times.
 There are several limitations with this approach.
@@
 This section introduces the nominal MILP model developed for optimizing the omnichannel multi-courier order fulfillment problem.
-A comprehensive list of the notations presented in this section is provided in Appendix for reference.
+A comprehensive list of the notations presented in this section is provided in Appendix A for reference.
 3.2.1 Main Decision Variables
@@
 The objective of the optimization problem for an order is to minimize the overall fulfillment costs while ensuring timely delivery.
-The fulfillment costs for an order include the fixed sourcing (non-parcel) costs and the shipping (parcel) costs, with consolidation discounts applied when multiple units are assigned to the same carrier-location pair.
+The fulfillment costs for an order include the fixed fulfillment (non-parcel) costs and the shipping (parcel) costs, with consolidation discounts applied when multiple units are assigned to the same carrier-location pair.
 Delivery timeliness is measured by deviations from the desired delivery date, whether early or late.
@@
 Let $c_{ik\ell}^{ship}>0$ be the per-unit shipping cost when using location $\ell$ and carrier $k$ to fulfill SKU $i$.
-For each location $\ell$, let $c_{\ell}^{fixed}>0$ be the per-unit fixed sourcing cost.
+For each location $\ell$, let $c_{\ell}^{fixed}>0$ be the per-unit fixed fulfillment cost.
 For each carrier $k$, let $\beta_{k}\in(0,1)$ be the discount percentage applied to shipping costs when multiple units are shipped using carrier $k$ from the same fulfillment center.
@@
 The weighted objective function for an order can be constructed as follows:
-$g(\mathbf{z},\mathbf{d})=\begin{cases}\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}[c_{\ell}^{fixed}+(1-\beta_{k})c_{ik\ell}^{ship}+\gamma^{+}d_{k\ell}^{+}+\gamma^{-}d_{k\ell}^{-}]z_{ik\ell}&\text{if }\sum_{i\in\mathcal{I}}z_{ik\ell}\geq 2,\\ \sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell}^{ship}+\gamma^{+}d_{k\ell}^{+}+\gamma^{-}d_{k\ell}^{-})z_{ik\ell}&\text{otherwise},\end{cases}$ (1)
+[…]
 where $d_{k\ell}^{+}=\max\{0,d_{k\ell}\}$ and $d_{k\ell}^{-}=\max\{0,-d_{k\ell}\}$.
@@
 This leads to the following contextual stochastic order fulfillment problem (CSOFP):
-(CSOFP) $\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s})}[g(\mathbf{z},\tilde{\mathbf{d}})].$ (6)
+[…]
 Solving Problem (6) in practice requires leveraging data from past orders.
@@
 The resulting dataset available for estimating the contextual distribution $\mathbb{P}(\mathbf{d}\mid\mathbf{s})$ is $\mathbb{D}=\{(\mathbf{s}_{o},\mathbf{D}_{o})\}_{o=1}^{|\mathcal{O}|}$.
-4 Methodology: A Generic CSO Framework for Observational Data
+4 Methodology: A CSO Framework for Observational Data
 Solving CSOFP under observational data requires: (1) learning a contextual distributional oracle from partial realizations, and (2) optimizing with that oracle in a tractable way.
-This section lays out a generic contextual stochastic optimization (CSO) framework that does exactly that.
-In Section 4.1, the abstract CSO problem (denoted by ${\cal P}$) is defined.
+This section lays out a contextual stochastic optimization (CSO) framework that does exactly that.
+In Section 4.1, a generic CSO problem (denoted by $\cal P$}) is defined.
 Section 4.2 shows how to construct a contextual distribution oracle $\mathcal{M}(\mathbf{s})$ from partial observations.
@@
 Finally, Section 4.6 maps all of the above back to the omnichannel multi-courier CSOFP introduced in Section 3.
+Figure 2 provides a schematic overview of the proposed framework, illustrating the flow from historical observational data through the learning and optimization stages to produce a final decision.
+Figure 2: A Schematic Overview of the CSO Framework for Observational Data.
 4.1 Generic Contextual Stochastic Optimization (CSO) Problem
 Problem $\cal P$} consists of fulfilling a set of incoming requests sequentially, each of which specifies demands for a subset of products.
-Let $\mathcal{I}$ be the set of products, $\mathcal{V}$ the set of selection options (e.g., sourcing or transport modes), and $\mathcal{Z}\subseteq\mathbb{Z}_{+}^{|\mathcal{I}|\times|\mathcal{V}|}$ the feasible assignment set.
+Let $\mathcal{I}$ be the set of products, $\mathcal{V}$ the set of selection options (e.g., fulfillment or transport modes), and $\mathcal{Z}\subseteq\mathbb{Z}_{+}^{|\mathcal{I}|\times|\mathcal{V}|}$ the feasible assignment set.
 Each request arrives with known demand quantities $\{q_{i}:i\in\mathcal{I}\}$ at decision time.
 The decision variable $\mathbf{z}=(z_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}\in\mathcal{Z}$ assigns quantity $z_{iv}$ of product $i$ to option $v$.
-Let $\tilde{\mathbf{c}}=(\tilde{c}_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}\sim\mathbb{P}$ denote the random cost vector.
-Assume that $\tilde{\mathbf{c}}$ is element-wise positive and the cost function $g(\mathbf{z},\tilde{\mathbf{c}})$ is piecewise linear in $\mathbf{z}$.
-Introduce, for $j=0,1,\dots,n$, the breakpoint vectors
 […]
+$\mathbf{b}^{j}=\bigl(b^{j}_{iv}\bigr)_{i\in\mathcal{I},\;v\in\mathcal{V}},~\mathbf{b}^{0}<\mathbf{b}^{1}<\cdots<\mathbf{b}^{n}\quad(\text{component-wise}),$
 so that each segment is defined as
-$[\mathbf{b}^{j-1},\,\mathbf{b}^{j}]=\bigl{\{}\mathbf{z}:b^{\,j-1}_{iv}\leq z_{iv}\leq b^{j}_{iv}\;\forall\,i\in\mathcal{I},\;v\in\mathcal{V}\bigr{\}}.$
+$\mathbf{b}^{j-1},\,\mathbf{b}^{j}]=\bigl\{\mathbf{z}:b^{\,j-1}_{iv}\leq z_{iv}\leq b^{j}_{iv}\;\forall\,i\in\mathcal{I},\;v\in\mathcal{V}\bigr\}.$
 Let $\mathbf{a}^{j}=(a^{j}_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}$ be the marginal‐cost adjustment on segment $j$. Then
@@
 For each request, the fulfillment decision solves the risk‐neutral stochastic optimization problem
-$\min_{\mathbf{z}\in\mathcal{Z}}\;\mathbb{E}_{\tilde{\mathbf{c}}\sim\mathbb{P}}\bigl{[}g(\mathbf{z},\tilde{\mathbf{c}})\bigr{]}.$ (8)
+[…]
 The piecewise‐linear cost in (7) partitions the decision space into shipment‐size tiers determined by the breakpoints $\{\mathbf{b}^{j}\}$.
@@
 The distribution $\mathbb{P}$ is then approximated by the conditional distribution $\mathbb{P}(\mathbf{c}\mid\mathbf{s})$, yielding the Contextual Stochastic Optimization (CSO) problem
-$\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s})}[g(\mathbf{z},\mathbf{\tilde{c}})].$ (9)
+[…]
 Let $\tilde{\mathbf{c}}_{v}=(\tilde{c}_{iv})_{i\in\mathcal{I}}$ and $\mathbf{s}_{v}=(\mathbf{s}_{iv})_{i\in\mathcal{I}}$ for each $v\in\mathcal{V}$.
@@
 Consequently, Problem (9) can be written as
-$\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{c}}\sim\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{v})}[g(\mathbf{z},\mathbf{\tilde{c}})].$ (10)
+[…]
 4.2 Constructing a Contextual Distribution Oracle from Observational Data
@@
 The robust counterpart of (11) becomes
-[…]
+$\text{(C-RO-D)}\quad\min_{\mathbf{z}\in\mathcal{Z},\;t\in\mathbb{R}}$ $t$ (12)
 s.t. $t\geq g(\mathbf{z},\bm{\xi}^{n}),\quad n=1,\dots,N.$
@@
 A valid choice for the “big‑M” constant is $M\;=\;\max_{i,v,j}\bigl\{\,b_{iv}^{n}-b_{iv}^{1},\;b_{iv}^{\,n-1},\;(\underline{\xi}_{iv}+a_{iv}^{j})\,b_{iv}^{n}\bigr\}.$ By taking the dual of the inner maximization problem and introducing $y_{j}$ and $w_{iv}^{j}$, a robust counterpart of (11) is obtained as
-[…]
+$\text{(C-RO-B)}\quad\min_{\mathbf{z}\in\mathcal{Z},y,w,\pi,\lambda}$ $(\sum_{i,v,j}w_{iv}^{j}+\pi B+\sum_{i,v}\lambda_{iv})$ (16)
 s.t. $\pi+\lambda_{iv}\geq(\overline{\xi}_{iv}-\underline{\xi}_{iv})z_{iv},\quad\forall i,v$
@@
 provided
-[…]
+$N\;\geq\;2|\mathcal{V}||\mathcal{I}|\;+\;\frac{2|\mathcal{V}||\mathcal{I}|}{\epsilon}\ln\!\Bigl(\frac{2}{\epsilon}\Bigr)\;+\;\frac{2}{\epsilon}\ln\!\Bigl(\frac{1}{\delta}\Bigr)$
 samples are drawn (Calafiore and Campi 2006, Campi and Garatti 2008, Bertsimas et al.
@@
 4.6 Application to the Omnichannel Multi-Courier CSOFP
-The generic CSO framework applies directly to CSOFP via the following identifications:
+The CSO framework applies directly to CSOFP via the following identifications:
 $\mathcal{I}\;\mapsto\;\{\text{SKUs}\},$ $\mathcal{V}\;\mapsto\;\mathcal{K}\times\mathcal{L},$
@@
 Hence CSOFP (Problem 6) can be rewritten as
-$\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{d}}\sim\Pi_{k,\ell}\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell})}[g(\mathbf{z},\tilde{\mathbf{d}})].$ (17)
-To learn each marginal distribution $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell})$, partition the historical dataset into
-$\mathbb{D}_{k\ell}=\bigl{\{}(\mathbf{s}_{k\ell,m},\,d_{k\ell,m})\bigr{\}}_{m=1}^{M_{k\ell}},$
-and train a per‑carrier model $\mathcal{M}_{k}$ on $\mathbb{D}_{k}=\bigcup_{\ell\in\mathcal{L}}\mathbb{D}_{k\ell}$.
-The full contextual distribution oracle is $\mathcal{M}(\mathbf{s})=\bigl{(}\mathcal{M}_{k}(\mathbf{s}_{k\ell})\bigr{)}_{k\in\mathcal{K},\,\ell\in\mathcal{L}}.$
-This per‑carrier decomposition simplifies learning, generalizes to new locations, and outperforms a single global model according to preliminary experiments.
+[…]
+To learn each marginal distribution $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell})$, the historical dataset is first partitioned into $\mathbb{D}_{k\ell}=\bigl\{(\mathbf{s}_{k\ell,m},\,d_{k\ell,m})\bigr\}_{m=1}^{M_{k\ell}}.$ This is followed by training a per‑carrier model $\mathcal{M}_{k}$ on the dataset $\mathbb{D}_{k}=\bigcup_{\ell\in\mathcal{L}}\mathbb{D}_{k\ell}$.
+This per-carrier approach is a practical adaptation of the generic per-option framework from Section 4.2.
+Rather than training a distinct model for each individual location-carrier pair (which can be inefficient and suffer from data sparsity), the data is pooled to train a single, more robust model for each carrier.
+This decomposition simplifies learning, allows the model to generalize to new locations, and was found to outperform a single global model in preliminary experiments.
+The full contextual distribution oracle is then assembled as $\mathcal{M}(\mathbf{s})=\bigl(\mathcal{M}_{k}(\mathbf{s}_{k\ell})\bigr)_{k\in\mathcal{K},\,\ell\in\mathcal{L}}.$
 4.6.1 Applying C-SAA and C-RO
-Computing the expectation in $\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathcal{M}(\mathbf{s})}\bigl{[}g(\mathbf{z},\tilde{\mathbf{d}})\bigr{]}$ directly would require either enumerating all $C^{|\mathcal{K}|\times|\mathcal{L}|}$ deviation vectors or numerically integrating a continuous CDF—both computationally intractable.
-Instead, the two solution paradigms use the contextual distribution oracle $\mathcal{M}(\mathbf{s})$ as follows:
+[…]
   • C-SAA replaces the expectation by a sample average: draw $N_{1}$ full‑vector scenarios $\{\tilde{\mathbf{d}}^{n}\}\sim\mathcal{M}(\mathbf{s})$ and solve
@@
     $\min_{\mathbf{z}\in\mathcal{Z}}\;\max_{\mathbf{d}\in\mathcal{U}(\mathbf{s})}g(\mathbf{z},\mathbf{d}).$
-Both approaches avoid the exponential support growth while remaining agnostic to the internal structure of $\mathcal{M}$. A full C‑RO‑B reformulation is provided in the Appendix.
+Both approaches avoid the exponential support growth while remaining agnostic to the internal structure of $\mathcal{M}$. A full C‑RO‑B reformulation is provided in the Appendix F.
 5 Learning the Contextual Distribution of Delivery Deviations for CSOFP
@@
 Specifically, MLR is a classic classifier that aims to estimate the class probabilities by maximizing the log-likelihood function, the negative of the log loss function.
-The log loss is defined as $-\sum_{m=1}^{M}\sum_{c=1}^{C}\mathbf{1}\{d_{m}=\xi_{c}\}\log(\hat{p}_{c}(\mathbf{s}_{m})).$ On the other hand, classification trees use a nonparametric approach that recursively partitions the data into subsets based on feature values in a top-down manner, resulting in a tree-like model of decisions.
+The log loss is defined as $\sum_{m=1}^{M}\sum_{c=1}^{C}\mathbf{1}\{d_{m}=\xi_{c}\}\log(\hat{p}_{c}(\mathbf{s}_{m})).$ On the other hand, classification trees use a nonparametric approach that recursively partitions the data into subsets based on feature values in a top-down manner, resulting in a tree-like model of decisions.
 Let $\mathcal{R}_{1},\ldots,\mathcal{R}_{R}$ be the regions partitioned by a classification tree. Classification tree estimates
@@
 To enhance predictive power, ensemble learning methods such as bagging and gradient boosting are employed, through the use of random forests (RF) and gradient boosted trees (GBT).
-The details for their calculation of the class probabilities are available in Appendix.
+The details for their calculation of the class probabilities are available in Appendix G.
 The performance of the MC-CLF models are evaluated by log loss and Brier score.
@@
 In this paper, a $k$-fold cross‐validation procedure is employed: for each fold $i$, the classifier is trained on its training split, raw scores are obtained on the held‐out calibration split, and an isotonic regression model $g^{(i)}$ is fit to those scores.
-On unseen test instance with covariates $\mathbf{s}$, each fold’s classifier and corresponding $g^{(i)}$ yield calibrated probabilities $\tilde{p}_{c}^{(i)}(\mathbf{s})$, which are then averaged across all $k$ folds to produce the final estimate $\tilde{p}_{c}(\mathbf{s})=\frac{1}{k}\sum_{i=1}^{k}\tilde{p}_{c}^{(i)}(\mathbf{s})\,.$
+On unseen test instance with covariates $\mathbf{s}$, each fold’s classifier and corresponding $g^{(i)}$ yield calibrated probabilities $\tilde{p}_{c}^{(i)}(\mathbf{s})$, which are then averaged across all $k$ folds to produce the final estimate $\tilde{p}_{c}(\mathbf{s})=\frac{1}{k}\sum_{i=1}^{k}\tilde{p}_{c}^{(i)}(\mathbf{s})\,.$ Since per-class calibration does not guarantee the final probabilities will sum to one, a normalization step is applied.
+The vector of calibrated class probabilities $\bigl(\tilde{p}_{c}(\mathbf{s})\bigr)_{c=1}^{C}$ is divided by its sum to yield a valid multi-class probability distribution.
 5.3 Tree-Based Quantile Regression
@@
 $\rho_{q}(d,\hat{d})\;=\;q\,\max(d-\hat{d},0)\;+\;(1-q)\,\max(\hat{d}-d,0),$ (19)
-where $\hat{d}$ denotes the model’s predicted $q$‑th quantile.
+where $\hat{d}$ denotes the model’s predicted qq‑th quantile.
 Collecting these quantile estimates across $q$ yields a stepwise approximation of the CDF.
@@
 CRPS evaluates the closeness of the estimated CDF $\hat{F}$ given covariates $\mathbf{s}$ to the observed ground truth value $d$
-[…]
+$\text{CRPS}(\hat{F},d,\mathbf{s})=\int_{-\infty}^{\infty}(\hat{F}(u\mid\mathbf{s})-\mathbf{1}\{u\geq d\})^{2}du.$ (22)
 Let $\xi$ and $\xi^{\prime}$ be two independent random variables with distribution $\hat{F}(u\mid\mathbf{s})$. The CRPS can alternatively be expressed as
@@
   1. 1.
-    Pure Cost-Driven Greedy Heuristic (Greedy): Mimics the industrial partner’s sourcing engine by selecting, for each unit of an SKU, the carrier–location pair that minimizes the fulfillment cost.
+    Pure Cost-Driven Greedy Heuristic (Greedy): Mimics the industrial partner’s fulfillment optimization engine by selecting, for each unit of an SKU, the carrier–location pair that minimizes the fulfillment cost.
 Formally, for each order it solves $\min_{\mathbf{z}\in\mathcal{Z}}\{\sum_{i,k,\ell}(c_{\ell}^{fixed}+c_{ik\ell}^{ship})z_{ik\ell}\}.$
@@
   4. 4.
-    Point Predict–Then–Optimize (PTO): Trains per-carrier regression models $h_{k}(\cdot)$ to estimate the conditional mean deviations for each $(k,\ell)\in\mathcal{K}\times\mathcal{L}$, producing the point‐prediction vector $\hat{\mathbf{d}}=\{h_{k}(\mathbf{s}_{k\ell})\}_{k\in\mathcal{K},\ell\in\mathcal{L}}$.
-It then solves the deterministic problem $\min_{\bm{z}\in\mathcal{Z}}g\bigl{(}\bm{z},\hat{\mathbf{d}}\bigr{)}.$
-    This method is effective only if the objective function $g(\mathbf{z},\tilde{\mathbf{d}})$ is linear in $\tilde{\mathbf{d}}$, in which case estimating the conditional distribution simplifies to estimating the conditional mean.
-This is because, by linearity of expectation, $\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}})]=g(\mathbf{z},\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s})}[\tilde{\mathbf{d}}])$.
-However, $g(\bm{z},\tilde{\mathbf{d}})$ is nonlinear due to asymmetric penalty costs of the deviations, hence replacing the distribution by its mean may bias the solution.
-For each carrier, the regression model with the lowest cross‐validated mean squared error is selected (see Appendix for details).
+[…]
+[…]
 6.4 Experimental Setting
@@
 The training and calibration procedures of the simulator closely followed the methodology outlined in Section 5.2.
-An evaluation of the simulator is provided in Appendix.
+An evaluation of the simulator is provided in Appendix J.
 For the following experiments, as mentioned in Section 6.2, all ML models were trained on order data from the first three weeks.
@@
 Lower coverage levels and smaller budgets yield less conservative, more cost‐efficient solutions (at the expense of risk protection), whereas higher values emphasize hedging against extreme deviations by sacrificing average performance.
-Figures 7 and 8 summarize how these parameters affect solution quality under both C‑RO‑B‑CLF and C‑RO‑B‑QRF.
+Figures 8 and 9 summarize how these parameters affect solution quality under both C‑RO‑B‑CLF and C‑RO‑B‑QRF.
 As the prediction interval coverage increases from 70% to 90%.
@@
 6.9 Perturbation Analysis under Varying Simulation Environments
-To assess the robustness of the proposed CSO methods, this section evaluates performance under varying simulation environments by introducing systematic perturbations to the simulated delivery time deviation distributions.
-These perturbations represent moderate to adversarial divergence from the test distribution, providing insight into the methods’ stability under distributional shifts.
-Two types of perturbations are considered: random uniform noise and delay-specific bias applied to the simulation probability distributions.
+A critical challenge in any simulation-based evaluation is that the simulator itself is an imperfect model of the real-world data-generating process.
+This is particularly true in this application, where the purely observational nature of the historical data, and the resulting lack of counterfactual outcomes, make it impossible to construct a perfect, clairvoyant simulator.
+Therefore, to rigorously assess the stability of the proposed solutions, this section performs a stress test designed to probe the framework’s performance under a misspecified ground truth.
+This analysis is conducted by introducing systematic perturbations to the simulated delivery time deviation distributions, representing moderate to adversarial divergences from the test data.
+Specifically, two types of perturbations are considered: random uniform noise and a delay-specific bias applied to the simulation probability distributions.
+Together, these tests provide critical insight into the methods’ stability under real-world distributional shifts and their robustness to the inevitable simulator-reality mismatch.
 Random Uniform Noise Perturbation
@@
 Parallel computation of the $Q$ subproblems can further reduce C‑SAA’s runtime.
-Since single-item orders account for the majority of order volume, multi-item orders (what the methods are designed for) can be routed to a dedicated sourcing engine without disrupting standard operations.
+For completeness, a comparison of computation times against the baseline models is provided (see Appendix K).
+The Greedy and PTO methods are computationally cheaper, solving most instances in fractions of a second.
+However, the proposed CSO methods, particularly the highly-efficient C-RO-B, remain well within practical time limits for real-world deployment (Figure 13(a)).
+The substantial improvements in solution quality (Table 3) justify this modest increase in computational overhead, representing a favorable trade-off between performance and speed.
+Since single-item orders account for the majority of order volume, multi-item orders (what the methods are designed for) can be routed to a dedicated fulfillment optimization engine without disrupting standard operations.
 Under normal load, any of the three methods may be applied; under peak loads—such as holiday surges—C‑RO‑B is recommended to preserve fast solve times while maintaining high solution quality.
@@
 6.11 Managerial and Operational Implications
-The proposed fulfillment framework offers actionable guidance for practitioners seeking to align sourcing strategies with demand variability, service-level requirements, and computational constraints.
+The proposed fulfillment framework offers actionable guidance for practitioners seeking to align fulfillment strategies with demand variability, service-level requirements, and computational constraints.
 Depending on the operating context, different methods are preferable.
@@
 7 Conclusion
-This paper introduces a generic contextual stochastic optimization (CSO) framework for data‑driven decision making when only observational data—hence missing counterfactual costs—are available.
+This paper introduces a generic contextual stochastic optimization (CSO) framework for data‑driven decision making when only observational data, hence missing counterfactual costs, are available.
 The framework is deliberately modular.
@@
 2021, Elmachtoub and Grigas 2022) to settings where data are purely observational.
-  • Multi‑period, inventory‑aware CSO: Extend the framework to jointly optimize sourcing, shipping, and stocking over rolling horizons, capturing inventory replenishment spillovers and elevate the model from a myopic to a long-term strategic planning tool.
+  • Multi‑period, inventory‑aware CSO: Extend the framework to jointly optimize fulfillment, shipping, and stocking over rolling horizons, capturing inventory replenishment spillovers and elevate the model from a myopic to a long-term strategic planning tool.
+  • Generalization to broader problem classes: Extending the framework to broader settings like network flow optimization is a key direction for future work.
+The sampling-based C-SAA and C-RO-D approaches can be directly generalized.
+However, the C-RO-B formulation’s tractability relies on the specific problem structure, and adapting it to more complex constraints would require new reformulations.
+{APPENDIX}
+8 MILP Model Notations
+Table 4 presents the notations used in the MILP optimization model.
+Table 4: Optimization Model Nomenclature
+Sets
+$\mathcal{I}$ Set of SKUs, indexed by $i$.
+$\mathcal{K}$ Set of carriers, indexed by $k$.
+$\mathcal{L}$ Set of locations, indexed by $\ell$.
+Parameters
+$\operatorname{inv}_{i\ell}$ Inventory level of SKU $i$ at location $\ell$ when the order is placed.
+$\operatorname{cap}_{\ell}$ Available capacity at location $\ell$ when the order is placed.
+$e_{ik\ell}$ 1 if (carrier $k$, location $\ell$) is an eligible candidate location-carrier pair for SKU $i$ in the order and 0 otherwise.
+$c_{ik\ell}^{ship}$ Per-unit shipping cost when using location $\ell$ and carrier $k$ to ship SKU $i$.
+$c_{\ell}^{fixed}$ Per-unit fixed sourcing (non-parcel) cost when using location $l$.
+$\beta_{k}$ Discount factor for shipping multiple units using carrier $k$.
+$q_{i}$ Quantity of SKU $i$ in the order.
+$d_{k\ell}$ Delivery time deviation when using location $\ell$ and carrier $k$.
+$\gamma^{+}$ Constant that converts late delivery penalty into a per-unit cost.
+$\gamma^{-}$ Constant that converts early delivery penalty into a per-unit cost.
+NN A large number = $\sum_{i\in\mathcal{I}}q_{i}+2$.
+$M_{k\ell}$ A large number = $\sum_{i\in\mathcal{I}}c_{ik\ell}^{ship}q_{i}$.
+Decision Variables
+$z_{ik\ell}$ Integer decision variable which measures the quantity of SKU $i$ in the order sourced from location $\ell$ and shipped by carrier $k$.
+$y_{k\ell}$ Binary decision variable which equals 1 if more than one unit is sourced from location $\ell$ and carrier $k$ within the order, and equals 0 otherwise.
+$w_{k\ell}$ Continuous decision variable which measures the cost to be discounted in the order using location-carrier pair $(k,\ell)$.
+9 Data Preprocessing
+To optimize the dataset for machine learning and optimization models, several preprocessing steps were implemented to clean the data.
+Duplicate entries and rows with missing values were removed.
+Geographic coordinates were generated from the zip codes of the fulfillment centers and customer locations, with any invalid zip codes discarded in the process.
+The geographic distance between customer and fulfillment center locations was approximated using the Haversine formula.
+Additionally, the dataset was filtered to include only carriers with sufficient order history for training prediction models, resulting in a subset of carriers.
+10 Descriptive Analytics and Insights
+The company receives tens of thousands of home delivery online orders daily, with an average ranging from 30,000 to 70,000 orders.
+During peak hours, 500 to 2,000 orders may be received within a 15-minute time window.
+Direct information regarding which carriers are available to ship specific SKUs from particular fulfillment centers is not available in the dataset.
+However, heuristic rules can be derived through analyses of historical orders.
+The customer’s location and the dimensions/weight of the SKU(s) in an order are critical factors in determining carrier suitability.
+Some carriers show a more scattered distribution of customer locations, indicating limited coverage in certain states of the U.S..
+Additionally, coverage is dependent on the location of the fulfillment center from which the order is shipped, as some carriers can only serve a limited range around that center.
+In contrast, other carriers have served locations spanning nearly the entire US.
+Furthermore, the upper bounds of SKU length, height, and weight in the historical data reflect the package size and weight limits of different carriers.
+Figure 14: Delivery Deviation Distribution of Orders Shipped by Three Example Carriers.
+The distribution of deviations varies significantly across carriers, as illustrated in Figure 14.
+For the first two carriers, the majority of orders are delivered on time.
+The first carrier shows nearly all on-time deliveries, with the remaining orders arriving late.
+The second carrier has approximately half of the orders delivered on time, with slightly more early deliveries than late ones.
+In contrast, orders shipped by the last carrier predominantly arrive one day early, and most deliveries arrive ahead of schedule.
+Given these pronounced differences, separate prediction models were developed for each carrier.
+11 Feature Importance
+To identify the most important features for predicting deviations, a feature importance analysis was conducted using a tree-based regression model.
+Two methods were examined: Mean Decrease in Impurity (MDI) and Shapley values (SHAP).
+MDI measures the importance of a feature by calculating the total reduction in impurity (e.g., Gini impurity or entropy) it brings when used in decision trees (Breiman 2001).
+SHAP values, on the other hand, provide a unified measure of feature importance based on cooperative game theory, attributing to each feature its contribution to the prediction by considering all possible feature combinations (Lundberg et al.
+2020).
+Figure 15 highlights the top 10 most important features evaluated, as evaluated using MDI and SHAP, for an example carrier under a tree-based regression model.
+For this carrier, the planned lead time (promised delivery time) emerges as the most significant factor.
+Additionally, order information such as the release time and SKU type, fulfillment center characteristics like on-hand capacity, and carrier-specific attributes such as promised time-in-transit and carrier zone also play important roles.
+(a) MDI Feature Importance.
+(b) SHAP Value Feature Importance.
+Figure 15: Top 10 Mean Regression Features for Delivery Deviation Measured by Two Feature Importance Metrics For an Example Carrier.
+12 Linearize the Objective Function of the Order Fulfillment Problem
+The nonlinear objective function $g(\mathbf{z},\mathbf{d})$ can be linearlized using auxiliary variables.
+Let $\mathbf{y}=(y_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}\in\{0,1\}^{(|\mathcal{K}||\mathcal{L}|)}$ be indicator variables, where $y_{k\ell}=1$ indicates more than one unit is sourced from location $\ell$ and shipped by carrier $k$ within the order.
+Additionally, let $\mathbf{w}=(w_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}\in\mathbb{R}^{(|\mathcal{K}||\mathcal{L}|)+}$ be continuous variables, where $w_{k\ell}$ represents the shipping costs to be discounted in the order using carrier-location pair ($k,\ell$).
+The linear objective function and the additional constraints for linearization are defined as follows:
+$g(\mathbf{z},\mathbf{d})=$ $(\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell}^{ship}+\gamma^{+}d_{k\ell}^{+}+\gamma^{-}d_{k\ell}^{-})z_{ik\ell}-\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\beta_{k}w_{k\ell})$ (24)
+$\begin{aligned} \text{s.t.}\quad&\sum_{i\in\mathcal{I}}z_{ik\ell}\geq 2-N(1-y_{k\ell})\\ &\sum_{i\in\mathcal{I}}z_{ik\ell}\leq 1+Ny_{k\ell}\\ &w_{k\ell}\leq M_{k\ell}y_{k\ell}\\ &w_{k\ell}\leq\sum_{i\in\mathcal{I}}c_{ik\ell}^{ship}z_{ik\ell}+(1-y_{k\ell})M_{k\ell}\\ &w_{k\ell}\geq\sum_{i\in\mathcal{I}}c_{ik\ell}^{ship}z_{ik\ell}-(1-y_{k\ell})M_{k\ell}\\ &y_{k\ell}\in\{0,1\}\\ &w_{k\ell}\geq 0\\ \end{aligned}\quad,\forall{k\in\mathcal{K}},{\ell\in\mathcal{L}},\\$ (25)
+where $N$ and $M_{k\ell}$ are large constants for the big-M constraints, with their strictest values being $\sum_{i\in\mathcal{I}}q_{i}+2$ and $\sum_{i\in\mathcal{I}}c_{ik\ell}^{ship}q_{i}$, respectively.
+13 Applying the C-RO-B to CSOFP
+[…]
+Substituting $\tilde{\mathbf{c}}$ leads to:
+$\min_{\mathbf{z}\in\mathcal{Z}}\max_{\tilde{\mathbf{c}}\in\mathcal{U}_{b}(\mathbf{s})}$ $(\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\tilde{c}_{ik\ell}z_{ik\ell}-\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\beta_{k}w_{k\ell}).$ (26)
+Given that the prediction interval for each $\tilde{d}_{k\ell}$ is $\underline{d}_{k\ell},\underline{d}_{k\ell}+\hat{d}_{k\ell}]$, where $\underline{d}_{k\ell}$ and ($\underline{d}_{k\ell}+\hat{d}_{k\ell}$) are the estimated lower and upper quantiles of the deviation for the carrier-location pair $(k,\ell)$, the prediction interval for $\tilde{\mathbf{c}}$ can be expressed as:
+[…]
+Let $B\in[0,(|\mathcal{K}||\mathcal{L}|)]$ be the uncertainty budget. The contextual budgeted interval uncertainty set $\mathcal{U}_{b}(\mathbf{s})$ is expressed as follows:
+$\mathcal{U}_{b}(\mathbf{s},B)=\{\bm{\xi}\in\mathbb{R}^{(|\mathcal{I}||\mathcal{K}||\mathcal{L}|)}:\xi_{ik\ell}=\underline{\xi}_{ik\ell}+\hat{\xi}_{ik\ell}\delta_{k\ell},\sum_{k\in\mathcal{K},\ell\in\mathcal{L}}\delta_{k\ell}\leq B,\bm{\delta}\in[0,1]^{(|\mathcal{K}||\mathcal{L}|)}\}.$ (28)
+Then, with fixed $\mathbf{z}$, the inner maximization problem of Problem (26) becomes:
+$\max\quad$ $\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\hat{\xi}_{ik\ell}z_{ik\ell}\delta_{k\ell}$ (29)
+s.t. $\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\delta_{k\ell}\leq B$ (30)
+$\delta_{k\ell}\in[0,1],\quad\forall k\in\mathcal{K},\ell\in\mathcal{L}.$ (31)
+The robust counterpart of Problem (26) is formulated as follows:
+$\min_{\mathbf{z}\in\mathcal{Z}}$ $(\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\underline{\xi}_{ik\ell}z_{ik\ell}-\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\beta_{k}w_{k\ell}+\pi B+\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\lambda_{k\ell})$ (32)
+s.t. $\pi+\lambda_{k\ell}\geq\hat{\xi}_{ik\ell}z_{ik\ell},\quad\forall i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}$
+$\pi\geq 0$
+$\lambda_{k\ell}\geq 0,\quad\forall k\in\mathcal{K},\ell\in\mathcal{L},$
+where $\pi$ and $(\lambda_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ are the dual variables of Constraints (30) and (31), respectively.
+14 Probabilistic Multi-Class Classification Using Ensemble Learning
+RF leverages multiple classification trees trained on different random subsets of the data, making predictions by averaging the results (Breiman 2001).
+Let $\mathcal{R}_{1}^{(t)},\ldots,\mathcal{R}_{R}^{(t)}$ correspond to the partitioned regions of the $t$-th tree.
+Given covariates $\mathbf{s}$, the predicted probability for class $c$ in an RF classifier is:
+$\hat{p}_{c}(\mathbf{s})=\frac{1}{T}\sum_{t=1}^{T}\frac{\sum_{r=1}^{R}\mathbf{1}\{\mathbf{s}\in\mathcal{R}_{r}^{(t)}\}\cdot\mathbf{1}\{d=\xi_{c}\}}{\sum_{r=1}^{R}\mathbf{1}\{\mathbf{s}\in\mathcal{R}_{r}^{(t)}\}}.$ (33)
+In contrast, GBT build classification trees sequentially, where each new tree is trained to optimize a loss function based on the residual errors of the previous trees (Friedman 2001).
+Let $R$ be the number of boosting iterations.
+Given covariates $\mathbf{s}$, for the $r$-th tree where $r\in[R]$, let $h_{r}(\mathbf{s})$ be the prediction and $\gamma_{cr}$ be the weight for class $c$.
+The predicted probability for class $c$ in a GBT classifier is obtained using the softmax function of the log-odds:
+$\hat{p}_{c}(\mathbf{s})=\frac{e^{G_{c}(\mathbf{s})}}{\sum_{c^{\prime}=1}^{C}e^{G_{c^{\prime}}(\mathbf{s})}},$ (34)
+with the log-odds defined as:
+$G_{c}(\mathbf{s})=\sum_{r=1}^{R}\gamma_{cr}h_{r}(\mathbf{s}).$ (35)
+15 Point Prediction ML Models
+The regression models employed to generate point predictions are presented below.
+  1. 1.
+    Linear Regression (LinReg): LinReg solves the Ordinary Least Squares (OLS) problem.
+  2. 2.
+    Ridge Regression (Ridge): Ridge adds an $L_{2}$ norm regularization term to the coefficients obtained from OLS.
+  3. 3.
+    Least Absolute Shrinkage and Selection Operator Regression (Lasso): Lasso adds an $L_{1}$ norm regularization term to the coefficients obtained from OLS.
+  4. 4.
+    Random Forest (RF-Reg): RF-Reg operates similarly to the Random Forest classifier described in Section 14, but it builds a collection of regression trees instead.
+  5. 5.
+    Gradient Boosted Trees (CatBoost-Reg and XGBoost-Reg): Two versions of Gradient Boosted Trees (GBT) are implemented for regression.
+These GBT models are similar to the GBT classifiers described in Appendix 14, except they use regression trees.
+CatBoost-Reg is implemented using the CatBoost package, which builds symmetric trees and has native support for categorical features (Prokhorenkova et al.
+2018).
+XGBoost-Reg is implemented using the XGBoost package (Chen and Guestrin 2016).
+Table 5: Average Out‐of‐Sample MSE of Regression Models.
+Model LinReg Ridge Lasso RF-Reg CatBoost-Reg XGBoost-Reg
+MSE 2.324 2.324 3.149 1.791 1.668 1.666
+16 Hyperparameters of the ML Models
+Tables 6, 7, and 8 detail the hyperparameters and corresponding search spaces employed during the model selection and tuning processes for the ML algorithms considered in this study.
+Table 6: Hyperparameters for Point Prediction Regression Models
+Model Hyperparameter Range
+Ridge $\alpha$ ($L_{2}$ regularization) $1e-4\leq\alpha\leq 10.0$
+Lasso $\alpha$ ($L_{1}$ regularization) $1e-4\leq\alpha\leq 10.0$
+RF-Reg n_estimators $100\leq\text{n\_estimators}\leq 1000$ (step: 100)
+max_depth $3\leq\text{max\_depth}\leq 10$
+min_samples_split $2\leq\text{min\_samples\_split}\leq 10$
+min_samples_leaf $1\leq\text{min\_samples\_leaf}\leq 20$
+CatBoost-Reg iterations $50\leq\text{iterations}\leq 300$
+depth $4\leq\text{depth}\leq 10$
+learning_rate $0.01\leq\text{learning\_rate}\leq 0.3$
+l2_leaf_reg $1\leq\text{l2\_leaf\_reg}\leq 10$
+bagging_temperature $0.0\leq\text{bagging\_temperature}\leq 1.0$
+random_strength $0.0\leq\text{random\_strength}\leq 10.0$
+border_count $1\leq\text{border\_count}\leq 255$
+colsample_bylevel $0.5\leq\text{colsample\_bylevel}\leq 1.0$
+XGBoost-Reg n_estimators $100\leq\text{n\_estimators}\leq 1000$ (step: 100)
+max_depth $3\leq\text{max\_depth}\leq 10$
+learning_rate $0.01\leq\text{learning\_rate}\leq 0.3$
+subsample $0.5\leq\text{subsample}\leq 1.0$
+colsample_bytree $0.5\leq\text{colsample\_bytree}\leq 1.0$
+gamma $0\leq\text{gamma}\leq 5$
+reg_alpha $0.0\leq\text{reg\_alpha}\leq 100.0$
+reg_lambda $0.0\leq\text{reg\_lambda}\leq 100.0$
+min_child_weight $1\leq\text{min\_child\_weight}\leq 10$
+max_delta_step $0\leq\text{max\_delta\_step}\leq 10$
+colsample_bylevel $0.5\leq\text{colsample\_bylevel}\leq 1.0$
+colsample_bynode $0.5\leq\text{colsample\_bynode}\leq 1.0$
+Table 7: Hyperparameters for Classification Models
+Model Hyperparameter Range
+MLR C (regularization) $1e-4\leq\text{C}\leq 1e2$
+Classification Tree max_depth $3\leq\text{max\_depth}\leq 10$
+min_samples_leaf $10\leq\text{min\_samples\_leaf}\leq 30$ (step: 5)
+RF-MC-CLF n_estimators $600\leq\text{n\_estimators}\leq 1000$ (step: 200)
+max_depth $3\leq\text{max\_depth}\leq 10$
+min_samples_split $2\leq\text{min\_samples\_split}\leq 10$
+min_samples_leaf $10\leq\text{min\_samples\_leaf}\leq 30$ (step: 5)
+CatBoost-MC-CLF iterations $50\leq\text{iterations}\leq 300$
+depth $4\leq\text{depth}\leq 10$
+learning_rate $0.01\leq\text{learning\_rate}\leq 0.3$
+l2_leaf_reg $1\leq\text{l2\_leaf\_reg}\leq 10$
+bagging_temperature $0.0\leq\text{bagging\_temperature}\leq 1.0$
+random_strength $0.0\leq\text{random\_strength}\leq 10.0$
+border_count $1\leq\text{border\_count}\leq 255$
+colsample_bylevel $0.5\leq\text{colsample\_bylevel}\leq 1.0$
+Table 8: Hyperparameters for Quantile Regression Models
+Model Hyperparameter Range
+Regression Tree max_depth $3\leq\text{max\_depth}\leq 10$
+min_samples_leaf $10\leq\text{min\_samples\_leaf}\leq 30$ (step: 10)
+QRF n_estimators $600\leq\text{n\_estimators}\leq 1000$ (step: 200)
+max_depth $3\leq\text{max\_depth}\leq 10$
+min_samples_split $2\leq\text{min\_samples\_split}\leq 10$
+min_samples_leaf $10\leq\text{min\_samples\_leaf}\leq 30$ (step: 5)
+17 Evaluation of the Simulator
+The simulator is compared against a baseline that uses multinomial logistic regression (MLR), a common approach in the literature (Salari et al. 2022).
+Table 9 reports the average CRPS across all carriers.
+The proposed calibrated MC-CLF simulator achieves a CRPS of 0.314, compared with 0.779 for the MLR baseline, indicating a substantial improvement in distributional accuracy at the individual‐instance level.
+Figure 16 overlays each simulator’s joint predicted distribution on the true empirical distribution from the test set.
+The proposed calibrated MC-CLF simulator closely matches the empirical mass at zero deviation and reproduces both early‐ and late‐delivery tails, whereas the MLR baseline systematically overpredicts late deviations and underpredicts on‐time deliveries.
+Figure 17 presents three carrier‐specific examples.
+In each case, the proposed calibrated MC-CLF simulator aligns tightly with the carrier’s empirical distribution across all deviation classes.
+By contrast, the MLR baseline exhibits clear misalignments—particularly in the majority classes and tails—underscoring the proposed simulator’s superior ability to capture carrier‐level heterogeneity.
+Table 9: Average In-Sample CRPS Across All Carriers.
+Model Calibrated MC-CLF Simulator MLR Simulator (Baseline)
+CRPS 0.314 0.779
+(a) Calibrated MC-CLF Simulator
+(b) Baseline
+Figure 16: Overall Joint Predicted Distribution of the Simulators vs. the True Empirical Distribution
+(a) Example Carrier 1: Calibrated MC-CLF Simulator
+(b) Example Carrier 1: Baseline
+(c) Example Carrier 2: Calibrated MC-CLF Simulator
+(d) Example Carrier 2: Baseline
+(e) Example Carrier 3: Calibrated MC-CLF Simulator
+(f) Example Carrier 3: Baseline
+Figure 17: Selected Carrier Examples: Joint Predicted Distribution of the Simulators vs. the True Empirical Distribution
+18 Additional Details on Computational Scalability
+Table 10 provides a comprehensive comparison of objective values across different instance categories for the six proposed CSO methods.
+Table 11 summarizes the computation times for the baseline models and the proposed CSO-CLF methods.
+Table 10: Realized Average ($\pm$ 95% CI margin of error) and Worst‐Case Objective Values Across Instance Categories.
+Instance C-SAA-CLF C-SAA-QRF C-RO-B-CLF C-RO-B-QRF C-RO-D-CLF C-RO-D-QRF Avg Worst Avg Worst Avg Worst Avg Worst Avg Worst Avg Worst Small $0.821\pm 0.003$ 0.790 $0.848\pm 0.004$ 0.813 $0.851\pm 0.004$ 0.794 $0.871\pm 0.003$ 0.813 $0.936\pm 0.004$ 0.757 $0.920\pm 0.004$ 0.766 Medium $0.799\pm 0.008$ 0.932 $0.806\pm 0.009$ 0.925 $0.797\pm 0.009$ 0.834 $0.815\pm 0.009$ 0.821 $0.890\pm 0.010$ 0.822 $0.883\pm 0.009$ 0.825 Large $0.847\pm 0.012$ 1.032 $0.871\pm 0.017$ 1.055 $0.823\pm 0.014$ 0.982 $0.836\pm 0.016$ 0.981 $0.910\pm 0.017$ 0.972 $0.904\pm 0.014$ 0.973 Total $0.820\pm 0.003$ 0.823 $0.844\pm 0.003$ 0.842 $0.842\pm 0.003$ 0.811 $0.861\pm 0.003$ 0.824 $0.928\pm 0.003$ 0.779 $0.914\pm 0.003$ 0.787
+Table 11: Computation Time Statistics (in seconds) Across Instance Categories.
+Instance C-SAA-CLF C-RO-B-CLF C-RO-D-CLF C-Empirical-SAA Empirical-SAA PTO Greedy Avg Std Max Avg Std Max Avg Std Max Avg Std Max Avg Std Max Avg Std Max Avg Std Max Small 1.620 2.267 28.021 0.016 0.024 0.374 0.199 0.301 5.475 2.520 3.552 46.90 2.510 3.547 46.86 0.014 0.021 0.684 0.001 0.001 0.008 Medium 27.831 10.245 107.347 0.281 0.106 0.860 3.229 2.008 39.222 43.742 16.475 165.717 43.864 16.678 179.577 0.243 0.103 1.027 0.013 0.007 0.060 Large 59.266 22.590 182.401 0.698 0.322 2.856 7.683 7.056 111.188 91.190 35.100 278.677 92.225 35.421 295.212 0.629 0.329 2.584 0.052 0.034 0.317 Total 8.974 17.433 182.401 0.096 0.203 2.856 1.101 2.791 111.188 13.959 27.052 278.677 14.033 27.304 295.212 0.086 0.187 2.584 0.006 0.016 0.317
+18.1 Sample Size Selection: Balancing Solution Quality and Computational Efficiency
+Figures 18 and 19 illustrate the trade-off between average realized objective value and computation time for the two C-SAA methods.
+Similarly, Figures 20 and 21 depict the trade-off between worst-case realized objective value and computation time for the two C-RO-D methods.
+(a) Small Instances.
+(b) Medium Instances.
+(c) Large Instances.
+Figure 18: Computation Time by C-SAA Sample Size.
+(a) Small Instances.
+(b) Medium Instances.
+(c) Large Instances.
+Figure 19: Average Realized Objective Value by C-SAA Sample Size.
+(a) Small Instances.
+(b) Medium Instances.
+(c) Large Instances.
+Figure 20: Computation Time by C-RO-D Sample Size.
+(a) Small Instances.
+(b) Medium Instances.
+(c) Large Instances.
+Figure 21: Worst-Case Realized Objective Value by C-RO-D Sample Size.